### v2

# A defect-chemistry-informed phase-field model of grain growth in oxide electroceramics

Abstract
Dopants can significantly affect the properties of oxide ceramics through their impact on the property-determined microstructure characteristics such as grain boundary segregation, space charge layer formation in the grain boundary vicinity, and the resultant microstructure features like bimodality due to abnormal grain growth. To support rational oxide ceramics design, we propose a multiphysics-based and defect-chemistry-informed phase-field grain growth model to simulate the microstructure evolution of oxide ceramics. It fully respects the thermodynamics of charged point defects (oxygen vacancies and dopants) in both the grain interior and boundaries and considers the competing kinetics of defect diffusion and grain boundary movement. The proposed phase-field model is benchmarked against well-known simplified bicrystal models, including the Mott-Schottky and Gouy-Chapman models. Various simulation results are presented to reveal the impacts of defect formation energy differences between the grain interior and the grain boundary core on the key microstructural aspects. In particular, simulation results confirm that the solute drag effect alone can lead to bimodal grain size distribution, without any contribution from grain misorientation and other anisotropy. Interestingly, abnormal grain growth simulations demonstrate that grain boundary potentials can vary substantially: grain boundaries of larger grains tend to have lower potentials than those of smaller grains. Such heterogeneous grain boundary potential distribution may inspire a new material optimization strategy through microstructure design. This study provides a comprehensive framework for defect-chemistry-consistent investigations of microstructure evolution in polycrystalline oxide ceramics, offering fundamental insights into in-situ processes during critical manufacturing stages.
keywords:
Phase-field method , Defect-chemistry , Space charge layer , Abnormal grain growth
{graphicalabstract}[Uncaptioned image]
1 Introduction
In recent decades, oxide ceramics, such as SrTiO3 (STO) and BaTiO3 (BTO), have undergone tremendous development and are increasingly crucial in applications, such as solid oxide fuel cells [1, 2], sensors [3, 4], capacitors [5, 6], actuators [7, 8], and memristors [9, 10]. Oxide ceramics can be fabricated via sintering or deposition processes [11, 12, 13, 14]. When manufactured through sintering, oxide ceramics exhibit a polycrystalline microstructure, with their macroscopic properties heavily influenced by the synthesis conditions [15]. To tailor the physical and chemical properties of oxide ceramics, heterovalent substitution (doping) is an effective method for introducing additional charged defects through ionic compensation during synthesis. The state-of-the-art designs are mostly empirical, since knowledge of the correlation among doping strategies, processing and the resultant microstructure is limited. The influential mechanisms of doping on microstructure in oxide ceramics are manifold. First of all, the addition of doping elements in the parent material systems triggers defect chemistry reaction and results in multiple types of charged point defects or defect complexes, which interact with the grain boundary (GB) chemically, electrically and elastically and form space charge layers (SCLs) around GB. More specifically, the equilibrium concentrations of the point defects within the bulk are thermodynamically determined through defect chemistry [16, 17]. The scenario near inside the GB core and in the GB vicinity are different. In undoped and acceptor doped perovskites such as Fe-doped STO, the GB core is positively charged with excess oxygen vacancies, notated $\text{V}_{\text{O}}^{\text{..}}$ in Kröger-Vink notation [18], stemming from the negative formation energy difference of oxygen vacancy between GB core and bulk, i.e. $\Delta g_{\text{V}_{\text{O}}^{..}}=\mu^{0}_{\text{V}_{\text{O}}^{..},\text{c}}-\mu^{0}_{\text{V}_{\text{O}}^{..},\text{b}}<0$. To compensate the positively charged GB core, a negatively charged region forms adjacent to it, called the space charge layer (SCL).
Sharp interface models have been used to describe the SCL and the related grain boundary potential. Among them are the Mott-Schottky (MS) model and the Gouy-Chapman (GC) model. The MS model assumes that the dopant concentration is constant, and only the oxygen vacancies are able to redistribute, in order to form the SCLs. In contrast, the GC model has both defects as mobile and thus allows both to contribute to charge compensation. Unlike oxygen vacancies, acceptor dopant forms an accumulation region within the SCL and in the GB core due to their negative charge [Fig. 1(e)]. The spatial inhomogeneity introduced by SCL formation exerts a profound impact on the transport dynamics of charge species. Evidence shows that cation diffusion parallel to the GB is expedited [19, 20], whereas charge transport across the GB is significantly hindered [21].
SCLs determine not only the GB local properties of the synthesized material, but also impact the grain growth dynamics already during the synthesis, e.g., the well-known solute drag effect. Grain growth in oxide ceramics, driven by a reduction in surface area and controlled by the diffusion of charged species, plays a crucial role in determining the final grain size distribution and the concentration distributions of various point defects. Segregation in a moving GB core necessitates additional free energy dissipation and results in the solute drag effect [22, 23]. This effect significantly interferes with the kinetics of grain growth and can independently trigger abnormal grain growth even in the absence of any texture or pinning sites [24]. Dopant segregation to GB core has been experimentally evidenced, especially at high temperatures [25, 26]. Experiments also show that even 0.2% amount of Fe in STO can pin certain portion of GBs [Fig. 1(b)]. This trend increases with 2% Fe [Fig. 1(c)]. When increasing the dopant concentration to 5% [Fig. 1(d)], the grain growth is completely suppressed [25, 27]. In addition, the grain growth coefficient is found to follow a non-Arrhenius behavior between $1350^{\circ}$ and $1425^{\circ}$. Two types of GB core with different mobilities are proposed to explain the non-Arrhenius behavior [28, 25]. Despite various evidence supporting the impact of SCLs on GB properties and grain growth dynamics, experimental characterization of the SCL remains a challenge [29]. The complexity arises from the inherent structural and compositional heterogeneities present in polycrystalline oxide ceramics. One of the primary obstacles in this characterization is the limitation of in-situ transmission electron microscopy (TEM) investigations. Such studies are typically unfeasible due to the high sintering temperatures required for these ceramics, which exceed the operational limits of most TEM setups.
It is hence desirable to employ theoretical models and numerical simulations to evaluate SCLs and abnormal grain growth. De Souza derived numerical solutions based on defect chemistry in acceptor-doped oxide systems at equilibrium state [30]. The impacts of different charge species, the formation energy difference of oxygen vacancy and dopant between GB core and bulk phase, temperature and partial pressure of oxygen on grain boundary potential are comprehensively studied in a bicrystal with planar GB core. However, applying these numerical solutions to polycrystalline cases is impractical. Under this circumstance, the phase-field method with a diffuse interface emerges as a powerful tool to solve different moving boundary problems [31, 32]. There are a series of phase-field sintering models, e.g. to consider the solute drag effect in metals [24] and ceramics [33] and for non-isothermal sintering [34, 35, 36, 37]. But these models do not consider the charged defects, electrostatics or defect chemistry, which are important for oxide ceramics. Recently, Vikrant et al. developed a multi-physics phase-field model with an emphasis on GB misorientation and investigate the SCL formation at equilibrium and quasi-equilibrium state [38]. Then, this model is applied to explain the abnormal grain growth with a bimodal pattern as observed in experiments [39]. However, in this phase-field model the defect chemistry is not explicitly included, and thus there is no simulation results dedicated to the influence of defect energetics on SCLs and abnormal grain growth.
[…]
In the present work, we develop a phase-field grain growth model founded on the principle of defect-chemistry to simulate SCL formation and abnormal grain growth during sintering of oxide electroceramics. By taking into account the formation energy differences and the equilibrium defect concentrations in bulk and GB core, the thermodynamic descriptions of point defects are included in the phase-field model, along with the kinetics of grain boundary motion and defect diffusion. The spatio-temporal distributions of defect concentration and electrostatic potential in the grain interior and at GB are then solved self-consistently.
The outline of this paper is as follows. Section 2 presents the fundamentals of defect chemistry and the classical sharp interface bicrystal models, i.e., the GC and MS models, for they serve as important benchmarks in later sections. Section 3 describes the grain growth phase-field model in the principle of defect chemistry and electrostatics. Then, we demonstrate the implementation of the phase-field model via the finite element method in Section 4. The capabilities of this model are discussed for bicrystal and polycrystalline cases in Section 5. Thereby the solute drag induced abnormal grain growth simulation results are systematically presented and discussed, particularly in terms of the impact of segregation energies. It is followed by a concluding Section 6.
2 Defect chemistry of oxide ceramics
The basics of defect chemistry and the sharp interface bicrystal models (the MS and the GC models) for the equilibrium state were well described in Refs. [30, 40, 16]. Thereby the GB is treated as sharp interface with GB core properties. Thus the concentration profiles of oxygen vacancy and dopant are discontinuous, with concentration jumps across the GB core. For the bicrystal, it becomes a one-dimensional case.
Assume a coordinate system $x$, with the origin $x=0$ located at the GB core. Moreover, point defects are assumed to be dilute and non-interacting in both GB core and bulk phases. Therefore, the electrochemical potentials of oxygen vacancy and dopant in the bulk and in the core region can be formulated [30] as
$\mu_{\text{V}_{\text{O}}^{..},\text{b}}(x)=\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{0}+k_{\text{B}}T\ln\left[\frac{c_{\text{V}_{\text{O}}^{..},\text{b}}(x)}{N_{\text{V}_{\text{O}}^{..},\text{b}}-c_{\text{V}_{\text{O}}^{..},\text{b}}(x)}\right]+z_{\text{V}_{\text{O}}^{..}}e\phi(x),$ (1)
$\mu_{\text{dop},\text{b}}(x)=\mu_{\text{dop},\text{b}}^{0}+k_{\text{B}}T\ln\left[\frac{c_{\text{dop},\text{b}}(x)}{N_{\text{dop},\text{b}}-c_{\text{dop},\text{b}}(x)}\right]+z_{\text{dop}}e\phi(x),$ (2)
$\mu_{\text{V}_{\text{O}}^{..},\text{c}}=\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{0}+k_{\text{B}}T\ln\left[\frac{c_{\text{V}_{\text{O}}^{..},\text{c}}}{N_{\text{V}_{\text{O}}^{..},\text{c}}-c_{\text{V}_{\text{O}}^{..},\text{c}}}\right]+z_{\text{V}_{\text{O}}^{..}}e\Phi_{0},$ (3)
$\mu_{\text{dop},\text{c}}=\mu_{\text{dop},\text{c}}^{0}+k_{\text{B}}T\ln\left[\frac{c_{\text{dop},\text{c}}}{N_{\text{dop},\text{c}}-c_{\text{dop},\text{c}}}\right]+z_{\text{dop}}e\Phi_{0},$ (4)
where $\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{0}$, $\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{0}$, $\mu_{\text{dop},\text{b}}^{0}$ and $\mu_{\text{dop},\text{c}}^{0}$ denote the standard formation energies of oxygen vacancy and dopant in the bulk and in the core, respectively. The symbols $N_{\text{V}_{\text{O}}^{..},\text{b}}^{0}$, $N_{\text{V}_{\text{O}}^{..},\text{c}}^{0}$, $N_{\text{dop},\text{b}}^{0}$ and $N_{\text{dop},\text{c}}^{0}$ indicate the number of available sites per unit volume of oxygen vacancy and dopant in the bulk and core, respectively. $z_{\text{V}_{\text{O}}^{..}}$ and $z_{\text{dop}}$ denote the charge numbers of oxygen vacancy and acceptor dopant. $\phi(x)$ is the electrostatic potential, and $\Phi_{0}$ is the electrostatic potential in the GB core. Similarly, $c_{\text{dop},\text{c}}$ and $c_{\text{V}_{\text{O}}^{..},\text{c}}$ are the concentrations of dopant and oxygen vacancy in the GB core. Thus, we have $\Phi_{0}=\phi(0)$, $c_{\text{dop},\text{c}}=c_{\text{dop},\text{b}}(0)$ and $c_{\text{V}_{\text{O}}^{..},\text{c}}=c_{\text{V}_{\text{O}}^{..},\text{b}}(0)$. Additionally, $k_{\text{B}}$, $e$ and $T$ are the Boltzmann constant, elementary charge and temperature, respectively. The formation of a SCL requires the segregation of oxygen vacancy in the core. The smaller value of $\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{0}$ than $\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{0}$, i.e. $\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{0}-\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{0}<0$, drives the formation of SCL. At the equilibrium state, when electrochemical potentials should be everywhere the same (in other words, no gradient), we have
$\mu_{\text{V}_{\text{O}}^{..},\text{c}}=\mu_{\text{V}_{\text{O}}^{..},\text{b}}(x)=\mu_{\text{V}_{\text{O}}^{..},\text{b}}(\infty),$ (5)
$\mu_{\text{dop},\text{c}}=\mu_{\text{dop},\text{b}}(x)=\mu_{\text{dop},\text{b}}(\infty),$ (6)
Hence, equilibrium concentration profiles of oxygen vacancy and dopant in the SCL take the form
$c_{\text{V}_{\text{O}}^{..},\text{b}}(x)=c_{\text{V}_{\text{O}}^{..},\text{b}}(\infty)\exp\left[\frac{-z_{\text{V}_{\text{O}}^{..}}e\{\phi(x)-\phi(\infty)\}}{k_{\text{B}}T}\right],$ (7)
$c_{\text{dop},\text{b}}(x)=\frac{c_{\text{dop},\text{b}}(\infty)N_{\text{dop},\text{b}}\exp\left[\frac{-z_{\text{dop}}e\{\phi(x)-\phi(\infty)\}}{k_{\text{B}}T}\right]}{N_{\text{dop},\text{b}}+c_{\text{dop},\text{b}}(\infty)\left\{\exp\left[\frac{-z_{\text{dop}}e\{\phi(x)-\phi(\infty)\}}{k_{\text{B}}T}\right]-1\right\}}.$ (8)
In analogy, equilibrium concentrations of oxygen vacancy and dopant in the core are
$c_{\text{V}_{\text{O}}^{..},\text{c}}=\frac{N_{\text{V}_{\text{O}}^{..},\text{c}}c_{\text{V}_{\text{O}}^{..},\text{b}}(\infty)\exp\left[-\frac{\Delta g_{\text{V}_{\text{O}}^{..}}+z_{\text{V}_{\text{O}}^{..}}e\Phi_{0}}{k_{\text{B}}T}\right]}{N_{\text{V}_{\text{O}}^{..},\text{b}}+c_{\text{V}_{\text{O}}^{..},\text{b}}(\infty)\exp\left[-\frac{\Delta g_{\text{V}_{\text{O}}^{..}}+z_{\text{V}_{\text{O}}^{..}}e\Phi_{0}}{k_{\text{B}}T}\right]},$ (9)
$c_{\text{dop},\text{c}}=\frac{N_{\text{dop},\text{c}}c_{\text{dop},\text{b}}(\infty)\exp\left[-\frac{\Delta g_{\text{dop}}+z_{\text{dop}}e\Phi_{0}}{k_{\text{B}}T}\right]}{N_{\text{dop},\text{b}}+c_{\text{dop},\text{b}}(\infty)\exp\left[-\frac{\Delta g_{\text{dop}}+z_{\text{dop}}e\Phi_{0}}{k_{\text{B}}T}\right]},$ (10)
where $\Delta g_{\text{V}_{\text{O}}^{..}}=\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{0}-\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{0}$ and $\Delta g_{\text{dop}}=\mu_{\text{dop},\text{c}}^{0}-\mu_{\text{dop},\text{b}}^{0}$. All parameters are tabulated in Table 1.
Table 1: Summary of the parameters in the sharp interface method
Parameter Definition Unit
$\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{0}$ Formation energy of oxygen vacancy in the bulk phase eV
$\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{0}$ Formation energy of oxygen vacancy in the core region eV
$\mu_{\text{dop},\text{b}}^{0}$ Formation energy of dopant in the bulk phase eV
$\mu_{\text{dop},\text{c}}^{0}$ Formation energy of dopant in the core region eV
$\Delta g_{\text{V}_{\text{O}}^{..}}$ The difference of formation energy of oxygen vacancy between core and bulk, i.e. $\Delta g_{\text{V}_{\text{O}}^{..}}=\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{0}-\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{0}$ eV
$\Delta g_{\text{dop}}$ The difference of formation energy of dopant between core and bulk, i.e. $\Delta g_{\text{dop}}=\mu_{\text{dop},\text{c}}^{0}-\mu_{\text{dop},\text{b}}^{0}$ eV
$N_{\text{V}_{\text{O}}^{..},\text{b}}$ Maximum number density of oxygen vacancy in the bulk phase $\frac{1}{\text{m}^{3}}$
$N_{\text{V}_{\text{O}}^{..},\text{c}}$ Maximum number density of oxygen vacancy in the core region $\frac{1}{\text{m}^{3}}$
$N_{\text{dop},\text{b}}$ Maximum number density of dopant in the bulk phase. $N_{\text{dop},\text{b}}=1/a^{3}$, with $a$ being the perovskite lattice constant $\frac{1}{\text{m}^{3}}$
$N_{\text{dop},\text{c}}$ Maximum number density of dopant in the core region $\frac{1}{\text{m}^{3}}$
$z_{\text{V}_{\text{O}}^{..}}$ Valency number of oxygen vacancy, $z_{\text{V}_{\text{O}}^{..}}=2$
$z_{\text{dop}}$ Valency number of dopant, $z_{\text{dop}}=-1$ for $\text{Fe}_{\text{Ti}}^{{}^{\prime}}$ in Fe-doped STO -
$c_{\text{dop},\text{b}}(\infty)$ Number density of dopant in the bulk far from the core region $\frac{1}{\text{m}^{3}}$
$c_{\text{V}_{\text{O}}^{..},\text{b}}(\infty)$ Number density of oxygen vacancy in the bulk far from the core region $\frac{1}{\text{m}^{3}}$
$k_{\text{B}}$ Boltzmann constant $k_{\text{B}}=1.38\times 10^{-23}$ $\frac{\text{J}}{\text{K}}$
$T$ Temperature K
$\epsilon_{0}$ Vacuum permittivity $\epsilon_{0}=8.854\times 10^{-12}$ $\frac{\text{C}}{\text{Vm}}$
$\epsilon_{\text{r}}$ Relative permittivity -
$w_{\text{c}}$ width of the core m
2.1 The Mott-Schottky model
The Mott-Schottky model assumes that the dopant is immobile and its concentration remains fixed and constant everywhere, i.e. $\nabla c_{\text{dop}}=0$. Note that this assumption simplifies the scenario but breaks the equilibrium condition of the electrochemical potential of the dopants. Based on the assumption, Poisson’s equation can be written as
$\begin{split}\epsilon_{0}\epsilon_{\text{r}}\frac{\partial^{2}\phi}{\partial x^{2}}&=-\rho(x)\\ &=-e[z_{\text{dop}}c_{\text{dop},\text{b}}(\infty)+z_{\text{V}_{\text{O}}^{..}}c_{\text{V}_{\text{O}}^{..},\text{b}}(x)].\end{split}$ (11)
[…]
$\epsilon_{0}\epsilon_{\text{r}}\frac{\partial^{2}\phi}{\partial x^{2}}=2ec_{\text{V}_{\text{O}}^{..},\text{b}}(\infty)\left\{1-\exp\left[\frac{-2e\{\phi(x)-\phi(\infty)\}}{k_{\text{B}}T}\right]\right\}.$ (12)
[…]
$\frac{\partial^{2}\phi}{\partial x^{2}}=\frac{1}{2\alpha l_{\text{D}}^{2}}\left\{1-\exp\left[\frac{-2e\{\phi(x)-\phi(\infty)\}}{k_{\text{B}}T}\right]\right\},$ (13)
2.2 The Gouy-Chapman model
In the Gouy-Chapman model both point defects (oxygen vacancies and dopant) are mobile. The electrochemical potential of the acceptor dopant is also considered to be constant, as is that of the oxygen vacancies. Thus, Poisson’s equation becomes
$\begin{split}\epsilon_{0}\epsilon_{\text{r}}\frac{\partial^{2}\phi}{\partial x^{2}}&=-\rho(x)\\ &=-e[z_{\text{dop}}c_{\text{dop},\text{b}}(x)+z_{\text{V}_{\text{O}}^{..}}c_{\text{V}_{\text{O}}^{..},\text{b}}(x)].\end{split}$ (14)
Substituting Eq. (7) and (8) into (14), we have
$\begin{split}\epsilon_{0}\epsilon_{\text{r}}\frac{\partial^{2}\phi}{\partial x^{2}}&=-e\left\{z_{\text{dop}}\frac{c_{\text{dop},\text{b}}(\infty)N_{\text{dop},\text{b}}\exp\left[\frac{-z_{\text{dop}}e\{\phi(x)-\phi(\infty)\}}{k_{\text{B}}T}\right]}{N_{\text{dop},\text{b}}+c_{\text{dop},\text{b}}(\infty)\left\{\exp\left[\frac{-z_{\text{dop}}e\{\phi(x)-\phi(\infty)\}}{k_{\text{B}}T}\right]-1\right\}}\right.\\ &\left.+z_{\text{V}_{\text{O}}^{..}}c_{\text{V}_{\text{O}}^{..},\text{b}}(\infty)\exp\left[\frac{-z_{\text{V}_{\text{O}}^{..}}e\{\phi(x)-\phi(\infty)\}}{k_{\text{B}}T}\right]\right\}.\end{split}$ (15)
The numerical solutions of the MS and GC models have been achieved through an iteration process [30]. In the present work, we propose an alternative numerical method via finite element method (FEM). By considering the symmetric configuration of a bicrysal (ignoring the grain misorientation), the semi-infinite simulation domain of $[0,\infty)$ is sufficient. The location of the GB core is located at the left boundary of the semi-infinite domain ($x=0$). As for the boundary conditions, the electrostatic field is grounded at the right side, i.e., $\phi|_{x=\infty}=0$, while at the left side, a mixed boundary condition is given by
$\frac{\partial\phi}{\partial x}\bigg{|}_{x=0}=-\frac{Q_{\text{c}}(\phi_{0})}{2\epsilon_{0}\epsilon_{\text{r}}},$ (16)
where $Q_{\text{c}}(\phi_{0})=ew_{\text{c}}(z_{\text{V}_{\text{O}}^{..}}c_{\text{V}_{\text{O}}^{..},\text{c}}+z_{\text{dop}}c_{\text{dop},\text{c}})$ and $\phi_{0}=\phi|_{x=0}$. We employed the boundary element in the finite element method to regard directly the mixed type of boundary condition defined in the above equation. The obtained numerical solutions agree fully with these from Ref. [30] (See Supplementary).
3 Phase-field grain growth model with defect chemistry and electrostatics
3.1 Free energy density functional and electrochemical potentials
We include here the defect chemistry into the grain growth model for oxide ceramics, with Fe-doped STO as a prototype. Upon grain contact in Fe-doped STO, the lower formation energy of oxygen vacancy in the core compels point defects to segregate towards the GB core. Afterwards, the depletion or accumulation zone appears in the vicinity of GB core. The SCL forms through a diffusion process. The handling of SCL is similarly addressed in Ref. [30]. The segregation of point defects in the core serves to dissipate the total free energy of the system and also induces solute drag effect when the core moves.
Assuming an Fe-doped STO polycrystalline system consisting of $n$ grains, we choose Kim-Kim-Suzuki (KKS) phase-field model and utilize a set of non-conserved order parameters (OPs) ${\eta_{i}}$ to distinguish between different grains [41]. Within grain $i$, $\eta_{i}=1$ and $\eta_{j}=0$ for $j<n$ and $j\neq i$. Within the core region of grain $i$ and $j$, the order parameters $\eta_{i}$ and $\eta_{j}$ spatially vary from 0 to 1. In the latter context, $\eta$ denotes the entire set of OPs $\eta_{i}$}. When considering electrochemical contributions, we employ two conserved concentration fields along with an electrostatic potential field. Other normalized field variables are $C_{\text{V}_{\text{O}}^{..}}$ for oxygen vacancy concentration, $C_{\text{dop}}$ for dopant concentration, and $\phi$ for the electrostatic potential. In Fe-doped STO, $\text{Fe}^{3+}$ and $\text{Fe}^{4+}$ can coexist. At high temperatures and low oxygen partial pressures, a higher proportion of $\text{Fe}^{3+}$ is observed and thus the dopant $\text{Fe}_{\text{Ti}}^{{}^{\prime}}$ has negative charge. For example, during the sintering process at 1623 K in air, approximately 90% of the iron is estimated to be in the $\text{Fe}^{3+}$ state, according to the bulk defect chemical model (see Fig. A1 in Ref. [42]). Thus, in this study, we assume the valence state of Fe to be 3+ and therefore the valance of dopant $\text{Fe}_{\text{Ti}}^{{}^{\prime}}$ is -1, i.e. $z_{\text{dop}}=-1$. Then, the formulation of total free energy density functional is given by
$\mathscr{F}=\int_{\Omega}(f^{\text{grad}}+f^{\text{loc}}+f^{\text{ech}})\text{d}\Omega.$ (17)
$f^{\text{grad}}$ represents the gradient term, it is given by
$f^{\text{grad}}=\frac{1}{2}\kappa\sum_{i=1}^{n}|\nabla\eta_{i}|^{2},$ (18)
where $\kappa$ is the coefficient of the gradient term. $f^{\text{loc}}$ is the well potential. According to Ref. [43], its formulation is given by
$f^{\text{loc}}=\omega\sum_{i=1}^{n}\left(\frac{\eta_{i}^{4}}{4}-\frac{\eta_{i}^{2}}{2}+\gamma\sum_{i=1}^{n}\sum_{j>i}^{n}\eta_{i}^{2}\eta_{j}^{2}+\frac{1}{4}\right),$ (19)
with $\gamma=1.5$ for producing symmetric profile of $\eta$ and isotropic grain growth.
The determination of parameters such as $\kappa$ and $\omega$ can be linked to characteristics such as core width ($w_{\text{c}}$) and interface energy ($\sigma$). Their relations can be expressed as
$\sigma=\frac{\sqrt{2}}{3}\sqrt{\kappa\omega},$ (20)
and
$w_{\text{c}}=\sqrt{\frac{8\kappa}{\omega}}.$ (21)
In the present work, we only consider the isotropic interface energy, i.e. it is not dependent on GB misorientation. The third term in free energy functional is the electrochemical contribution. Here, we treat the electrochemical free energy density as a mixture of bulk phase and core by an interpolation function. Thus, $f^{\text{ech}}$ is
$f^{\text{ech}}=[1-h(\eta)]f^{\text{ech}}_{\text{b}}(C_{\text{V}_{\text{O}}^{..},\text{b}},C_{\text{dop},\text{b}})+h(\eta)f^{\text{ech}}_{\text{c}}(C_{\text{V}_{\text{O}}^{..},\text{c}},C_{\text{dop},\text{c}})-\frac{\epsilon_{0}\epsilon_{\text{r}}}{2}(\nabla\phi)^{2},$ (22)
where the subscripts b and c indicate the bulk phase and core, respectively. $h(\eta)$ is the interpolation function and is given by
$h(\eta)=\frac{4}{3}\left[1-4\sum_{i=1}^{n}\eta_{i}^{3}+3\left(\sum_{i=1}^{n}\eta_{i}^{2}\right)^{2}\right].$ (23)
Here, $h(\eta)$ exhibits non-monotonic behavior. Specifically, for an $i/j$ GB, $h(\eta_{i})=0$ when $\eta_{i}=1$ (representing the bulk phase), whereas $h(\eta_{i})=1$ when $\eta_{i}=0.5$ (representing the core). This interpolation function allows us to differentiate between the bulk phase and the core region. $\epsilon_{0}$ is the vacuum permittivity; $\epsilon_{\text{r}}$ is the relative permittivity. For STO, $\epsilon_{\text{r}}=9\times 10^{4}/(T-35)$ [30]. $\epsilon_{\text{r}}$ is only dependent on temperature in the present work. Hence, $\epsilon_{\text{r}}$ is constant across the SCL zone for specific temperature. Furthermore, $f^{\text{ech}}_{\text{b}}(C_{\text{V}_{\text{O}}^{..},\text{b}},C_{\text{dop},\text{b}})$ and $f^{\text{ech}}_{\text{c}}(C_{\text{V}_{\text{O}}^{..},\text{c}},C_{\text{dop},\text{c}})$ in Eq. 22 denote the electrochemical free energy contributions from the bulk and core, respectively, with $C_{\text{V}_{\text{O}}^{..},\text{b}}$ and $C_{\text{dop},\text{b}}$ being the phase concentrations of oxygen vacancy and dopant in the bulk phase, while $C_{\text{V}_{\text{O}}^{..},\text{c}}$ and $C_{\text{dop},\text{c}}$ being the phase concentrations of oxygen vacancy and dopant in the core. When $h(\eta)=1$, $C_{\text{dop},\text{c}}$ and $C_{\text{V}_{\text{O}}^{..},\text{c}}$ correspond to $c_{\text{dop},\text{c}}$ and $c_{\text{V}_{\text{O}}^{..},\text{c}}$. When $h(\eta)=0$, $C_{\text{dop},\text{b}}$ and $C_{\text{V}_{\text{O}}^{..},\text{b}}$ are same as $c_{\text{dop},\text{b}}(x)$ and $c_{\text{V}_{\text{O}}^{..},\text{b}}(x)$. Local concentration is treated as the mixture of phase concentrations in bulk and core. The relations between local concentration and phase concentration are expressed as
$C_{\text{V}_{\text{O}}^{..}}=[1-h(\eta)]C_{\text{V}_{\text{O}}^{..},\text{b}}+h(\eta)C_{\text{V}_{\text{O}}^{..},\text{c}},$ (24)
$C_{\text{dop}}=[1-h(\eta)]C_{\text{dop},\text{b}}+h(\eta)C_{\text{dop},\text{c}}.$ (25)
In principle of the thermodynamic model extended by Ref. [30], the expressions for the electrochemical free energies $f^{\text{ech}}_{\text{c}}(C_{\text{V}_{\text{O}}^{..},\text{c}},C_{\text{dop},\text{c}})$ and $f^{\text{ech}}_{\text{b}}(C_{\text{V}_{\text{O}}^{..},\text{b}},C_{\text{dop},\text{b}})$ are given by
$\begin{split}f^{\text{ech}}_{\text{c}}(C_{\text{V}_{\text{O}}^{..},\text{c}},C_{\text{dop},\text{c}})&=\frac{1}{V_{\text{m}}}\left\{g^{0}_{\text{V}_{\text{O}}^{..},\text{c}}C_{\text{V}_{\text{O}}^{..},\text{c}}+g^{0}_{\text{dop},\text{c}}C_{\text{dop},\text{c}}\right.\\ &+RT\left[C_{\text{V}_{\text{O}}^{..},\text{c}}\ln(C_{\text{V}_{\text{O}}^{..},\text{c}})+(\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}-C_{\text{V}_{\text{O}}^{..},\text{c}})\ln(\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}-C_{\text{V}_{\text{O}}^{..},\text{c}})\right.\\ &\left.\left.+C_{\text{dop},\text{c}}\ln(C_{\text{dop},\text{c}})+(\tilde{N}_{\text{dop},\text{c}}-C_{\text{dop},\text{c}})\ln(\tilde{N}_{\text{dop},\text{c}}-C_{\text{dop},\text{c}})\right]\right\}\\ &+\frac{\mathcal{F}}{V_{\text{m}}}\left(z_{\text{V}_{\text{O}}^{..}}C_{\text{V}_{\text{O}}^{..},\text{c}}+z_{\text{dop}}C_{\text{dop},\text{c}}\right)\phi,\end{split}$ (26)
$\begin{split}f^{\text{ech}}_{\text{b}}(C_{\text{V}_{\text{O}}^{..},\text{b}},C_{\text{dop},\text{b}})&=\frac{1}{V_{\text{m}}}\left\{g^{0}_{\text{V}_{\text{O}}^{..},\text{b}}C_{\text{V}_{\text{O}}^{..},\text{b}}+g^{0}_{\text{dop},\text{b}}C_{\text{dop},\text{b}}\right.\\ &+RT\left[C_{\text{V}_{\text{O}}^{..},\text{b}}\ln(C_{\text{V}_{\text{O}}^{..},\text{b}})+(\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}-C_{\text{V}_{\text{O}}^{..},\text{b}})\ln(\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}-C_{\text{V}_{\text{O}}^{..},\text{b}})\right.\\ &\left.\left.+C_{\text{dop},\text{b}}\ln(C_{\text{dop},\text{b}})+(\tilde{N}_{\text{dop},\text{b}}-C_{\text{dop},\text{b}})\ln(\tilde{N}_{\text{dop},\text{b}}-C_{\text{dop},\text{b}})\right]\right\}\\ &+\frac{\mathcal{F}}{V_{\text{m}}}\left(z_{\text{V}_{\text{O}}^{..}}C_{\text{V}_{\text{O}}^{..},\text{b}}+z_{\text{dop}}C_{\text{dop},\text{b}}\right)\phi,\end{split}$ (27)
where $g^{0}_{\text{V}_{\text{O}}^{..},\text{c}}$, $g^{0}_{\text{V}_{\text{O}}^{..},\text{b}}$, $g^{0}_{\text{dop},\text{c}}$, $g^{0}_{\text{dop},\text{b}}$ are standard formation energies of oxygen vacancy and dopant in bulk and core. $\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}$, $\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}$, $\tilde{N}_{\text{dop},\text{c}}$, $\tilde{N}_{\text{dop},\text{b}}$ are dimensionless numbers of available sites per unit volume for oxygen vacancy and dopant in bulk and core. $z_{\text{V}_{\text{O}}^{..}}$ and $z_{\text{dop}}$ are valance numbers of oxygen vacancy and dopant. $\mathcal{F}$, $V_{\text{m}}$, $R$, $T$ are the Faraday constant, molar volume, gas constant and temperature, respectively. The electrochemical potentials of oxygen vacancy and dopant in bulk and core are obtained by
$\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{\text{PF}}=\frac{\delta\mathscr{F}}{\delta C_{\text{V}_{\text{O}}^{..},\text{b}}}=\frac{g_{\text{V}_{\text{O}}^{..},\text{b}}^{0}}{V_{\text{m}}}+\frac{RT}{V_{\text{m}}}\ln{\left[\frac{C_{\text{V}_{\text{O}}^{..},\text{b}}}{\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}-C_{\text{V}_{\text{O}}^{..},\text{b}}}\right]}+\frac{\mathcal{F}}{V_{\text{m}}}z_{\text{V}_{\text{O}}^{..}}\phi,$ (28)
$\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{\text{PF}}=\frac{\delta\mathscr{F}}{\delta C_{\text{V}_{\text{O}}^{..},\text{c}}}=\frac{g_{\text{V}_{\text{O}}^{..},\text{c}}^{0}}{V_{\text{m}}}+\frac{RT}{V_{\text{m}}}\ln{\left[\frac{C_{\text{V}_{\text{O}}^{..},\text{c}}}{\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}-C_{\text{V}_{\text{O}}^{..},\text{c}}}\right]}+\frac{\mathcal{F}}{V_{\text{m}}}z_{\text{V}_{\text{O}}^{..}}\phi,$ (29)
$\mu_{\text{dop},\text{b}}^{\text{PF}}=\frac{\delta\mathscr{F}}{\delta C_{\text{dop},\text{b}}}=\frac{g_{\text{dop},\text{b}}^{0}}{V_{\text{m}}}+\frac{RT}{V_{\text{m}}}\ln{\left[\frac{C_{\text{dop},\text{b}}}{\tilde{N}_{\text{dop},\text{b}}-C_{\text{dop},\text{b}}}\right]}+\frac{\mathcal{F}}{V_{\text{m}}}z_{\text{dop}}\phi,$ (30)
$\mu_{\text{dop},\text{c}}^{\text{PF}}=\frac{\delta\mathscr{F}}{\delta C_{\text{dop},\text{c}}}=\frac{g_{\text{dop},\text{c}}^{0}}{V_{\text{m}}}+\frac{RT}{V_{\text{m}}}\ln{\left[\frac{C_{\text{dop},\text{c}}}{\tilde{N}_{\text{dop},\text{c}}-C_{\text{dop},\text{c}}}\right]}+\frac{\mathcal{F}}{V_{\text{m}}}z_{\text{dop}}\phi,$ (31)
which are consistent with the standard formations of electrochemical potentials in defect chemistry, i.e. Eq. 1 to Eq. 4.
3.2 Governing equations
In the framework of KKS phase-field model, the evolution of phase-field OPs obeys the Allen-Cahn equation. For grain $i$, the evolution of $\eta_{i}$ in time is
$\frac{1}{L}\frac{\partial\eta_{i}}{\partial t}=\frac{\delta\mathscr{F}}{\delta\eta_{i}},$ (32)
where $L$ is the GB mobility. Substituting Eq. (17) into (32), we can obtain
$\begin{split}\frac{1}{L}\frac{\partial\eta_{i}}{\partial t}&=\nabla\cdot\kappa\nabla\eta_{i}-\omega\frac{\partial f^{\text{loc}}(\eta)}{\partial\eta_{i}}\\ &+\frac{\partial h(\eta)}{\partial\eta_{i}}\left[f_{\text{b}}^{\text{ech}}-f_{\text{c}}^{\text{ech}}-\frac{\partial f_{\text{b}}^{\text{ech}}}{\partial C_{\text{V}_{\text{O}}^{..},\text{b}}}(C_{\text{V}_{\text{O}}^{..},\text{b}}-C_{\text{V}_{\text{O}}^{..},\text{c}})-\frac{\partial f_{\text{b}}^{\text{ech}}}{\partial C_{\text{dop},\text{b}}}(C_{\text{dop},\text{b}}-C_{\text{dop},\text{c}})\right].\end{split}$ (33)
At the quasi-equilibrium state, the core moves with a constant velocity, $v_{c}$, we have $\dot{\eta}=-v_{c}\eta^{\prime}(x)$, thus Eq. (32) becomes
$\begin{split}\frac{v_{c}}{L}\frac{d\eta}{dx}&=\nabla\cdot\kappa\nabla\eta_{i}-\omega\frac{\partial f^{\text{loc}}(\eta)}{\partial\eta_{i}}\\ &+\frac{\partial h(\eta)}{\partial\eta_{i}}\left[f_{\text{b}}^{\text{ech}}-f_{\text{c}}^{\text{ech}}-\frac{\partial f_{\text{b}}^{\text{ech}}}{\partial C_{\text{V}_{\text{O}}^{..},\text{b}}}(C_{\text{V}_{\text{O}}^{..},\text{b}}-C_{\text{V}_{\text{O}}^{..},\text{c}})-\frac{\partial f_{\text{b}}^{\text{ech}}}{\partial C_{\text{dop},\text{b}}}(C_{\text{dop},\text{b}}-C_{\text{dop},\text{c}})\right],\end{split}$ (34)
Conserved concentration fields $C_{\text{V}_{\text{O}}^{..}}$ and $C_{\text{dop}}$ follow Fick’s law, which are expressed as
$\frac{\partial C_{\text{V}_{\text{O}}^{..}}}{\partial t}=\nabla\left(M_{\text{V}_{\text{O}}^{..}}\nabla\frac{\delta\mathscr{F}}{\delta C_{\text{V}_{\text{O}}^{..}}}\right),$ (35)
$\frac{\partial C_{\text{dop}}}{\partial t}=\nabla\left(M_{\text{dop}}\nabla\frac{\delta\mathscr{F}}{\delta C_{\text{dop}}}\right),$ (36)
where $M_{\text{V}_{\text{O}}^{..}}$ and $M_{\text{dop}}$ are mobilities of oxygen vacancy and dopant, which are given by $M_{\text{V}_{\text{O}}^{..}}=\frac{D_{\text{V}_{\text{O}}^{..}}}{\partial^{2}f^{\text{ech}}/\partial C_{\text{V}_{\text{O}}^{..}}^{2}}$ and $M_{\text{dop}}=\frac{D_{\text{dop}}}{\partial^{2}f^{\text{ech}}/\partial C_{\text{dop}}^{2}}$, with $D_{\text{V}_{\text{O}}^{..}}$ and $D_{\text{dop}}$ being the diffusivities of oxygen vacancy and dopant, respectively. Moreover, local equilibrium between bulk phases and core should be satisfied everywhere. Therefore, we have following constrains
$\delta\mathscr{F}/\delta C_{\text{V}_{\text{O}}^{..}}=\delta\mathscr{F}/\delta C_{\text{V}_{\text{O}}^{..},\text{b}}=\delta\mathscr{F}/\delta C_{\text{dop},\text{b}},$ (37)
$\delta\mathscr{F}/\delta C_{\text{dop}}=\delta\mathscr{F}/\delta C_{\text{V}_{\text{O}}^{..},\text{c}}=\delta\mathscr{F}/\delta C_{\text{dop},\text{c}}.$ (38)
Therefore, $\delta\mathscr{F}/\delta C_{\text{V}_{\text{O}}^{..}}$ and $\delta\mathscr{F}/\delta C_{\text{dop}}$ in Eqs. 35 and 36 can be substituted by $\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{\text{PF}}$, $\mu_{\text{dop},\text{b}}^{\text{PF}}$ or $\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{\text{PF}}$, $\mu_{\text{dop},\text{c}}^{\text{PF}}$. In addition, the governing equation of electrostatic potential field is obtained by Poisson’s equation
$\epsilon_{0}\epsilon_{\text{r}}\nabla^{2}\phi+\frac{\mathcal{F}}{V_{\text{m}}}(z_{\text{V}_{\text{O}}^{..}}C_{\text{V}_{\text{O}}^{..}}+z_{\text{dop}}C_{\text{dop}})=0,$ (39)
The parameters utilized in phase-field model are tabulated in table 2.
Table 2: Summary of the parameters in the phase-field model
DOFs Definition Unit
$\eta_{i}$ phase-field order parameter of grain $i$, $i=1,2,3,...,n$
$C_{\text{V}_{\text{O}}^{..}}$ Normalized local concentration of oxygen vacancy ($C_{\text{V}_{\text{O}}^{..}}=a^{3}c_{\text{V}_{\text{O}}^{..}}$, with $a$ being the lattice constant) -
$C_{\text{dop}}$ Normalized local concentration of dopant, $C_{\text{dop}}=a^{3}c_{\text{dop}}$
$C_{\text{V}_{\text{O}}^{..},\text{b}}$ Normalized concentration of oxygen vacancy in bulk phase, $C_{\text{V}_{\text{O}}^{..},\text{b}}=a^{3}c_{\text{V}_{\text{O}}^{..},\text{b}}$
$C_{\text{V}_{\text{O}}^{..},\text{c}}$ Normalized concentration of oxygen vacancy in core region, $C_{\text{V}_{\text{O}}^{..},\text{c}}=a^{3}c_{\text{V}_{\text{O}}^{..},\text{c}}$
$C_{\text{dop},\text{b}}$ Normalized concentration of dopant in bulk phase, $C_{\text{dop},\text{b}}=a^{3}c_{\text{dop},\text{b}}$
$C_{\text{dop},\text{c}}$ Normalized concentration of dopant in core region, $C_{\text{dop},\text{c}}=a^{3}c_{\text{dop},\text{c}}$
${\phi}$ Electrostatic potential V
Parameters Definition Unit
$\kappa$ Coefficient of gradient term $\frac{\text{J}}{\text{m}}$
$\omega$ Heigh of well potential $\frac{\text{J}}{\text{m}^{3}}$
$L$ GB mobility $\frac{\text{m}^{3}}{\text{Js}}$
$g^{0}_{\text{V}_{\text{O}}^{..},\text{b}}$ Formation energy of oxygen vacancy in the bulk phase. $g^{0}_{\text{V}_{\text{O}}^{..},\text{b}}=\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{0}N_{\text{A}}$, with $N_{\text{A}}$ being Avogadro constant $\frac{\text{J}}{\text{mol}}$
$g_{\text{V}_{\text{O}}^{..},\text{c}}^{0}$ Formation energy of oxygen vacancy in the core region. $g^{0}_{\text{V}_{\text{O}}^{..},\text{c}}=\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{0}N_{\text{A}}$ $\frac{\text{J}}{\text{mol}}$
$g^{0}_{\text{dop},\text{b}}$ Formation energy of dopant in the bulk phase. $g^{0}_{\text{dop},\text{b}}=\mu_{\text{dop},\text{b}}^{0}N_{\text{A}}$ $\frac{\text{J}}{\text{mol}}$
$g^{0}_{\text{dop},\text{c}}$ Formation energy of dopant in the core region. $g^{0}_{\text{dop},\text{c}}=\mu_{\text{dop},\text{c}}^{0}N_{\text{A}}$ $\frac{\text{J}}{\text{mol}}$
$V_{\text{m}}$ Molar volume $\frac{\text{m}^{3}}{\text{mol}}$
$\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}$ Dimensionless available sites of oxygen vacancy in core region. $\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}=N_{\text{V}_{\text{O}}^{..},\text{c}}a^{3}$, with $a$ being the lattice constant -
$\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}$ Dimensionless available sites of oxygen vacancy in bulk phase. $\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}=N_{\text{V}_{\text{O}}^{..},\text{b}}a^{3}$
$\tilde{N}_{\text{dop},\text{c}}$ Dimensionless available sites of dopant in core region. $\tilde{N}_{\text{dop},\text{c}}=N_{\text{dop},\text{c}}a^{3}$
$\tilde{N}_{\text{dop},\text{b}}$ Dimensionless available sites of dopant in bulk phase. $\tilde{N}_{\text{dop},\text{b}}=N_{\text{dop},\text{b}}a^{3}$
$R$ Gas constant $\frac{\text{J}}{\text{K}}$
$T$ Temperature K
$\mathcal{F}$ Faraday constant $\frac{\text{C}}{\text{mol}}$
$D_{\text{V}_{\text{O}}^{..}}$ Diffusivity of oxygen vacancy $\frac{\text{m}^{2}}{\text{s}}$
$D_{\text{dop}}$ Diffusivity of dopant $\frac{\text{m}^{2}}{\text{s}}$
4 Finite element implementation of the phase-field model
In this section, we present the detailed finite element implementation for the defect-chemistry informed phase-field model. According to the model presented in Section 3 and Table 2, we treat the OPs $\eta_{i}$ and the field variables $C_{\text{V}_{\text{O}}^{..}}$, $C_{\text{dop}}$, $C_{\text{V}_{\text{O}}^{..},\text{b}}$, $C_{\text{V}_{\text{O}}^{..},\text{c}}$, $C_{\text{dop},\text{b}}$, $C_{\text{dop},\text{c}}$, $\phi$ as degrees-of-freedom (DOFs). Moreover, to solve Eqs. (35) and (36), we use the mixed finite element formulation and treat the electro-chemical potentials $\mu_{\text{V}_{\text{O}}^{..}}$ and $\mu_{\text{dop}}$ as additional DOFs [44].
The strong forms of the corresponding governing equations have already been given in Eqs. (24), (25) and (33) to (39). According to the FEM [45], we formulate the residuals from the weak forms of these governing equations by introducing corresponding test functions $\psi_{C_{\text{V}_{\text{O}}^{..},\text{c}}}$, $\psi_{C_{\text{dop},\text{c}}}$, $\psi_{\eta_{i}}$, $\psi_{C_{\text{V}_{\text{O}}^{..}}}$, $\psi_{C_{\text{dop}}}$, $\psi_{\mu_{\text{V}_{\text{O}}^{..}}}$, $\psi_{\mu_{\text{dop}}}$, $\psi_{C_{\text{V}_{\text{O}}^{..},\text{b}}}$, $\psi_{C_{\text{dop},\text{b}}}$ and $\psi_{\phi}$. Firstly, the residuals for the weak forms of Eqs. (24), (25) and (33) are
$R_{C_{\text{V}_{\text{O}}^{..},\text{c}}}=\int_{V}\psi_{C_{\text{V}_{\text{O}}^{..},\text{c}}}\{C_{\text{V}_{\text{O}}^{..}}-[1-h(\eta)]C_{\text{V}_{\text{O}}^{..},\text{b}}-h(\eta)C_{\text{V}_{\text{O}}^{..},\text{c}}\}\text{d}V,$ (40)
$R_{C_{\text{dop},\text{c}}}=\int_{V}\psi_{C_{\text{dop},\text{c}}}\{C_{\text{dop}}-[1-h(\eta)]C_{\text{dop},\text{b}}-h(\eta)C_{\text{dop},\text{c}}\}\text{d}V,$ (41)
$\begin{split}R_{\eta_{i}}&=\int_{V}{\psi_{\eta_{i}}}\frac{\eta_{i}^{t_{n+1}}-\eta_{i}^{t_{n}}}{\Delta t}\text{d}V+L\int_{V}\psi_{\eta_{i}}\omega\frac{\partial f^{\text{loc}}(\eta)}{\partial\eta_{i}}\text{d}V+L\int_{V}\kappa\eta_{i,j}\psi_{\eta_{i},j}\text{d}V-L\int_{\Gamma}\kappa\eta_{i,j}\psi_{\eta_{i}}\hat{n}_{j}\text{d}\Gamma\\ &-L\int_{V}\psi_{\eta_{i}}\frac{\partial h(\eta)}{\partial\eta_{i}}\left[f_{\text{b}}^{\text{ech}}-f_{\text{c}}^{\text{ech}}-\frac{\partial f_{\text{b}}^{\text{ech}}}{\partial C_{\text{V}_{\text{O}}^{..},\text{b}}}(C_{\text{V}_{\text{O}}^{..},\text{b}}-C_{\text{V}_{\text{O}}^{..},\text{c}})-\frac{\partial f_{\text{b}}^{\text{ech}}}{\partial C_{\text{dop},\text{b}}}(C_{\text{dop},\text{b}}-C_{\text{dop},\text{c}})\right]\text{d}V.\end{split}$ (42)
Note that we use the backward Euler method for the time integration. In the very last equation, $\eta_{i}^{t_{n}}$ and $\eta_{i}^{t_{n+1}}$ indicate $\eta_{i}$ at the time step $t_{n}$ and $t_{n+1}$, respectively. Thereby, $\Delta t=t_{n+1}-t_{n}$, and $\hat{n}_{i}$ is the normal vector to the boundary $\Gamma$ of the subdomain. Then, through introducing the addition coupling fields $\mu_{\text{V}_{\text{O}}^{..}}=\delta\mathscr{F}/\delta C_{\text{V}_{\text{O}}^{..}}$ and $\mu_{\text{dop}}=\delta\mathscr{F}/\delta C_{\text{dop}}$, we obtain the weak forms of Eqs. (35) and (36) as
$R_{C_{\text{V}_{\text{O}}^{..}}}=\int_{V}\psi_{C_{\text{V}_{\text{O}}^{..}}}\left(\mu_{\text{V}_{\text{O}}^{..}}-\frac{\delta\mathscr{F}}{\delta C_{\text{V}_{\text{O}}^{..},\text{b}}}\right)\text{d}V.$ (43)
$R_{C_{\text{dop}}}=\int_{V}\psi_{C_{\text{dop}}}\left(\mu_{\text{dop}}-\frac{\delta\mathscr{F}}{\delta C_{\text{dop},\text{b}}}\right)\text{d}V.$ (44)
$R_{\mu_{\text{V}_{\text{O}}^{..}}}=\int_{V}\psi_{\mu_{\text{V}_{\text{O}}^{..}}}\frac{C_{\text{V}_{\text{O}}^{..}}^{t_{n+1}}-C_{\text{V}_{\text{O}}^{..}}^{t_{n}}}{\Delta t}\text{d}V+\int_{V}M_{\text{V}_{\text{O}}^{..}}\mu_{V_{O}^{..},i}\psi_{\mu_{\text{V}_{\text{O}}^{..}},i}\text{d}V-\int_{\Gamma}M_{\text{V}_{\text{O}}^{..}}\mu_{V_{O}^{..},i}\psi_{\mu_{\text{V}_{\text{O}}^{..}}}\hat{n}_{i}\text{d}\Gamma.$ (45)
$R_{\mu_{\text{dop}}}=\int_{V}\psi_{\mu_{\text{dop}}}\frac{C_{\text{dop}}^{t_{n+1}}-C_{\text{dop}}^{t_{n}}}{\Delta t}\text{d}V+\int_{V}M_{\text{dop}}\mu_{dop,i}\psi_{\mu_{\text{dop}},i}\text{d}V-\int_{\Gamma}M_{\text{dop}}\mu_{dop,i}\psi_{\mu_{\text{dop}}}\hat{n}_{i}\text{d}\Gamma.$ (46)
The weak forms of the local equilibrium constrain in Eqs. (37) and (38) are
$R_{C_{\text{V}_{\text{O}}^{..},\text{b}}}=\int_{V}\psi_{C_{\text{V}_{\text{O}}^{..},\text{b}}}\left(\frac{\delta\mathscr{F}}{\delta C_{\text{V}_{\text{O}}^{..},\text{b}}}-\frac{\delta\mathscr{F}}{\delta C_{\text{V}_{\text{O}}^{..},\text{c}}}\right)\text{d}V.$ (47)
$R_{C_{\text{dop},\text{b}}}=\int_{V}\psi_{C_{\text{dop},\text{b}}}\left(\frac{\delta\mathscr{F}}{\delta C_{\text{dop},\text{b}}}-\frac{\delta\mathscr{F}}{\delta C_{\text{dop},\text{c}}}\right)\text{d}V.$ (48)
Finally, the weak form of Eq (39) is
$R_{\phi}=\int_{V}\epsilon_{0}\epsilon_{\text{r}}\phi_{,i}\psi_{\phi,i}\text{d}V-\int_{\Gamma}\epsilon_{0}\epsilon_{\text{r}}\phi_{,i}\psi_{\phi}\hat{n}_{i}\text{d}\Gamma-\int_{V}\psi_{\phi}\frac{\mathcal{F}}{V_{\text{m}}}(z_{\text{V}_{\text{O}}^{..}}C_{\text{V}_{\text{O}}^{..}}+z_{\text{dop}}C_{\text{dop}})\text{d}V.$ (49)
Following the Galerkin approach [46], the test functions are discretized as $\psi_{C_{\text{V}_{\text{O}}^{..},\text{c}}}=N_{C_{\text{V}_{\text{O}}^{..},\text{c}}}^{I}\psi_{C_{\text{V}_{\text{O}}^{..},\text{c}}}^{I}$, $\psi_{C_{\text{dop},\text{c}}}=N_{C_{\text{dop},\text{c}}}^{I}\psi_{C_{\text{dop},\text{c}}}^{I}$, $\psi_{\eta_{i}}=N_{\eta_{i}}^{I}\psi_{\eta_{i}}^{I}$, $\psi_{C_{\text{V}_{\text{O}}^{..}}}=N_{C_{\text{V}_{\text{O}}^{..}}}^{I}\psi_{C_{\text{V}_{\text{O}}^{..}}}^{I}$, $\psi_{C_{\text{dop}}}=N_{C_{\text{dop}}}\psi_{C_{\text{dop}}}^{I}$, $\psi_{\mu_{\text{V}_{\text{O}}^{..}}}=N_{\mu_{\text{V}_{\text{O}}^{..}}}\psi_{\mu_{\text{V}_{\text{O}}^{..}}}^{I}$, $\psi_{\mu_{\text{dop}}}=N_{\mu_{\text{dop}}}\psi_{\mu_{\text{dop}}}^{I}$, $\psi_{C_{\text{V}_{\text{O}}^{..},\text{b}}}=N_{C_{\text{V}_{\text{O}}^{..},\text{b}}}\psi_{C_{\text{V}_{\text{O}}^{..},\text{b}}}^{I}$, $\psi_{C_{\text{dop},\text{b}}}=N_{C_{\text{dop},\text{b}}}\psi_{C_{\text{dop},\text{b}}}^{I}$ and $\psi_{\phi}=N_{\phi}\psi_{\phi}^{I}$, where $I$ denotes the node index and Einstein summation convention is used. $\psi_{C_{\text{V}_{\text{O}}^{..},\text{c}}}^{I}$, $\psi_{C_{\text{dop},\text{c}}}^{I}$, $\psi_{\eta_{i}}^{I}$, $\psi_{C_{\text{V}_{\text{O}}^{..}}}^{I}$, $\psi_{C_{\text{dop}}}^{I}$, $\psi_{\mu_{\text{V}_{\text{O}}^{..}}}^{I}$, $\psi_{\mu_{\text{dop}}}^{I}$, $\psi_{C_{\text{V}_{\text{O}}^{..},\text{b}}}^{I}$, $\psi_{C_{\text{dop},\text{b}}}^{I}$ and $\psi_{\phi}^{I}$ are the nodal weights, while $N_{C_{\text{V}_{\text{O}}^{..},\text{c}}}^{I}$, $N_{C_{\text{dop},\text{c}}}^{I}$, $N_{\eta_{i}}^{I}$, $N_{C_{\text{V}_{\text{O}}^{..}}}^{I}$, $N_{C_{\text{dop}}}^{I}$, $N_{\mu_{\text{V}_{\text{O}}^{..}}}^{I}$, $N_{\mu_{\text{dop}}}^{I}$, $N_{C_{\text{V}_{\text{O}}^{..},\text{b}}}^{I}$, $N_{C_{\text{dop},\text{b}}}^{I}$ and $N_{\phi}^{I}$ are the shape functions for the corresponding variables. Then we can write down the discretized residuals as
$R_{C_{\text{V}_{\text{O}}^{..},\text{c}}}^{I}=\int_{V}N_{C_{\text{V}_{\text{O}}^{..},\text{c}}}^{I}\{C_{\text{V}_{\text{O}}^{..}}-[1-h(\eta)]C_{\text{V}_{\text{O}}^{..},\text{b}}-h(\eta)C_{\text{V}_{\text{O}}^{..},\text{c}}\}\text{d}V$ (50)
$R_{C_{\text{dop},\text{c}}}^{I}=\int_{V}N_{C_{\text{dop},\text{c}}}^{I}\{C_{\text{dop}}-[1-h(\eta)]C_{\text{dop},\text{b}}-h(\eta)C_{\text{dop},\text{c}}\}\text{d}V$ (51)
$R_{\eta_{i}}^{I}&=\int_{V}\frac{N_{\eta_{i}}^{I}}{\Delta t}\text{d}V+L\int_{V}N_{\eta_{i}}^{I}\omega\frac{\partial f^{\text{loc}}(\eta)}{\partial\eta_{i}}\text{d}V+L\int_{V}\kappa\eta_{i,j}N_{\eta_{i},j}^{I}\text{d}V-L\int_{\Gamma}\kappa\eta_{i,j}N_{\eta_{i}}^{I}\hat{n}_{j}\text{d}\Gamma\\ &-L\int_{V}N_{\eta_{i}}^{I}\frac{\partial h(\eta)}{\partial\eta_{i}}\left[f_{\text{b}}^{\text{ech}}-f_{\text{c}}^{\text{ech}}-\frac{\partial f_{\text{b}}^{\text{ech}}}{\partial C_{\text{V}_{\text{O}}^{..},\text{b}}}(C_{\text{V}_{\text{O}}^{..},\text{b}}-C_{\text{V}_{\text{O}}^{..},\text{c}})-\frac{\partial f_{\text{b}}^{\text{ech}}}{\partial C_{\text{dop},\text{b}}}(C_{\text{dop},\text{b}}-C_{\text{dop},\text{c}})\right]\text{d}V\end{aligned}$ (52)
$R_{C_{\text{V}_{\text{O}}^{..}}}^{I}=\int_{V}N_{C_{\text{V}_{\text{O}}^{..}}}^{I}\left(\mu_{\text{V}_{\text{O}}^{..}}-\frac{\delta\mathscr{F}}{\delta C_{\text{V}_{\text{O}}^{..},\text{b}}}\right)\text{d}V$ (53)
$R_{C_{\text{dop}}}^{I}=\int_{V}N_{C_{\text{dop}}}^{I}\left(\mu_{\text{dop}}-\frac{\delta\mathscr{F}}{\delta C_{\text{dop},\text{b}}}\right)\text{d}V$ (54)
$R_{\mu_{\text{V}_{\text{O}}^{..}}}^{I}=\int_{V}\frac{N_{\mu_{\text{V}_{\text{O}}^{..}}}^{I}}{\Delta t}\text{d}V+\int_{V}M_{\text{V}_{\text{O}}^{..}}\mu_{\text{V}_{\text{O}}^{..},i}N_{\mu_{\text{V}_{\text{O}}^{..}},i}^{I}\text{d}V-\int_{\Gamma}M_{\text{V}_{\text{O}}^{..}}\mu_{V_{O}^{..},i}N_{\mu_{\text{V}_{\text{O}}^{..}}}^{I}\hat{n}_{i}\text{d}\Gamma$ (55)
$R_{\mu_{\text{dop}}}^{I}=\int_{V}\frac{N_{\mu_{\text{dop}}}^{I}}{\Delta t}\text{d}V+\int_{V}M_{\text{dop}}\mu_{\text{dop},i}N_{\mu_{\text{dop}},i}^{I}\text{d}V-\int_{\Gamma}M_{\text{dop}}\mu_{\text{dop},i}N_{\mu_{\text{dop}}}^{I}\hat{n}_{i}\text{d}\Gamma$ (56)
$R_{C_{\text{V}_{\text{O}}^{..},\text{b}}}^{I}=\int_{V}N_{C_{\text{V}_{\text{O}}^{..},\text{b}}}^{I}\left(\frac{\delta\mathscr{F}}{\delta C_{\text{V}_{\text{O}}^{..},\text{b}}}-\frac{\delta\mathscr{F}}{\delta C_{\text{V}_{\text{O}}^{..},\text{c}}}\right)\text{d}V$ (57)
$R_{C_{\text{dop},\text{b}}}^{I}=\int_{V}N_{C_{\text{dop},\text{b}}}^{I}\left(\frac{\delta\mathscr{F}}{\delta C_{\text{dop},\text{b}}}-\frac{\delta\mathscr{F}}{\delta C_{\text{dop},\text{c}}}\right)\text{d}V$ (58)
$R_{\phi}^{I}=\int_{V}\epsilon_{0}\epsilon_{\text{r}}\phi_{,i}N_{\phi,i}^{I}\text{d}V-\int_{\Gamma}\epsilon_{0}\epsilon_{\text{r}}\phi_{,i}N_{\phi}^{I}\hat{n}_{i}\text{d}\Gamma-\int_{V}N_{\phi}^{I}\frac{\mathcal{F}}{V_{\text{m}}}(z_{\text{V}_{\text{O}}^{..}}C_{\text{V}_{\text{O}}^{..}}+z_{\text{dop}}C_{\text{dop}})\text{d}V$ (59)
The linearization of the residuals can be written in the element-level linear formulation, in which the terms in the tangent matrix are calculated by $K_{\xi,\zeta}^{IJ}=\partial R_{\xi}^{I}/\partial\zeta^{J}$, where ${\xi,\zeta}$ indicates individual DOF. The non-zero terms in the tangent matrix are presented in supplementary. The numerical implementation of the proposed phase-field model is carried out in Multiphysics Object-Oriented Simulation Environment (MOOSE) framework [47, 48].
5 Phase-field simulation results of Fe-doped SrTiO3
We present first in Section 5.1 one-dimensional phase-field simulation results to demonstrate its capability of reproducing the formation of a SCL at equilibrium state. Subsequently, we investigate the influence of solute drag effects induced by the segregation of point defects in the core on the formation of an asymmetric SCL under quasi-equilibrium state in Section 5.2. Finally, we conduct two-dimensional (2D) phase-field simulations to study the microstructure evolution process in STO polycrystalline material during sintering and analyze the resulting grain size distribution in Section 5.3.
Note that in the paper we ignore the GB orientation dependency or other anisotropy of the segregation energies $\Delta g_{\text{V}_{\text{O}}^{..}}$ and $\Delta g_{\text{dop}}$, for we intend to reveal the most fundamental mechanisms of SCL formation and abnormal grain growth. Without loss of generality, the proposed phase-field model can be extended accordingly to include such anisotropy aspects.
5.1 GB space charge layers at equilibrium
To demonstrate the capability of the phase-field model at thermodynamically equilibrium state, one-dimensional phase-field simulations are conducted and compared with the sharp interface predictions [30]. The semi-finite simulation domain has the size of 780 nm, with the core located at the left boundary. The grain boundary thickness is prescribed to be 0.78 nm [30]. The domain is discretized into a mesh consisting of 2000 elements. The simulation temperature is set at 600 K, and the relative permittivity $\epsilon_{\text{r}}$ of strontium titanate (STO) is specified as 160, according to [30], which is calculated via $\epsilon_{\text{r}}=9\times 10^{4}/(T-35)$ as mentioned in section 3.1. The interface energy, molar volume and interface mobility of STO are 0.6 $\text{J}\text{m}^{-2}$, 3.6×$\times 10^{-5}$ $\text{m}^{3}\text{mol}^{-1}$ and 1×$\times 10^{-7}$ $\text{m}^{3}\text{J}^{-1}\text{s}^{-1}$, respectively [38]. The relative tolerance for convergence is set to $1\times 10^{-8}$.
5.1.1 Benchmark case 1: The MS Model
By fixing a uniform dopant concentration, the proposed phase-field model can be used to simulate the MS case. The results are given in Fig. 2. The initial concentrations of oxygen vacancies and dopant far from the core (denoted as $c_{\text{dop},\text{b}}(\infty)$ and $c_{\text{V}_{\text{O}}^{..},\text{b}}(\infty)$, respectively) are $7\times 10^{24}\text{ m}^{-3}$ and $3.5\times 10^{24}\text{ m}^{-3}$ [30]. The number of available sites of oxygen vacancy in bulk phase and GB core are $N_{\text{V}_{\text{O}}^{..},\text{b}}=5.12\times 10^{28}$ $\text{m}^{-3}$ and $N_{\text{V}_{\text{O}}^{..},\text{c}}=3\times 10^{26}$ $\text{m}^{-3}$, respectively[30, 49]. The grain boundary potential $\Phi_{0}$ is plotted as a function of $\Delta g_{\text{V}_{\text{O}}^{..}}$ in Fig. 2(a). Due to the right boundary being grounded, the grain boundary potential is directly obtained at the interface position when $\eta=0.5$, as defined in the phase-field method. For $\Delta g_{\text{V}_{\text{O}}^{..}}$ larger than -1.25 eV, the phase-field results exhibit good agreement with analytical calculations. As $\Delta g_{\text{V}_{\text{O}}^{..}}$ becomes more negative, particularly at -1.5 eV and -2.0 eV, the available oxygen vacancy sites are completely exhausted. Therefore, the concentrations of oxygen vacancies in the core are close to $N_{\text{V}_{\text{O}}^{..},\text{c}}$. Although the phase-field results demonstrate convergent behavior, a discrepancy between the phase-field results and the analytical solution emerges. To explain the aforementioned discrepancy, we plot the concentration of oxygen vacancies in the core as a function of $\Delta g_{\text{V}_{\text{O}}^{..}}$ in Fig. 2(b). Similar to the grain boundary potential, $c_{\text{V}_{\text{O}}^{..}}$ is also extracted at $\eta=0.5$. We observe that the phase-field results coincide with the analytical predictions, even when $\Delta g_{\text{V}_{\text{O}}^{..}}=-2.0$ eV. Next, we examine the concentration of oxygen vacancy in the bulk in terms of $x$. Here, we plot the phase-field results for $\Delta g_{\text{V}_{\text{O}}^{..}}$=-0.5 eV, -1.0 eV and -1.5 eV in Fig. 2 (c), (d) and (e) in the range of 0 to 50 nm, which includes half of the interface region. For the analytical solution, we shift it to the right by a distance of $w_{\text{c}}/2$ to match the bulk phase in the phase-field results. When ${\Delta g_{\text{V}_{\text{O}}^{..}}}=-0.5$ eV, $c_{\text{V}_{\text{O}}^{..},\text{b}}$ obtained by phase-field simulations agrees well with analytical predictions as shown in Fig. 2(c1). However, in Fig. 2(d1) and (e1), for $\Delta g_{\text{V}_{\text{O}}^{..}}=-1.5$ and $-2.0$ eV, the phase-field results produce higher $c_{\text{V}_{\text{O}}^{..},\text{b}}$. This comparison indicates that less oxygen vacancy is segregated in the core in the phase-field simulations when choosing more negative $\Delta g_{\text{V}_{\text{O}}^{..}}$. Additionally, we plot the area density of the net charge, $Q_{\mathrm{net}}=\int_{-w_{\text{c}}/2}^{w_{\text{c}}/2}(z{{}_{\text{dop}}}c_{{}_{\text{dop},\text{c}}}+z_{{}_{\text{V}_{\text{O}}^{..}}}c_{{}_{\text{V}_{\text{O}}^{..},\text{c}}})\text{d}x$, in the core in Fig. 2(f). Less net charge is reproduced via phase-field simulation when $\Delta g_{\text{V}_{\text{O}}^{..}}$ is smaller. In short, this discrepancy is attributed to the diffuse interface in the phase-field simulations. Due to its finite interface thickness and smooth interface characteristics, the core is less positively charged compared to analytical solutions. Consequently, the electrostatic potential profile obtained via phase-field simulations decays faster than analytical calculations as shown in Fig. 2(c2), (d2) and (e2). Up to this point, we clearly observe a good agreement between the oxygen vacancy concentrations in the GB core and bulk phase as predicted by the phase-field results and the sharp interface predictions. The different approaches to handling interfaces in the phase-field model and the sharp interface method can result in discrepancies when the available densities of point defects are fully occupied.
Figure 2: Comparison of phase-field (PF) simulation outcomes with those of the sharp interface (SI) model predictions obtained via FEM and taken from Ref. [30] in the MS model. Figure (a) illustrates the grain boundary potential as a function of $\Delta g_{\text{V}_{\text{O}}^{..}}$. Figure (b) depicts the relationship between the oxygen vacancy concentration and $\Delta g_{\text{V}_{\text{O}}^{..}}$. Figures (c), (d), and (e) present the concentration distribution of dopant and the electrostatic potential distribution for different oxygen vacancy segregation energy. Figure (f) demonstrates the net area density of charge in the core as a function of $\Delta g_{\text{V}_{\text{O}}^{..}}$.
5.1.2 Benchmark case 2: The GC Model
By setting the electrochemical potentials of the acceptor dopant and of the oxygen vacancies to be constant, the presented phase-field model is then reduced to the GC model, in order to reproduce the equilibrium SCL formation in the GC case. The number of available sites of dopant and oxygen vacancy in bulk and GB core are $N_{\text{V}_{\text{O}}^{..},\text{c}}=1\times 10^{27}$ $\text{m}^{-3}$, $N_{\text{V}_{\text{O}}^{..},\text{b}}=5.12\times 10^{28}$ $\text{m}^{-3}$, $N_{\text{dop},\text{c}}=1.68\times 10^{27}$ $\text{m}^{-3}$ and $N_{\text{dop},\text{b}}=1.69\times 10^{28}$ $\text{m}^{-3}$, respectively [30, 49]. The Debye lengths $l_{\text{D}}=\sqrt{\frac{\epsilon_{0}\epsilon_{\text{r}}k_{\text{B}}T}{2e^{2}c_{\text{dop},\text{b}}(\infty)}}$ in both MS and GC models are 5.71 nm. In Fig. 3(a), the grain boundary potentials obtained via the phase-field simulations are compared with the sharp interface predictions. $\Delta g_{\text{V}_{\text{O}}^{..}}$ is fixed at -1.0 eV here.
When $\Delta g_{\text{dop}}$ is positive, dopant cations are rejected from the core, thereby causing a more pronounced grain boundary potential. Conversely, as $\Delta g_{\text{dop}}$ becomes negative, dopant segregation to the core occurs, reducing the net charge and consequently decreasing the grain boundary potential. Moreover, as $\Delta g_{\text{dop}}$ becomes sufficiently negative, the available number sites for dopant segregating within the core become scarce. Therefore, the grain boundary potential profile is flat when $\Delta g_{\text{dop}}$ is positive or sufficiently negative. The comparison indicates the phase-field results agree with the predictions of analytical solution. The discrepancy is smaller than 7%. The concentrations of vacancies and acceptor dopants in the core as a function of $\Delta g_{\text{dop}}$ are plotted in Fig. 3(b).
In Fig. 3(c), (d), and (e), we present the concentration profiles of oxygen vacancies and dopant, as well as the distribution of electrostatic potential in the bulk phase for $\Delta g_{\text{dop}}$ = 0 eV, -0.125 eV and -0.375 eV, respectively. Due to symmetry, we only depict half of the SCL in one bulk phase.
At $\Delta g_{\text{dop}}$ = 0 eV, phase-field results predict the formation of a dopant accumulation zone near the core. The dopant concentration in the core can be much lower than in the accumulation zone, leading to the formation of a nonmonotonic curve as shown in Fig. 3(c1). Contrarily, in Virkant’s findings [38], the expected accumulation zone at $\Delta g_{\text{dop}}$=0 eV is replaced by a depletion zone. As $\Delta g_{\text{dop}}$ decreases to -0.125 eV, dopant concentration in the core increases significantly, causing the formation of monotonic concentration profile. Further decreasing $\Delta g_{\text{dop}}$ to -0.375 eV, the dopant concentration in the core increases to $N_{\text{dop},\text{b}}$. It is evident that the accumulation zone becomes lower than 0 eV, more exactly -0.125 eV. Additionally, near the core, the concentration profile of dopant exhibits a distinct kink, indicating a notable change in the gradient.
The oxygen vacancy concentration profiles of are plotted in Fig. 3(c2), (d2) and (e2). Depletion appears for all $\Delta g_{\text{dop}}$. Notably, as $\Delta g_{\text{dop}}$ decreases, there is an increase in the concentrations within these depletion zones. In Fig. 3(c3), (d3) and (e3), we present a comparison of the electrostatic potential across the SCL with analytical solutions. The phase-field simulation results agree well with these analytical predictions. Decreasing $\Delta g_{\text{dop}}$ causes lower electrostatic potential distribution near the core. Moreover, we also present the net charge area density $Q_{\mathrm{net}}$ in Fig. 3(f). The trend of the curve obtained via phase-field results agrees with the analytical predictions. Analogous to the MS case, phase-field results are smaller than analytical predictions.
Figure 3: Comparison of phase-field (PF) simulation results with the sharp interface (SI) results via FEM method for the GC case. Figures (a) and (b) showcase the variations in grain boundary potential, dopant concentration, and oxygen vacancy concentration in the core as functions of $\Delta g_{\text{dop}}$. Figures (c), (d), and (e) detail the spatial distributions of dopant and oxygen vacancy concentrations in the bulk phase and the electrostatic potential. Figure (f) depicts the area density of net charge in the core, also as a function of $\Delta g_{\text{dop}}$. $\Delta g_{\text{V}_{\text{O}}^{..}}$ was kept at -1.0 eV for all calculations.
5.2 GB space charge layers at quasi-equilibrium
In this section, we investigate systematically the solute drag effect caused by an asymmetric SCL at a quasi-statically moving GB in a STO bicrystal. The solute drag phenomenon arises when point defects, segregated within a moving core in a quasi-equilibrium state, lead to additional free energy dissipation to maintain motion. At low core velocities, these defects can keep pace with the core, significantly increasing the necessary driving force. However, beyond a critical velocity, the defects lag behind, remaining in the bulk phases and causing the breakdown of symmetric concentration and electrostatic potential profiles across the SCL.
To investigate the formation of asymmetric SCL due to solute drag effects under quasi-equilibrium conditions, a constant core velocity (${v}_{\text{c}}$) along the simulation domain is assigned as a parameter in the simulations. The velocity of the core ranges from 0.01 $\text{nm}\,\text{s}^{-1}$ to 100 $\text{nm}\,\text{s}^{-1}$ to study its influence on the SCL. The defect chemistry parameters $N_{\text{V}_{\text{O}}^{..},\text{c}}$, $N_{\text{V}_{\text{O}}^{..},\text{b}}$, $N_{\text{dop},\text{c}}$ and $N_{\text{dop},\text{b}}$ in this case are $1\times 10^{27}$ $\text{m}^{-3}$, $5.12\times 10^{28}$ $\text{m}^{-3}$, $1.68\times 10^{27}$ $\text{m}^{-3}$ and $1.69\times 10^{28}$ $\text{m}^{-3}$, respectively. Simulations are conducted at 1623 K, with a relative permittivity ($\epsilon_{\text{r}}$) of STO set to 56, based on [38]. The diffusivities of oxygen vacancies and dopant in the bulk phase are specified as $1.35\times 10^{-9}$ $\text{m}^{2}\text{s}^{-1}$ and $1.3\times 10^{-11}$ $\text{m}^{2}\text{s}^{-1}$, respectively [38]. The simulation domain and boundary conditions remain consistent with the previous case.
In alignment with Vikrant [38], the total driving force $F_{\text{T}}$ for the core motion is given by
$F_{\text{T}}=\int_{-\infty}^{\infty}\frac{v_{c}}{L}\left(\frac{\partial\eta}{\partial x}\right)^{2}\text{d}x+\int_{-\infty}^{\infty}{\mu_{\text{V}_{\text{O}}^{..}}}\frac{\partial C_{\text{V}_{\text{O}}^{..}}}{\partial x}\text{d}x+\int_{-\infty}^{\infty}{\mu_{\text{dop}}}\frac{\partial C_{\text{dop}}}{\partial x}\text{d}x.$ (60)
The second and third terms in Eq. 60 denote as the electrochemical drag force, which accounts for the electrostatic and chemical contributions. The integration in Eq. 60 are solved numerically.
In Fig. 4, we illustrate the concentration profiles of oxygen vacancies and dopant, as well as the charge density and electrostatic potential as functions of $x$. From Fig. 4(a) to Fig. 4(e), we progressively increase the velocity of the core from 0 $\text{nm}\,\text{s}^{-1}$ to 100 $\text{nm}\,\text{s}^{-1}$. The direction of the core motion is from left to right along $x$-axis. In the following context, the left side of the core is named as the growing grain, while the right side is called the shrinking grain.
In Fig. 4(a), the symmetric SCL formation is presented at $v_{c}=0$ $\text{nm}\,\text{s}^{-1}$. However, the symmetric pattern begins to deteriorate even at very low velocity, e.g. 0.01 $\text{nm}\,\text{s}^{-1}$, as shown in Fig. 4(b1), (b2) and (b3). Compared with the concentration profiles at an equilibrium state, the accumulation zone of dopant in the growing grain is reduced, while that of the shrinking grain increases. This asymmetry arises, since during the core motion, dopant cations in the shrinking grain are captured by core due to their low diffusivity. Driving the motion of extra dopant cations within the core demands extra free energy dissipation. Therefore, the dopant cations flowing from the core region into the growing grain lead to the enhancement of the accumulation zone in the growing grain. However, the concentration of dopant in the core decreases from $7.8\times 10^{25}$ $\text{m}^{-3}$ ($v_{c}=0$ in Fig. 4(a1)) to $6.8\times 10^{28}$ $\text{m}^{-3}$ at $v_{c}=0.01$ $\text{nm}\,\text{s}^{-1}$. In the growing grain far from the core, dopant concentration reaches $7.2\times 10^{25}$ $\text{m}^{-3}$, which is higher than equilibrium concentration, i.e. $c_{\text{dop},\text{b}}(\infty)$, as shown in Fig. 4(a1). In Fig. 4(b2), we plot the calculated charge density $\rho=e(z_{\text{dop}}c_{\text{dop}}+z_{\text{V}_{\text{O}}^{..}}c_{\text{V}_{\text{O}}^{..}})$. The charge density in the core decreases. Compared with the equilibrium case, we also observe the formation of less (more) negative charged zone in the shrinking (growing) grain, respectively. In Fig. 4(b3), the grain boundary potential increases from 0.248 V ($v_{c}=0$ $\text{nm}\,\text{s}^{-1}$) to 0.265 V at $v_{c}=0.01$ $\text{nm}\,\text{s}^{-1}$. Moreover, an electrostatic potential different of 0.011 V develops between the shrinking and growing grains.
When we increase the velocity of the core to 0.5 $\text{nm}\,\text{s}^{-1}$, dopant cations are not capable of catching the movement of core, as depicted in Fig. 4(c1). Consequently, the dopant concentration within the core decreases substantially, resulting in a smaller depletion zone ahead of the core and a reduced accumulation zone behind it. Similarly, the concentration of oxygen vacancies within the core also decreases significantly. Correspondingly, as shown in Fig. 4(c2), the charge density in the core drops to $8.6\times 10^{-7}$ $\text{C}\text{m}^{-3}$, which contrasts with the increase in charge density in the core observed in Virkant’s work [38] at intermediate velocities. We attribute the difference to the defect chemistry incorporated in the present phase field model.
When the core velocity increases to 2 $\text{nm}\,\text{s}^{-1}$, the segregation of dopant in the core becomes negligible, as shown in Fig. 4(d1). The concentration of dopant on both the left and right sides of the bulk phase returns to the equilibrium value, i.e. $c_{\text{V}_{\text{O}}^{..},\text{b}}(\infty)$. In Fig. 4(d2), due to the continuous decrease in charge density within the core, the negatively charged zone is diminished. Furthermore, the grain boundary potential increases to 0.446 V and the electrostatic potential difference between the left and right sides reaches 0.122 V.
At a high core velocity, such as 100 $\text{nm}\,\text{s}^{-1}$, the dopant concentration profile becomes completely flat, while the oxygen vacancy concentration profile exhibits strong asymmetry, as depicted in Fig. 4(e1). The concentration of oxygen vacancies within the core decreases significantly to $2.46\times 10^{26}$ $\text{m}^{-3}$. In the shrinking grain, the depletion zone of oxygen vacancies remains sharp, about 2.5 nm, but it extends to approximately 20 nm in the growing grain. Consequently, as shown in Fig. 4(e2), the negatively charged zone in the shrinking grain is much smaller, and the negative charge region in the growing grain is negligible. In Fig. 4(e3), the electrostatic potential is illustrated. The grain boundary potential increases to 1.02 V and decays very slowly in the growing grain. In addition, the electrostatic potential difference becomes more pronounced, reaching 0.995 V.
In summary, at small core velocity, the asymmetric formation of SCL is mainly caused by dopant coming apart from the core. The charge density in the negatively charged SCL and in the positively charged core both decrease. As a result, the electrostatic potential in the shrinking grain increases to a positive value. At high core velocity, dopant cations do not segregate to the core. The oxygen vacancy concentration profile is strongly asymmetric. This asymmetry causes an extended diffusion layer in the growing grain and elevates the electrostatic potentials in both the core and the shrinking grain.
Figure 4: Asymmetric SCL formation in a STO bicrystal system with $c_{\text{dop},\text{b}}(\infty)=7.0\times 10^{25}$ $\text{m}^{-3}$, $\Delta g_{\text{V}_{\text{O}}^{..}}=-1.5$ eV and $\Delta g_{\text{dop}}=-0.5$ eV at quasi-equilibrium state under different constant core velocity (a) $v_{\text{c}}=0$ $\text{nm}\,\text{s}^{-1}$, (b) $v_{\text{c}}=0.01$ $\text{nm}\,\text{s}^{-1}$, (c) $v_{\text{c}}=0.5$ $\text{nm}\,\text{s}^{-1}$, (d) $v_{\text{c}}=2$ $\text{nm}\,\text{s}^{-1}$, (e) $v_{\text{c}}=100$ $\text{nm}\,\text{s}^{-1}$. Each row has three sub-figures, depicting the distribution of dopant and oxygen vacancy concentrations, total charge density and electrostatic potential, respectively.
In order to further investigate solute drag effects and identify the critical core velocity at which a velocity jump occurs during sintering process, we present the core velocity as a function of the total driving force calculated via Eq. 60 in Fig. 5. The segregation energy difference, $\Delta g_{\text{dop}}$ and $\Delta g_{\text{V}_{\text{O}}^{..}}$, allow direct tuning of the concentrations of oxygen vacancy and dopant in the core, thereby impacting the critical total driving force required for a specific core velocity. In Fig. 5(a), we hold $\Delta g_{\text{dop}}$ and $c_{\text{dop},\text{b}}(\infty)$ constant at -0.5 eV and $7\times 10^{25}$ $\text{m}^{-3}$ and progressively decrease $\Delta g_{\text{V}_{\text{O}}^{..}}$ from -0.5 eV to -1.75 eV. During the grain shrinking process, the driving force resulting from the GB curvature becomes increasingly significant. Consequently, the velocity of the GB core evolves from a low value to a high value. At low core velocities, a more negative $\Delta g_{\text{V}_{\text{O}}^{..}}$ induces higher concentration of oxygen vacancy in the core consequently demanding a larger driving force to maintain the same velocity. As the core velocity increases to 1 $\text{nm}\,\text{s}^{-1}$, the dopant cations are unable to keep pace with the core movement, resulting in a nearly flat dopant concentration within the core. Hence, this leads to the same linear relationship between the total driving force and the core velocity for different $\Delta g_{\text{V}_{\text{O}}^{..}}$ at high velocity. On the other hand, for fixed $\Delta g_{\text{V}_{\text{O}}^{..}}$, such as $\Delta g_{\text{V}_{\text{O}}^{..}}$=-1.75 eV, we clearly observe the velocity jumps from 0.065 $\text{nm}\,\text{s}^{-1}$ to approximate 1.5 $\text{nm}\,\text{s}^{-1}$ at $1.45\times 10^{7}$ $\text{N}\text{m}^{-2}$. As we increase $\Delta g_{\text{V}_{\text{O}}^{..}}$ to -1.0 eV, we hardly observe the velocity jump. When $\Delta g_{\text{V}_{\text{O}}^{..}}$ = -0.5 eV, both the concentrations of dopant and oxygen vacancy in the core decrease significantly. The solute drag effect therefore becomes negligible. Fig. 5(b) shows the influence of $\Delta g_{\text{dop}}$ on solute drag effects when $\Delta g_{\text{V}_{\text{O}}^{..}}$ and $c_{\text{dop},\text{b}}$ are held constant at -1.5 eV and $7\times 10^{25}$ $\text{m}^{-3}$. In contrast to Fig. 5(a), we can evidently observe velocity jumps for all $\Delta g_{\text{dop}}$. As we decrease $\Delta g_{\text{dop}}$ to less negative value, such as from -0.2 eV to -0.3 eV, the enhancement of the solute drag effect is not significant. In Fig. 5(c), we explore the impact of the dopant concentration in the bulk far from the core, denoted as $c_{\text{dop},\text{b}}(\infty)$. We fix $\Delta g_{\text{V}_{\text{O}}^{..}}$ and $\Delta g_{\text{dop}}$ at -1.5 eV and -0.5 eV. As $c_{\text{dop},\text{b}}(\infty)$ increases, a larger driving force is necessary to counterbalance the solute drag effect to achieve the same core velocity. Additionally, a lower critical velocity is observed when the velocity jump occurs at smaller values of $c_{\text{dop},\text{b}}(\infty)$. Specifically, the critical velocity increases from 0.065 $\text{nm}\,\text{s}^{-1}$ to 0.095 $\text{nm}\,\text{s}^{-1}$, as $c_{\text{dop},\text{b}}(\infty)$ is raised from $4\times 10^{25}$ $\text{m}^{-3}$ to $2\times 10^{26}$ $\text{m}^{-3}$.
In the sintering process, the driving forces are initially large and decrease as the grain size increases. In other words, the GB core velocity starts on the upper right side of the S-shaped curves in Fig. 5 at the initial stage of the grain growth. As the grain size increases, the GB core velocity decreases to the critical velocity. At this point, a transition from high velocity to low velocity occurs at the critical value. In Fig. 5(a), we observe GB core velocity jumps when $\Delta g_{\text{V}_{\text{O}}^{..}}=$1.5 eV and -1.75 eV. The critical velocities for different $\Delta g_{\text{V}_{\text{O}}^{..}}$ are approximately 0.35 $\text{nm}\,\text{s}^{-1}$. Additionally, the GB core velocity jumps to a lower value when $\Delta g_{\text{V}_{\text{O}}^{..}}$ is more negative. In Fig. 5(b), the critical GB core velocity increases from 0.15 $\text{nm}\,\text{s}^{-1}$ to 0.65 $\text{nm}\,\text{s}^{-1}$ as $\Delta g_{\text{dop}}$ decreases from -0.3 eV to -0.7 eV. In Fig. 5(c), a larger $c_{\text{dop},\text{b}}(\infty)$ leads to a higher critical velocity.
Figure 5: The total driving force on a GB with a constant velocity with (a) $\Delta g_{\text{dop}}=-0.5$ eV, $c_{\text{dop},\text{b}}(\infty)=7\times 10^{25}$ $\text{m}^{-3}$ and different $\Delta g_{\text{V}_{\text{O}}^{..}}$, (b) $\Delta g_{\text{V}_{\text{O}}^{..}}=-1.5$ eV, $c_{\text{dop},\text{b}}(\infty)=7\times 10^{25}$ $\text{m}^{-3}$ and different $\Delta g_{\text{dop}}$, (c) $\Delta g_{\text{V}_{\text{O}}^{..}}=-1.5$ eV, $\Delta g_{\text{dop}}=-0.5$ eV and different $c_{\text{dop},\text{b}}(\infty)$. The pentagram symbols in subfigure b correspond to the different GB core velocities studied in Fig. 4. The green, blue and saffron pentagram symbols indicate $v_{\text{c}}$=2 $\text{nm}\,\text{s}^{-1}$, 0.5 $\text{nm}\,\text{s}^{-1}$ and 0.01 $\text{nm}\,\text{s}^{-1}$, respectively. Due to the range limit of y-axis, $v_{\text{c}}$=0 $\text{nm}\,\text{s}^{-1}$ (Fig. 4 a) and 100 $\text{nm}\,\text{s}^{-1}$ (Fig. 4 e) are not marked.
5.3 Abnormal grain growth
After systematically benchmarks and knowledge gained from quasi-static single GB cases, we investigate different abnormal grain growth patterns of materials with different defect segregation energies $\Delta g_{\text{dop}}$, $\Delta g_{\text{V}_{\text{O}}^{..}}$ and bulk defect concentrations such as $c_{\text{dop},\text{b}}$. The segregation of point defects in the core plays a significant role in influencing grain growth kinetics. To investigate the grain size distribution of STO polycrystalline material after sintering, we conduct 2D phase-field simulations. Note that three-dimensional phase-field simulations of the proposed model are achievable but require significantly more computational resources, which will be considered in the near future. In fact, to reveal the fundamental mechanisms, 2D simulations should be sufficient and allow for a comprehensive parameter study given the limited computational resources.
The simulation box size is set to 390 nm $\times$ 390 nm with a grid resolution of 1000 $\times$ 1000 points. The polycrystalline microstructure is initialized using a Voronoi tessellation method, where each grain was assigned a unique grain index. We started with an initial number of 1000 grains, and the random seed used for generating the polycrystalline microstructure is set to 42. To optimize memory usage during simulations, we also utilize grain tracker algorithm [34]. In the simulations, we set $N_{\text{V}_{\text{O}}^{..},\text{c}}$, $N_{\text{V}_{\text{O}}^{..},\text{b}}$, $N_{\text{dop},\text{c}}$ and $N_{\text{dop},\text{b}}$ as $1\times 10^{27}$ $\text{m}^{-3}$, $5.12\times 10^{28}\text{m}^{-3}$, $1.68\times 10^{27}$ $\text{m}^{-3}$ and $1.69\times 10^{28}$ $\text{m}^{-3}$. The simulations are conducted at a temperature of 1623 K, where the non-Arrhenius behavior of the grain growth is found [28]. Periodic boundary conditions are applied to all edges of the simulation domain for the phase-fields and concentration fields. For the electrostatic potential field, the bottom-left corner is grounded, while the other edges are set with the periodic boundary conditions. Furthermore, all the parameters used in the simulation are normalized. The lattice constant $a$ = $3.9\times 10^{-10}$ m is used for the length normalization. Other normalization factors include e.g. $f_{0}=\kappa/a^{2}$=$2.3\times 10^{9}$ $\text{J}\text{m}^{-3}$ for the energy density, $t_{0}=1/(Lf_{0})=0.0043$ s for the interface mobility, and $\phi_{0}=f_{0}V_{\text{m}}/\mathcal{F}$= 0.858 V for the electrostatic potential.
We present first the microstructure evolution of the grain growth process under specific conditions $\Delta g_{\text{V}_{\text{O}}^{..}}$ = -1.5 eV, $\Delta g_{\text{dop}}$=-0.4 eV and $c_{\text{dop},\text{b}}=7\times 10^{25}$ $\text{m}^{-3}$ when the sintering time is $t/\text{h}=\{0.067,0.33,0.67,1.0\}$. The average area of grain at time $t=0$ is <S>, i.e., 152.1 $\text{nm}^{2}$. When $t$=0.067 h, 0.33 h and 0.67 h, the grain size increases while the area fraction of small grain decreases. The grain size distribution starts to transition from a mono-modal to a bimodal pattern. When $t$=1 h, the largest grain area is about 32 times larger than <S>. The bimodal grain size distribution is clearly visible.
Figure 6: The microstructure evolution of grain growth process at (a) 0.067 h, (b) 0.33 h, (c) 0.67 h and (d) 1 h, with defect chemistry parameters set at $\Delta g_{\text{V}_{\text{O}}^{..}}$ = -1.5 eV, $\Delta g_{\text{dop}}$=-0.4 eV and $c_{\text{dop},\text{b}}=7\times 10^{25}$ $\text{m}^{-3}$. These parameters are consistent with those used in Fig. 7(e). By 1h, a bimodal grain size distribution becomes apparent.
5.3.1 Influence of segregation energies and bulk defect concentration
Subsequently, we investigate the influence of defect chemistry parameters on the microstructure evolution during sintering processes. The microstructure of grains is presented in Fig. 7(a)-(g) after 1 hour of sintering at 1623 K. In Fig. 7(a), (b) and (c), we increase $\Delta g_{\text{V}_{\text{O}}^{..}}$ from -1.5 eV to -0.5 eV when $\Delta g_{\text{dop}}$=-0.5 eV and $c_{\text{dop},\text{b}}=7\times 10^{25}$ $\text{m}^{-3}$. Bimodal grain size distribution is clearly observed in Fig. 7(a2), (b2) and (c2). As $\Delta g_{\text{V}_{\text{O}}^{..}}$ becomes more negative, the peak in the range of $40<\frac{S}{<S>}<$ 4 induced by smaller grains becomes more pronounced. As predicted in Fig. 5(a), solute drag effect is almost negligible for $\Delta g_{\text{V}_{\text{O}}^{..}}$=-0.5 eV. Nonetheless, the formation of solute cloud imprinted by vanished GBs is still detectable in Fig. 7(c3) even though the solute drag effect is very weak. Concentration of oxygen vacancy is detailed in Fig. 7(a4), (b4) and (c4). Segregation is more significant for $\Delta g_{\text{V}_{\text{O}}^{..}}=-1.5$ eV. Compared to the dopant, solute cloud of oxygen vacancies are less distinct. In view of the higher diffusivity of oxygen vacancy, the contribution of oxygen vacancy to solute drag force is also weak. Solute clouds of oxygen vacancies are only observed near the junctions where grains shrink rapidly with a small radius. The electrostatic potential distributions are depicted in Fig. 7(a5), (b5) and (c5). The negative potential bands can be also observed near the GB cores. As all the considered cases show, the electrostatic potentials in small shrinking grains are significantly higher than in the surrounding growing grains, leading to a stronger electrostatic contribution to the solute drag effects. Consequently, the shrinking speed of small grains is reduced, which induces pinning effects on the surrounding growing grains. Note that the heterogeneous electrostatic potential distribution arises from the velocity difference of individual GB, in accordance to the observation in the quasi-static cases in Fig. 4, rather than any anisotropy which is absent in the current simulations.
In Fig. 7(a), (d) and (e), we keep $\Delta g_{\text{V}_{\text{O}}^{..}}$ and $c_{\text{dop},\text{b}}$ at -1.5 eV and $7\times 10^{25}$ $\text{m}^{-3}$ and test different $\Delta g_{\text{dop}}$ values. Whereas, in Fig. 7(a), (f) and (g), $\Delta g_{\text{V}_{\text{O}}^{..}}$ and $\Delta g_{\text{dop}}$ are constant while $c_{\text{dop},\text{b}}$ increases from $2\times 10^{25}$ $\text{m}^{-3}$ to $2\times 10^{26}$ $\text{m}^{-3}$. For all simulations, the solute drag effects from dopant are pronounced. The dopant clouds imprinted by vanished GBs are clearly observed behind GB cores. In Fig. 7(e2) and (g2), the bimodal grain size distributions are presented. However, when $\Delta g_{\text{dop}}=-0.6$ eV in Fig. 7(d2) and $c_{\text{dop},\text{b}}=2\times 10^{26}$ $\text{m}^{-3}$ in Fig. 7(f2), the unimodal grain size distribution exhibits, due to the more pronounced probability of small grain. The simulation results in Fig. 7(f) demonstrate the high grain number, where the large grains are surrounded by small grains.
Figure 7: Phase-field simulation results of abnormal grain growth in Fe-doped STO after 1 h sintering under the influence of different defect chemistry parameters.
5.3.2 Grain boundary potential distribution
In Fig. 8, we present the statistics of the average grain boundary potential in the GB core, denoted as $\bar{\phi}_{\text{core}}$. The value of $\bar{\phi}_{\text{core}}$ is calculated as an average of the electrostatic potential along the GB core indicated by $\eta_{i}=0.5$ over individal GB length. Fig. 8(a) illustrates the distribution of $\bar{\phi}_{\text{core}}$ for $\Delta g_{\text{V}_{\text{O}}^{..}}=-0.5$ eV, $\Delta g_{\text{dop}}=-0.5$ eV, $c_{\text{dop}}(\infty)=7\times 10^{25}$ $\text{m}^{-3}$ (refer to Fig. 7(c) for the associated phase-field simulation result). When the solute drag effect is negligible, $\bar{\phi}_{\text{core}}$ is smaller than 0.05 V and approximately follows the normal distribution. In Fig. 8(b), the distribution of $\bar{\phi}_{\text{core}}$ is presented for $\Delta g_{\text{V}_{\text{O}}^{..}}=-1.5$ eV, $\Delta g_{\text{dop}}=-0.5$ eV, $c_{\text{dop}}(\infty)=7\times 10^{25}$ $\text{m}^{-3}$ (refer to Fig. 7(a)). Here, $\bar{\phi}_{\text{core}}$ significantly increases. In addition, the distribution of $\bar{\phi}_{\text{core}}$ changes from mono-modal to bimodal, with one peak at 0.31 V and another at 0.35 V. In Fig. 8(c), the parameters $\Delta g_{\text{V}_{\text{O}}^{..}}=-1.5$ eV, $\Delta g_{\text{dop}}=-0.5$ eV, $c_{\text{dop}}(\infty)=2\times 10^{26}$ $\text{m}^{-3}$ are chosen (refer to Fig. 7(f)). In this case, the solute drag effect is the most pronounced in this study. Not only $\bar{\phi}_{\text{core}}$ increases but also the bimodal distribution of $\bar{\phi}_{\text{core}}$ becomes more distinct.
[…]
In addition, two distinct types of GBs can be observed Fig. 9. The first type, characterized by a lower moving velocity, exhibits nearly symmetric concentration profiles across the GB core with a significantly higher segregation concentration of dopant. However, within a polycrystalline system, the accumulation zone of dopant at the space charge region transitions into a depletion zone. Conversely, the second type of GB core, which moves at a higher velocity, shows lower dopant concentrations due to the break of the symmetric SCL. Similarly, the oxygen vacancy distribution in this second GB core is also asymmetric, as illustrated in Fig. 7(e). Fig. 7(f) presents the electrostatic potential distributions across these two types of GB cores, where the second GB core displays a larger grain boundary potential and a more pronounced electrostatic potential difference.
Figure 9: Two types of GBs are observed in the rectangular region highlighted in Fig. 7(f). The distributions of dopant concentration, oxygen vacancy concentration, and electrostatic potential in this region are shown in (a), (b), and (c), respectively. Particularly, the dopant concentration profile within the grain interior in (a) exhibits strong imprints of vanished GBs during sintering. Two types of GB cores are distinguished by dashed yellow lines in (a). The first type, characterized by a low moving velocity, has a higher segregation concentration of dopants. In contrast, the second type, with a higher core velocity, shows a lower and asymmetric dopant concentration across the GB core. Furthermore, the dopant concentration, oxygen vacancy concentration, and electrostatic potential along the two types of GB cores are plotted in (d), (e), and (f), respectively.
In order to further evidence the two types of GB core, 14 different GBs are selected from the rectangular region in Fig. 7 (f). The relations between $v_{\text{c}}$ and $F_{\text{T}}$ as well as $\bar{\phi}_{\text{core}}$ and $F_{\text{T}}$ for each GB core are represented in Fig. 10(b). Here, $F_{\text{T}}$ is the total driving force, numerically calculated via Eq. 60. The solid lines indicate the phase-field simulation results in the bicrystal case (refer to the purple line in Fig. 5(c)), while the symbols represent different GBs in the polycrystalline case. In Fig. 10(b), the relationships of $F_{\text{T}}$ vs. $v_{\text{c}}$ and $F_{\text{T}}$ vs. $\bar{\phi}_{\text{core}}$ in the polycrystalline case almost follow those in the bicrystal case. There is also a velocity jump and an electrostatic potential jump when the total driving force is approximately $2\times 10^{7}$ $\text{N}/\text{m}^{2}$. In the high GB velocity region, GB6, GB7 and GB8, with less dopant segregation, have much higher electrostatic potentials than the others, aligning well with the predictions in the bicrystal case. In the low GB velocity region, GB3, GB13 and GB14 also agree well with the results from the bicrystal quasi-static case. Whereas, GBs with higher velocities, such as GB2, GB4, GB11 and GB12, do not fit well into the S curve of a bicrystal case. This discrepancy can be attributed to multiple reasons. For instance, these GBs have higher GB curvatures and can interact with other GBs, which go beyond the bicystal model and can only be resolved by the grain growth simulations presented here.
Figure 10: GB core velocity $v_{\text{c}}$ and average grain boundary potential $\bar{\phi}_{\text{core}}$ as functions of the total driving force $F_{\text{T}}$. The 14 different GBs, shown in Fig. 9(a), are chosen from the rectangular region in Fig. 7(f) in polycrystalline simulations. The relations between $v_{\text{c}}$ and $F_{\text{T}}$ as well as $\bar{\phi}_{\text{core}}$ and $F_{\text{T}}$ are obtained from phase-field simulations in bicrystal and polycrystalline Fe-doped STO. Solid lines indicate the simulation results for the bicrystal case, while dashed lines are the extrapolations of the simulation results in bicystal case at very low GB velocity regime. Other symbols represent 14 different GBs chosen in the polycrystalline case.
6 Conclusion
In this work, we develop a phase-field model to investigate SCL formation during grain growth in oxide electroceramics based firmly on principles of the defect chemistry. The main results of the paper include:
  • 1.
    The proposed phase-field model reproduces a series of benchmark results at equilibrium of Fe-doped STO bicrystal. Simulation results, including the symmetric SCLs, grain boundary potentials, concentrations of oxygen vacancies and acceptor dopants in the GB core and the bulk phase as well as the electrostatic distributions, show good agreement with the analytical predictions in the MS and GC models.
  • 2.
    Simulations are carried out comprehensively for a GB core moving with constant velocities under quasi-equilibrium state. At low core velocity, the accumulation zone in the shrinking grain decreases, while it increases in the growing grain, leading to the asymmetric SCL formation. With increasing GB core velocity, the dopant segregation diminishes. However, the movement of oxygen vacancy within the GB core only breaks at extremely high velocity due to the high diffusivity of oxygen vacancy. In addition, the electrostatic potential in the shrinking grain becomes higher than that in the growing grain.
  • 3.
    Abnormal grain growth is observed even without distinguishing any grain orientation or considering any anisotropic GB mobility in the phase-field simulations at 1623K. It indicates that the solute drag effect alone can lead to bimodal grain size distribution. As the formation energy differences of oxygen vacancies and acceptor dopants ($\Delta g_{\text{V}_{\text{O}}^{..}}$ and $\Delta g_{\text{dop}}$) become more negative, the grain growth velocity decreases. This trend is also observed with increasing dopant concentration ($c_{\text{dop},\text{b}}$). Notably, the dopant concentration profile within the grain interior distinctly reflects the significant imprint of the grain boundaries that disappeared during sintering.
  • 4.
    Simulation results reveal two distinct types of GB cores. The first type, with nearly symmetric concentration profiles at the GB, exhibits stronger solute drag effects, resulting in a much lower moving velocity and a tendency to remain pinned in position. The second type of GB core shows asymmetric concentration profiles and a higher moving velocity, contributing predominantly to grain coarsening.
  • 5.
    Abnormal grain growth resulting from solute drag effects leads to a large variation of grain boundary potentials. Typically, smaller grains exhibit higher grain boundary potentials at the later stages of grain growth and form more blocking GBs in term of ionic conductivity. When considering the total conductivity of an electrolyte material, the electrical current needs to find a ”detour” around these blocking GBs by passing through the surrounding larger grains, with less blocking GBs. It implies that to increase the conductivity one may also optimize the microstructure so that blocking GBs can be circumvented more easily. In contrast, if blocking GBs are desired, such as in capacitors, the grain size should be kept small to prevent the development of detour paths.
The proposed phase-field model offers naturally a series of promising extensions. We can e.g. incorporate additional species, such as neutral dopants, electrons, and electron holes, into the current phase-field model to quantitatively investigate the formation of the SCL. Furthermore, GB orientation plays a crucial role in SCL formation, and both the interface energy and the defect chemistry parameters $\Delta g_{\text{V}_{\text{O}}^{..}}$ and $\omega_{c}N_{\text{V}_{\text{O}}^{..},\text{c}}$ depend on GB orientation [49]. These complexities can be included in the phase-field model in the near future. Moreover, we assume in this paper only the $\text{Fe}^{3+}$ and ignore $\text{Fe}^{4+}$, which is reasonable for high temperature scenarios. But, the situation regarding solute drag effect changes at low temperature when a significant amount of $\text{Fe}^{4+}$ is present. The $\text{Fe}^{3+}$ ions in the GB core can not easily follow the movement of the GB core due to their low diffusivity. In contrast, the rapid movement of electron holes can adjust the local $\text{Fe}^{3+}$/$\text{Fe}^{4+}$ ratio and establish an equilibrium charge distribution. The solute drag effect may differ when a significant amount of $\text{Fe}^{4+}$ is considered.
Acknowledgement
The financial support of German Science Foundation (DFG) in the framework of the Collaborative Research Centre 1548 (CRC 1548, project number 463184206) and the Project 471260201 are acknowledged. The authors K. Wang and B.-X. Xu greatly appreciate the access to the Lichtenberg II High-Performance Computer (HPC) and the technique supports from the HHLR, Technische Universität Darmstadt. The computing time on the HPC is granted by the NHR4CES Resource Allocation Board under the project “special00007”.

### v3

@@
-A defect-chemistry-informed phase-field model of grain growth in oxide electroceramics
+A defect-chemistry-informed phase-field model of grain growth in oxide ceramics: application to Fe-doped SrTiO3
 Abstract
-Dopants can significantly affect the properties of oxide ceramics through their impact on the property-determined microstructure characteristics such as grain boundary segregation, space charge layer formation in the grain boundary vicinity, and the resultant microstructure features like bimodality due to abnormal grain growth.
-To support rational oxide ceramics design, we propose a multiphysics-based and defect-chemistry-informed phase-field grain growth model to simulate the microstructure evolution of oxide ceramics.
-It fully respects the thermodynamics of charged point defects (oxygen vacancies and dopants) in both the grain interior and boundaries and considers the competing kinetics of defect diffusion and grain boundary movement.
-The proposed phase-field model is benchmarked against well-known simplified bicrystal models, including the Mott-Schottky and Gouy-Chapman models.
-Various simulation results are presented to reveal the impacts of defect formation energy differences between the grain interior and the grain boundary core on the key microstructural aspects.
-In particular, simulation results confirm that the solute drag effect alone can lead to bimodal grain size distribution, without any contribution from grain misorientation and other anisotropy.
-Interestingly, abnormal grain growth simulations demonstrate that grain boundary potentials can vary substantially: grain boundaries of larger grains tend to have lower potentials than those of smaller grains.
+Dopants can significantly affect the properties of oxide ceramics through their impact on the property-determined microstructure characteristics such as grain boundary (GB) segregation, space charge layer formation in the GB vicinity, and the grain growth deviating from normal patterns.
+To support the rational design of oxide ceramics, we propose a defect-chemistry-informed phase-field grain growth model to simulate the microstructure evolution of oxide ceramics.
+It fully respects the defect-chemistry theory by accounting for the distinct segregation energies and available site densities of charged point defects (oxygen vacancies and acceptor dopants) in both the grain interior and boundaries, and it considers the competing kinetics of defect diffusion and GB movement.
+The proposed phase-field model is benchmarked against well-known bicrystal models, including the Mott-Schottky and Gouy-Chapman models.
+Various simulation results are presented to reveal the effect of different defect-chemistry parameters on the space charge layer formation and key microstructural aspects.
+In particular, simulation results confirm that the solute drag effect alone can lead to skewed grain size distribution that do not follow the log-normal distribution, without any contribution from grain misorientation and other anisotropy.
+Interestingly, simulations also demonstrate that grain boundary potentials can vary substantially: grain boundaries of larger grains tend to have lower potentials than those of smaller grains.
 Such heterogeneous grain boundary potential distribution may inspire a new material optimization strategy through microstructure design.
-This study provides a comprehensive framework for defect-chemistry-consistent investigations of microstructure evolution in polycrystalline oxide ceramics, offering fundamental insights into in-situ processes during critical manufacturing stages.
+This study provides a comprehensive framework for defect-chemistry-consistent investigations of microstructure evolution in polycrystalline oxide ceramics, offering fundamental insights into microscopic processes during critical manufacturing stages.
 keywords:
-Phase-field method , Defect-chemistry , Space charge layer , Abnormal grain growth
+Phase-field method , Defect-chemistry , Space charge layer , Grain growth
 {graphicalabstract}[Uncaptioned image]
 1 Introduction
-In recent decades, oxide ceramics, such as SrTiO3 (STO) and BaTiO3 (BTO), have undergone tremendous development and are increasingly crucial in applications, such as solid oxide fuel cells [1, 2], sensors [3, 4], capacitors [5, 6], actuators [7, 8], and memristors [9, 10].
-Oxide ceramics can be fabricated via sintering or deposition processes [11, 12, 13, 14].
-When manufactured through sintering, oxide ceramics exhibit a polycrystalline microstructure, with their macroscopic properties heavily influenced by the synthesis conditions [15].
-To tailor the physical and chemical properties of oxide ceramics, heterovalent substitution (doping) is an effective method for introducing additional charged defects through ionic compensation during synthesis.
-The state-of-the-art designs are mostly empirical, since knowledge of the correlation among doping strategies, processing and the resultant microstructure is limited.
+In recent decades, oxide ceramics, such as SrTiO3 (STO) and BaTiO3 (BTO), have undergone tremendous development and are increasingly crucial in applications, such as capacitors lee2011atomic , hou2017ultrahigh , actuators kursumovic2013new , baek2013epitaxial , sensors wang2022stretchable , chan2014highly , memristors muenstermann2010coexistence , hu2013ferroelectric and electrolytes in fuel cells ishihara1994doped , kreuer2003proton , li2014family , Oxide ceramics can be fabricated via sintering or deposition mishra2022ultra , reavley2022ultrafast , sayer1990ceramic , li2003two .
+When manufactured through sintering, oxide ceramics exhibit a polycrystalline microstructure, with their macroscopic properties heavily influenced by the synthesis conditions chaim2008sintering .
+To tailor the physical and chemical properties of oxide ceramics, heterovalent substitution (doping) is an effective method for introducing additional charged defects through ionic or electronic compensation during synthesis.
+The state-of-the-art designs are mostly empirical, since knowledge of the relationships between doping strategies, processing and the resultant microstructure is limited.
 The influential mechanisms of doping on microstructure in oxide ceramics are manifold.
-First of all, the addition of doping elements in the parent material systems triggers defect chemistry reaction and results in multiple types of charged point defects or defect complexes, which interact with the grain boundary (GB) chemically, electrically and elastically and form space charge layers (SCLs) around GB.
-More specifically, the equilibrium concentrations of the point defects within the bulk are thermodynamically determined through defect chemistry [16, 17].
+First of all, the addition of doping elements in the parent material systems triggers defect chemistry reactions and results in multiple types of charged point defects or defect complexes, which interact with the grain boundary (GB) chemically, electrically and elastically and form space charge layers (SCLs) around GB.
+More specifically, the equilibrium concentrations of the point defects within the bulk are thermodynamically determined through defect chemistry maier2023physical , smyth2000defect .
 The scenario near inside the GB core and in the GB vicinity are different.
-In undoped and acceptor doped perovskites such as Fe-doped STO, the GB core is positively charged with excess oxygen vacancies, notated $\text{V}_{\text{O}}^{\text{..}}$ in Kröger-Vink notation [18], stemming from the negative formation energy difference of oxygen vacancy between GB core and bulk, i.e.
+In undoped and acceptor doped perovskites such as Fe-doped STO, the GB core is positively charged with excess oxygen vacancies, $\text{V}_{\text{O}}^{\text{..}}$ in Kröger-Vink notation kroger1956relations , stemming from the negative formation energy difference of oxygen vacancy between GB core and bulk, i.e.
 $\Delta g_{\text{V}_{\text{O}}^{..}}=\mu^{0}_{\text{V}_{\text{O}}^{..},\text{c}}-\mu^{0}_{\text{V}_{\text{O}}^{..},\text{b}}<0$.
-To compensate the positively charged GB core, a negatively charged region forms adjacent to it, called the space charge layer (SCL).
+To compensate the positively charged GB core, negatively charged regions form adjacent to it, called the space charge layers (SCLs).
 Sharp interface models have been used to describe the SCL and the related grain boundary potential.
-Among them are the Mott-Schottky (MS) model and the Gouy-Chapman (GC) model.
+Among them are the Mott-Schottky (MS) model mott1939theory , mott1938note , schottky1939halbleitertheorie and the Gouy-Chapman (GC) model guoy1910constitution , chapman1913li .
 The MS model assumes that the dopant concentration is constant, and only the oxygen vacancies are able to redistribute, in order to form the SCLs.
 In contrast, the GC model has both defects as mobile and thus allows both to contribute to charge compensation.
-Unlike oxygen vacancies, acceptor dopant forms an accumulation region within the SCL and in the GB core due to their negative charge [Fig.
+While positively charged oxygen vacancies are depleted in the case of a positive core, acceptor dopants are depleted in the case of a positive core within the SCL and in the GB core due to their negative charge [Fig.
 1(e)].
 The spatial inhomogeneity introduced by SCL formation exerts a profound impact on the transport dynamics of charge species.
-Evidence shows that cation diffusion parallel to the GB is expedited [19, 20], whereas charge transport across the GB is significantly hindered [21].
-SCLs determine not only the GB local properties of the synthesized material, but also impact the grain growth dynamics already during the synthesis, e.g., the well-known solute drag effect.
+Evidence shows that cation diffusion parallel to the GB is expedited parras2020grain , savzinas2017134ba , whereas charge transport across the GB is significantly hindered waser2000grain .
+SCLs determine not only the GB local properties of the synthesized material, but also impact the grain growth dynamics already during the synthesis through the well-known solute drag effect cahn1962impurity .
 Grain growth in oxide ceramics, driven by a reduction in surface area and controlled by the diffusion of charged species, plays a crucial role in determining the final grain size distribution and the concentration distributions of various point defects.
-Segregation in a moving GB core necessitates additional free energy dissipation and results in the solute drag effect [22, 23].
-This effect significantly interferes with the kinetics of grain growth and can independently trigger abnormal grain growth even in the absence of any texture or pinning sites [24].
-Dopant segregation to GB core has been experimentally evidenced, especially at high temperatures [25, 26].
-Experiments also show that even 0.2% amount of Fe in STO can pin certain portion of GBs [Fig.
+Segregation in a moving GB core necessitates additional free energy dissipation and results in the solute drag effect cahn1962impurity , hillert1976treatment .
+This effect strongly affects the kinetics of grain growth and can independently trigger abnormal grain growth even in the absence of any texture or pinning sites kim2008grain .
+Dopant segregation to GB core has been experimentally evidenced, especially at high temperatures zahler2023grain , jennings2024grain .
+Experiments also show that even 0.2% amount of Fe in STO can pin a certain portion of GBs [Fig.
 1(b)].
-This trend increases with 2% Fe [Fig.
+This trend increases at 2% Fe [Fig.
 1(c)].
 When increasing the dopant concentration to 5% [Fig.
-1(d)], the grain growth is completely suppressed [25, 27].
-In addition, the grain growth coefficient is found to follow a non-Arrhenius behavior between $1350^{\circ}$ and $1425^{\circ}$.
-Two types of GB core with different mobilities are proposed to explain the non-Arrhenius behavior [28, 25].
-Despite various evidence supporting the impact of SCLs on GB properties and grain growth dynamics, experimental characterization of the SCL remains a challenge [29].
+1(d)], the grain growth is completely suppressed zahler2023grain , rheinheimer2016grain .
+In addition, the grain growth coefficient is found to follow a non-Arrhenius behavior between $1350^{\circ}$C and $1425^{\circ}$C.
+Two types of GB core with different mobilities have been proposed to explain the non-Arrhenius behavior rheinheimer2015non , zahler2023grain .
+Despite various pieces of evidence supporting the impact of SCLs on GB properties and grain growth dynamics, experimental characterization of the SCL remains a challenge fleig2002grain .
 The complexity arises from the inherent structural and compositional heterogeneities present in polycrystalline oxide ceramics.
@@
 It is hence desirable to employ theoretical models and numerical simulations to evaluate SCLs and abnormal grain growth.
-De Souza derived numerical solutions based on defect chemistry in acceptor-doped oxide systems at equilibrium state [30].
-The impacts of different charge species, the formation energy difference of oxygen vacancy and dopant between GB core and bulk phase, temperature and partial pressure of oxygen on grain boundary potential are comprehensively studied in a bicrystal with planar GB core.
-However, applying these numerical solutions to polycrystalline cases is impractical.
-Under this circumstance, the phase-field method with a diffuse interface emerges as a powerful tool to solve different moving boundary problems [31, 32].
-There are a series of phase-field sintering models, e.g.
-to consider the solute drag effect in metals [24] and ceramics [33] and for non-isothermal sintering [34, 35, 36, 37].
-But these models do not consider the charged defects, electrostatics or defect chemistry, which are important for oxide ceramics.
+Kliewer and Koehler derived the local density of charge carriers in the presence of electrostatic potentials from a global approach kliewer1965space .
+Through minimizing the free energy of the entire crystal, the equations for the densities of various defects are derived in the type of Boltzmann distribution.
+Blakely and Macdonald extended the global thermodynamic approach with the consideration of an exhaustible number of defect sites and derived Fermi-Dirac-like distribution functions of defects blakely1973space , mukhopadhyay1991ionic , macdonald1980interfacial .
+Jamnik employed a local thermodynamic formalism, in which the interface region was divided into a core and a space charge region with an assumption that the material constants, i.e.
+the standard chemical potentials and mobilities of defects, behave as step functions jamnik1995interfaces .
+De Souza extended the local thermodynamic approach to abrupt GB||SCL model based on defect chemistry in acceptor-doped oxide systems at equilibrium state de2009formation .
+The impacts of different charge species, the formation energy difference of oxygen vacancy and dopant between GB core and bulk phase, the number of available sites of oxygen vacancy and dopants, temperature and partial pressure of oxygen on grain boundary potential are comprehensively studied in a STO bicrystal with planar GB core de2009formation .
+Then, the abrupt GB||SCL model was implemented to reproduce the SCL in STO to investigate the impact of negative electrode on the defects distribution, GB potential and grain growth kinetics rheinheimer2019grain .
+In addition, this model was also employed to investigate the SCL formation in fluorite-type oxide parras2020grainJACS , tong2020analyzing , parras2020grain .
+However, applying the abrupt GB||SCL model to polycrystalline sintering scenarios is challenging.
+The complex microstructure, intricate GB distribution and GB kinetics affected by solute drag effects hinder the practical use of the sharp interface methods.
+Under this circumstance, the phase-field method with a diffuse interface emerges as a powerful tool to solve different moving boundary problems wang2020modeling , wang2021quantitative , hu2024data .
+Kobayashi et al.
+developed the Kobayashi-Warren-Carter (KWC) phase-field model to consider crystallographic orientation during grain growth kobayashi2000continuum .
+Moelans et al.
+proposed a homogeneous multi-well potential in the free energy density functional to obtain the quantitative relations between phase-field model parameters and GB properties moelans2008quantitative .
+Additionally, there are also a series of phase-field sintering models for non-isothermal sintering yang20193d , yang2020investigation , yang2022diffuse , oyedeji2023variational .
+Several researchers have developed phase-field models to capture solute drag effects during grain growth processes.
+The first phase-field simulation was performed by Fan et al.
+fan1999computer .
+Their model revealed that solute drag introduces a nonlinear relationship between GB velocity and driving force and influences the growth exponent in polycrystalline systems.
+Cha et al.
+proposed a phase-field model that treats the GB as a distinct phase and studied the impact of solute drag on migrating grain boundaries in a binary alloy system cha2002phase .
+The effect of solute drag is automatically incorporated into the model and it is capable of reproducing the equilibrium solute segregation and the free energy dissipation.
+Grönhagen et al.
+introduced a concentration dependency in the height of the double-well potential in the Gibbs-energy expression and verified the possibility of modeling solute drag on a moving boundary gronhagen2007grain .
+This model was then adopted by Kim kim2008grain and Li li2009phase .
+Kim et.
+al investigated the grain growth in association with GB segregation in two-dimensional polycrystalline systems kim2008grain .
+They concluded abnormal grain growth can be triggered by the solute drag effect.
+Li et al.
+observed that the solute concentration at the moving GB may increase with increasing velocity and becomes larger than the equilibrium value, which is unexpectedly predicted by solute drag theory li2009phase .
+They also found non-linear relationship between the curvature driving force and the GB velocity can exist not only in the transition regime of the velocity, but also in the low velocity regime.
+Greenwood et al.
+studied the atomic scale interaction of solute with a migrating GB by using a phase-field crystal model greenwood2012phase .
+They developed a new formalism to allow for the application of an external driving pressure for the growth of one grain.
+Extending from grain growth studies, phase-field models for solid-state sintering in the initial and final stage considering various diffusion paths are proposed by Wang et al.
+wang2006computer , Kumar et al.
+kumar2010phase , Hötzer et al.
+hotzer2019phase and Rehn et al.
+rehn2019phase .
+Rheinheimer et.
+al employed a multi-phase-field model to simulate the abnormal grain growth in STO ceramics by introducing two types of GB mobilities rheinheimer2019non .
+However, the electrostatic effects related SCL is not considered.
+In order to consider the electrostatic contribution in the phase-field model, Guyer et al.
+developed a phase-field model to capture the charge distribution in an equilibrium double layer at the electrochemical interface with mass and volume constraints, Poisson’s equation, ideal solution thermodynamics, and a simple description of the competing energies in the interface guyer2004equilibrium .
+By using the same model, Guyer et al.
+also explored its kinetic behavior for electrodeposition and electrodissolution conditions guyer2004kinetic .
+They demonstrated the ohmic conduction in the electrode and ionic conduction in the electrolyte and found the nonlinear relationship between current and overpotential as predicted by Butler-Volmer equation.
+Bishop et al.
+formulated a Helmholtz free energy functional with the electrostatic contribution and studied the phase separation of charged species by spinodal decomposition bishop2003effect .
+García et al.
+incorporated the laws of thermodynamics and Maxwell’s equations and proposed a theoretical framework including thermodynamically consistent equilibrium equations and kinetic driving forces to describe the time evolution for electrically and magnetically active materials garcia2004thermodynamically .
+Lund et al.
+proposed a generalized framework that integrates the free energy contributions of thermochemical, structural, mechanical, and electrical fields lund2021thermodynamically .
+They calculated the SCL in gadolinium-doped cerium oxide with high dopant concentration and predicted the macroscopic ionic conductivity with different segregation energy of oxygen vacancy.
+Aiming to investigate the sintering process with electrostatics, multi-physics phase-field model needs to be developed.
+Vikrant et al.
+developed a phase-field model including the electro-chemo-mechanical effects to physically describe the equilibrium and transport properties of charged interfaces in ion-conducting solids vikrant2018charged .
+They simulated gadolinium-doped cerium oxide and investigated the influence of gadolinia substitution concentration on the SCL.
+Then, Vikrant et al.
+extended this model to understand the effects of the intrinsic and extrinsic ionic species and point defects on the structural and electrochemical stability of grain boundaries in polycrystalline ceramics vikrant2019charged .
+They studied the impacts of different crystallographic misorientations on SCL in the cubic Yttria Stabilized Zirconia.
 Recently, Vikrant et al.
-developed a multi-physics phase-field model with an emphasis on GB misorientation and investigate the SCL formation at equilibrium and quasi-equilibrium state [38].
-Then, this model is applied to explain the abnormal grain growth with a bimodal pattern as observed in experiments [39].
-However, in this phase-field model the defect chemistry is not explicitly included, and thus there is no simulation results dedicated to the influence of defect energetics on SCLs and abnormal grain growth.
+developed a phase-field model based on KWC grain growth model and investigate the SCL formation at equilibrium and quasi-equilibrium state with the consideration of the GB misorientation and solute drag effect vikrant2020electrochemical .
+They considered the negative segregation formation energy of oxygen vacancies for SCL formation (and unfortunately treated oxygen vacancies and cation dopants as sharing the same crystallographic site).
+Then, this model was applied to explain the abnormal grain growth with a bimodal pattern as observed in experiments vikrant2020electrochemicallyActa .
+In addition to free energy based phase-field model, Aagesen et al.
+developed a grand potential-based phase-field model and included includes the effects of charged vacancies and the associated interactions between internal and applied electric fields aagesen2024electrochemical .
+In oxide ceramics such as Fe-doped STO, Fe atoms substitute for Ti4+ ions on the B-site of the perovskite sublattice, while charge compensation typically occurs via the formation of oxygen vacanies on the anion site.
+As a result, in the bulk phase, the number of available sites for an acceptor dopant corresponds to the number density of B-sites, which is $1/a^{3}$, where $a$ is the lattice parameter.
+The number of available sites indicates the maximum number of sites occupied by defects.
+In contrast, the number of available sites for an oxygen vacancy is three times higher, i.e.
+$3/a^{3}$.
+Even more important, B-site dopant and oxygen vacancy belong to different sub-lattices and thus have to be accounted for separately, in order to correctly represent the respective configurational entropy maier2023physical .
+In addition, the local structure within the GB core differs from that of the bulk.
+These structural differences can significantly alter the number of available sites for both dopants and oxygen vacancy.
+Therefore, it is necessary to treat the number of available sites for oxygen vacancies and dopants in the GB core as distinct from those in the bulk and these quantities play a critical role in determining the SCL.
+Variations in the number of available sites for oxygen vacancy and dopant give rise to diverse SCL formation behaviors de2009formation .
+Although Vikrant and Aagesen performed landmark studies on the space charge layer formation and grain microstructure evolution during the sintering process, the accurate coupling of these defect chemistry parameters into phase-field simulations was not addressed.
 […]
-In the present work, we develop a phase-field grain growth model founded on the principle of defect-chemistry to simulate SCL formation and abnormal grain growth during sintering of oxide electroceramics.
-By taking into account the formation energy differences and the equilibrium defect concentrations in bulk and GB core, the thermodynamic descriptions of point defects are included in the phase-field model, along with the kinetics of grain boundary motion and defect diffusion.
-The spatio-temporal distributions of defect concentration and electrostatic potential in the grain interior and at GB are then solved self-consistently.
+Therefore, defect-chemistry consistent SCLs are characterized by a number of inequivalent defects, each with its unique segregation energy and number of available sites.
+Some GB cores exhibit not only a negative segregation formation energy for oxygen vacancies ($\Delta_{\text{seg}}g_{\text{V}_{\text{O}}^{..}}<0$), but also a combined negative segregation formation energy for oxygen vacancies and acceptor cations ($\Delta_{\text{seg}}g_{\text{V}_{\text{O}}^{..}}<0$ and $\Delta_{\text{seg}}g_{\text{dop}}<0$) usler2024point .
+Meanwhile, different defects in the bulk and GB core require different numbers of available sites usler2024point , usler2024space , de2019effect .
+Moreover, the dominant effect leading to abnormal grain growth during the sintering process in oxide electroceramics remains debatable.
+The combined effects of GB misorientation and solute drag have been explored in Fe-doped STO vikrant2020electrochemical , vikrant2020electrochemicallyActa .
+Nevertheless, an investigation of GB energy anisotropy by the GB grooving technique indicates that the reduction of GB energy due to anisotropy can only lead to moderate (about 15%) changes in grain growth rate kelly2018anti .
+Another investigation of GB mobility anisotropy pointed out the GB mobility anisotropy transition temperature lies in the abnormal grain growth regime and can impact the fast-type and low-type GBs.
+However, it does not seem as the grain growth transition is caused by a change in the GB anisotropy.
+Instead, studies suggest that the transition is associated with a change in the GB stoichiometry, while the atomistic GB structure remains largely unchanged through the grain growth transition sternlicht2016mechanism , shih2010investigation , baurer2010abnormal , sternlicht2019characterization , sternlicht2019characterization2 .
+Additionally, very recent experimental observations demonstrate that the abnormal grain growth in Fe-doped STO during sintering is strongly influenced by dopant concentration zahler2023grain .
+Nominally undoped STO, exhibits a monomodal grain size distribution despite the presence of GB misorientation.
+Hence, this concentration-dependent grain size distribution highlights the critical role of solute drag in driving abnormal grain growth, rather than GB misorientation.
+Accordingly, we develop a phase-field grain growth model grounded in the principles of defect chemistry to reproduce SCL formation and investigate the role of solute drag effects during the sintering process.
+We exclude additional influences, such as GB misorientation on the grain growth process, on purpose, and in this way, we demonstrate that, even in the absence of these factors, the essential features of skewed grain size distribution which do not follow the normal grain growth can still be observed solely through the solute drag effect.
+The novelty of this work lies in three key aspects.
+First, this phase-field model accounts for unique segregation energies and distinct site densities of different defects in both the bulk phase and GB core, ensuring that the defect chemistry of both bulk and core are treated in a thermodynamically consistent description.
+Second, the capabilities of the phase-field model are rigorously benchmarked against sharp interface calculations, providing validation and accuracy.
+Third, the phase-field results reveal that solute drag alone is sufficient to lead to the skewed grain size distribution, confirming its role as the essential mechanism underlying this phenomenon.
 The outline of this paper is as follows.
@@
 The capabilities of this model are discussed for bicrystal and polycrystalline cases in Section 5.
-Thereby the solute drag induced abnormal grain growth simulation results are systematically presented and discussed, particularly in terms of the impact of segregation energies.
+Thereby the solute drag induced skewed grain size distribution are systematically presented and discussed, particularly in terms of the impact of different defect chemistry parameters.
 It is followed by a concluding Section 6.
 2 Defect chemistry of oxide ceramics
+In this section, we review the classical defect chemistry theory in determining the SCL formation under equilibrium state.
+The performed defect chemistry calculations are used to benchmark the phase-field simulations.
 The basics of defect chemistry and the sharp interface bicrystal models (the MS and the GC models) for the equilibrium state were well described in Refs.
-[30, 40, 16].
-Thereby the GB is treated as sharp interface with GB core properties.
-Thus the concentration profiles of oxygen vacancy and dopant are discontinuous, with concentration jumps across the GB core.
+de2009formation , tong2020analyzing , maier2023physical .
+According to abrupt GB||SCL model of Jamnik et al.
+jamnik1995interfaces , the interface region is divided into a GB core and SCL region, in which the materials constants behave as step functions.
+The two phases (GB core and bulk) are essentially treated as two phases, each with their own properties, and separated by a sharp interface.
+Thus the concentration profiles of oxygen vacancy and dopant , for $\mu^{0}_{\text{V}_{\text{O}}^{..}}\neq 0$ and $\mu^{0}_{\text{dop}}\neq 0$, are discontinuous, with concentration jumps across the GB core.
 For the bicrystal, it becomes a one-dimensional case.
-Assume a coordinate system $x$, with the origin $x=0$ located at the GB core.
-Moreover, point defects are assumed to be dilute and non-interacting in both GB core and bulk phases.
-Therefore, the electrochemical potentials of oxygen vacancy and dopant in the bulk and in the core region can be formulated [30] as
+Assume a symmetric bicrystal in a coordinate system $x$, with the origin $x=0$ located at the GB core.
+We introduce $c_{\text{def}}$ point defects into $N_{\text{def}}$ (the subscript def denotes different types of point defects) available number sites in a unit volume.
+We should note that $N_{\text{def}}$ varies for different sublattices and different phases.
+When $N_{\text{def}}\gg c_{\text{def}}$, point defects are assumed to be dilute and non-interacting in both GB core and bulk phases.
+In Ref.
+maier2023physical , the electrochemical potential of defect is formulated as
+$\mu_{\text{def}}=\mu^{0}_{\text{def}}+k_{\text{B}}T\ln\frac{c_{\text{def}}}{N_{\text{def}}-c_{\text{def}}}+z_{\text{def}}e\phi,$ (1)
+where $\mu^{0}_{\text{def}}$ is the standard formation energy of defect.
+$k_{\text{B}}$, $T$, $z_{\text{def}}$, $e$ and $\phi$ are the Boltzmann constant, temperature, valance state of the defect, the elementary charge and the electrostatic potential, respectively.
+Then, based on the abrupt GB||SCL model aforementioned, the electrochemical potentials of oxygen vacancy and dopant in the bulk and in the core region can be formulated de2009formation as
 $\mu_{\text{V}_{\text{O}}^{..},\text{b}}(x)=\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{0}+k_{\text{B}}T\ln\left[\frac{c_{\text{V}_{\text{O}}^{..},\text{b}}(x)}{N_{\text{V}_{\text{O}}^{..},\text{b}}-c_{\text{V}_{\text{O}}^{..},\text{b}}(x)}\right]+z_{\text{V}_{\text{O}}^{..}}e\phi(x),$ (2)
@@
 $z_{\text{V}_{\text{O}}^{..}}$ and $z_{\text{dop}}$ denote the charge numbers of oxygen vacancy and acceptor dopant.
-$\phi(x)$ is the electrostatic potential, and $\Phi_{0}$ is the electrostatic potential in the GB core.
+$\phi(x)$ is the spatially various electrostatic potential in the bulk phase, and $\Phi_{0}$ is the electrostatic potential in the GB core, which taken to be constant.
 Similarly, $c_{\text{dop},\text{c}}$ and $c_{\text{V}_{\text{O}}^{..},\text{c}}$ are the concentrations of dopant and oxygen vacancy in the GB core.
 Thus, we have $\Phi_{0}=\phi(0)$, $c_{\text{dop},\text{c}}=c_{\text{dop},\text{b}}(0)$ and $c_{\text{V}_{\text{O}}^{..},\text{c}}=c_{\text{V}_{\text{O}}^{..},\text{b}}(0)$.
-Additionally, $k_{\text{B}}$, $e$ and $T$ are the Boltzmann constant, elementary charge and temperature, respectively.
 The formation of a SCL requires the segregation of oxygen vacancy in the core.
@@
 $\mu_{\text{dop},\text{c}}^{0}$ Formation energy of dopant in the core region eV
-$\Delta g_{\text{V}_{\text{O}}^{..}}$ The difference of formation energy of oxygen vacancy between core and bulk, i.e.
-$\Delta g_{\text{V}_{\text{O}}^{..}}=\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{0}-\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{0}$ eV
-$\Delta g_{\text{dop}}$ The difference of formation energy of dopant between core and bulk, i.e. $\Delta g_{\text{dop}}=\mu_{\text{dop},\text{c}}^{0}-\mu_{\text{dop},\text{b}}^{0}$ eV
-$N_{\text{V}_{\text{O}}^{..},\text{b}}$ Maximum number density of oxygen vacancy in the bulk phase $\frac{1}{\text{m}^{3}}$
-$N_{\text{V}_{\text{O}}^{..},\text{c}}$ Maximum number density of oxygen vacancy in the core region $\frac{1}{\text{m}^{3}}$
-$N_{\text{dop},\text{b}}$ Maximum number density of dopant in the bulk phase. $N_{\text{dop},\text{b}}=1/a^{3}$, with $a$ being the perovskite lattice constant $\frac{1}{\text{m}^{3}}$
-$N_{\text{dop},\text{c}}$ Maximum number density of dopant in the core region $\frac{1}{\text{m}^{3}}$
-$z_{\text{V}_{\text{O}}^{..}}$ Valency number of oxygen vacancy, $z_{\text{V}_{\text{O}}^{..}}=2$
+$\Delta g_{\text{V}_{\text{O}}^{..}}$
+The difference of formation energy of oxygen vacancy between core and bulk,
+i.e. $\Delta g_{\text{V}_{\text{O}}^{..}}=\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{0}-\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{0}$
+eV
+$\Delta g_{\text{dop}}$
+The difference of formation energy of dopant between core and bulk,
+i.e. $\Delta g_{\text{dop}}=\mu_{\text{dop},\text{c}}^{0}-\mu_{\text{dop},\text{b}}^{0}$
+eV
+$N_{\text{V}_{\text{O}}^{..},\text{b}}$ Number of available sites of oxygen vacancy in the bulk phase $\frac{1}{\text{m}^{3}}$
+$N_{\text{V}_{\text{O}}^{..},\text{c}}$ Number of available sites of oxygen vacancy in the core region $\frac{1}{\text{m}^{3}}$
+$N_{\text{dop},\text{b}}$
+Number of available sites of dopant in the bulk phase.
+$N_{\text{dop},\text{b}}=1/a^{3}$, with $a$ being the perovskite lattice constant
+$\frac{1}{\text{m}^{3}}$
+$N_{\text{dop},\text{c}}$ Number of available sites of dopant in the core region $\frac{1}{\text{m}^{3}}$
+$z_{\text{V}_{\text{O}}^{..}}$ Valency number of oxygen vacancy, $z_{\text{V}_{\text{O}}^{..}}=2$ -
 $z_{\text{dop}}$ Valency number of dopant, $z_{\text{dop}}=-1$ for $\text{Fe}_{\text{Ti}}^{{}^{\prime}}$ in Fe-doped STO -
@@
 $w_{\text{c}}$ width of the core m
+$l_{\text{D}}$ Debye length m
 2.1 The Mott-Schottky model
@@
 Note that this assumption simplifies the scenario but breaks the equilibrium condition of the electrochemical potential of the dopants.
-Based on the assumption, Poisson’s equation can be written as
+Based on this assumption, Poisson’s equation can be written as
 $\begin{split}\epsilon_{0}\epsilon_{\text{r}}\frac{\partial^{2}\phi}{\partial x^{2}}&=-\rho(x)\\ &=-e[z_{\text{dop}}c_{\text{dop},\text{b}}(\infty)+z_{\text{V}_{\text{O}}^{..}}c_{\text{V}_{\text{O}}^{..},\text{b}}(x)].\end{split}$ (12)
-[…]
+Substituting $c_{\text{dop},\text{b}}(\infty)=2c_{\text{V}_{\text{O}}^{..},\text{b}}(\infty)$ into Eq. (12), we have
 $\epsilon_{0}\epsilon_{\text{r}}\frac{\partial^{2}\phi}{\partial x^{2}}=2ec_{\text{V}_{\text{O}}^{..},\text{b}}(\infty)\left\{1-\exp\left[\frac{-2e\{\phi(x)-\phi(\infty)\}}{k_{\text{B}}T}\right]\right\}.$ (13)
-[…]
+With $l_{\text{D}}=\sqrt{\frac{\epsilon_{0}\epsilon_{\text{r}}k_{\text{B}}T}{2e^{2}c_{\text{dop},\text{b}}(\infty)}}$ being the Debye length and $\alpha=\frac{e}{k_{\text{B}}T}$, Eq.(13) becomes
 $\frac{\partial^{2}\phi}{\partial x^{2}}=\frac{1}{2\alpha l_{\text{D}}^{2}}\left\{1-\exp\left[\frac{-2e\{\phi(x)-\phi(\infty)\}}{k_{\text{B}}T}\right]\right\},$ (14)
@@
 $\begin{split}\epsilon_{0}\epsilon_{\text{r}}\frac{\partial^{2}\phi}{\partial x^{2}}&=-e\left\{z_{\text{dop}}\frac{c_{\text{dop},\text{b}}(\infty)N_{\text{dop},\text{b}}\exp\left[\frac{-z_{\text{dop}}e\{\phi(x)-\phi(\infty)\}}{k_{\text{B}}T}\right]}{N_{\text{dop},\text{b}}+c_{\text{dop},\text{b}}(\infty)\left\{\exp\left[\frac{-z_{\text{dop}}e\{\phi(x)-\phi(\infty)\}}{k_{\text{B}}T}\right]-1\right\}}\right.\\ &\left.+z_{\text{V}_{\text{O}}^{..}}c_{\text{V}_{\text{O}}^{..},\text{b}}(\infty)\exp\left[\frac{-z_{\text{V}_{\text{O}}^{..}}e\{\phi(x)-\phi(\infty)\}}{k_{\text{B}}T}\right]\right\}.\end{split}$ (16)
-The numerical solutions of the MS and GC models have been achieved through an iteration process [30].
+The numerical solutions of the MS and GC models have been achieved through an iterative process de2009formation .
 In the present work, we propose an alternative numerical method via finite element method (FEM).
-By considering the symmetric configuration of a bicrysal (ignoring the grain misorientation), the semi-infinite simulation domain of $[0,\infty)$ is sufficient.
+By considering the symmetric configuration of a bicrysal (ignoring the grain misorientation), the semi-infinite simulation domain of $0,\infty)$ is sufficient.
 The location of the GB core is located at the left boundary of the semi-infinite domain ($x=0$).
@@
 The obtained numerical solutions agree fully with these from Ref.
-[30] (See Supplementary).
+de2009formation (See Supplementary).
 3 Phase-field grain growth model with defect chemistry and electrostatics
 3.1 Free energy density functional and electrochemical potentials
-We include here the defect chemistry into the grain growth model for oxide ceramics, with Fe-doped STO as a prototype.
-Upon grain contact in Fe-doped STO, the lower formation energy of oxygen vacancy in the core compels point defects to segregate towards the GB core.
-Afterwards, the depletion or accumulation zone appears in the vicinity of GB core.
-The SCL forms through a diffusion process.
-The handling of SCL is similarly addressed in Ref.
-[30].
-The segregation of point defects in the core serves to dissipate the total free energy of the system and also induces solute drag effect when the core moves.
-Assuming an Fe-doped STO polycrystalline system consisting of $n$ grains, we choose Kim-Kim-Suzuki (KKS) phase-field model and utilize a set of non-conserved order parameters (OPs) ${\eta_{i}}$ to distinguish between different grains [41].
-Within grain $i$, $\eta_{i}=1$ and $\eta_{j}=0$ for $j<n$ and $j\neq i$.
-Within the core region of grain $i$ and $j$, the order parameters $\eta_{i}$ and $\eta_{j}$ spatially vary from 0 to 1.
+In this part, we aim to develop a defect-chemistry informed phase-field model including not only the dominant driving force of SCL formation, i.e., segregation energy of oxygen vacancy, but also the segregation energy of acceptor dopant.
+The segregation energy of acceptor dopant can not govern the SCL formation, whereas it plays an important role on GB potential usler2024point , which is barely investigated by phase-field simulations reported in previous literature.
+In addition, we also include different available number sites of oxygen vacancy and acceptor dopant in bulk and GB core, i.e.
+$N_{\text{V}_{\text{O}}^{..},\text{b}}\neq N_{\text{dop},\text{b}}$ and $N_{\text{V}_{\text{O}}^{..},\text{c}}\neq N_{\text{dop},\text{c}}$.
+Assuming an Fe-doped STO polycrystalline system consisting of $n$ grains, we choose free energy based Kim-Kim-Suzuki (KKS) phase-field model and utilize a set of non-conserved order parameters (OPs) $\eta_{i}$} to distinguish between different grains kim1999phase .
+Then, by treating the GB core and bulk as distinct phases with separate segregation energies cha2002phase , the phase-field model presented here aligns with the assumptions of the abrupt GB||SCL model (see Section 2) and inherently incorporates solute drag effects.
+Therefore, within grain $i$, we have $\eta_{i}=1$ and $\eta_{j}=0$ for $j<n$ and $j\neq i$.
+Within the GB core region of grain $i$ and $j$, the order parameters $\eta_{i}$ and $\eta_{j}$ spatially vary from 0 to 1.
 In the latter context, $\eta$ denotes the entire set of OPs $\eta_{i}$}.
@@
 A1 in Ref.
-[42]).
+merkle2020water ).
 Thus, in this study, we assume the valence state of Fe to be 3+ and therefore the valance of dopant $\text{Fe}_{\text{Ti}}^{{}^{\prime}}$ is -1, i.e.
@@
 $\mathscr{F}=\int_{\Omega}(f^{\text{grad}}+f^{\text{loc}}+f^{\text{ech}})\text{d}\Omega.$ (18)
-$f^{\text{grad}}$ represents the gradient term, it is given by
+The free energy density functional contains three contributions. The first term $f^{\text{grad}}$ represents the gradient energy of the diffuse interface and is given by
 $f^{\text{grad}}=\frac{1}{2}\kappa\sum_{i=1}^{n}|\nabla\eta_{i}|^{2},$ (19)
-where $\kappa$ is the coefficient of the gradient term. $f^{\text{loc}}$ is the well potential. According to Ref. [43], its formulation is given by
+where $\kappa$ is the coefficient of the gradient term.
+The second term $f^{\text{loc}}$ is the multi-well potential and it reaches minimum value in the bulk phase.
+According to Ref.
+moelans2008quantitative , aagesen2024electrochemical , its formulation is given by
 $f^{\text{loc}}=\omega\sum_{i=1}^{n}\left(\frac{\eta_{i}^{4}}{4}-\frac{\eta_{i}^{2}}{2}+\gamma\sum_{i=1}^{n}\sum_{j>i}^{n}\eta_{i}^{2}\eta_{j}^{2}+\frac{1}{4}\right),$ (20)
-with $\gamma=1.5$ for producing symmetric profile of $\eta$ and isotropic grain growth.
-The determination of parameters such as $\kappa$ and $\omega$ can be linked to characteristics such as core width ($w_{\text{c}}$) and interface energy ($\sigma$). Their relations can be expressed as
+with $\omega$ being the free energy barrier coefficient, $\gamma=1.5$ for producing symmetric profile of $\eta$ and isotropic grain growth.
+The determination of parameters such as $\kappa$ and $\omega$ can be linked to characteristics such as GB core width ($w_{\text{c}}$) and GB energy ($\sigma$) moelans2008quantitative .
+Their relations can be expressed as
 $\sigma=\frac{\sqrt{2}}{3}\sqrt{\kappa\omega},$ (21)
@@
 $w_{\text{c}}=\sqrt{\frac{8\kappa}{\omega}}.$ (22)
-In the present work, we only consider the isotropic interface energy, i.e.
+In the present work, we only consider the isotropic GB energy, i.e.
 it is not dependent on GB misorientation.
 The third term in free energy functional is the electrochemical contribution.
-Here, we treat the electrochemical free energy density as a mixture of bulk phase and core by an interpolation function.
-Thus, $f^{\text{ech}}$ is
+Here, we treat the electrochemical free energy density as a mixture of bulk phase and GB core by an interpolation function.
+Thus, with the consideration of electrostatic free energy according to Refs.
+bishop2003effect , garcia2004thermodynamically , $f^{\text{ech}}$ is
 $f^{\text{ech}}=[1-h(\eta)]f^{\text{ech}}_{\text{b}}(C_{\text{V}_{\text{O}}^{..},\text{b}},C_{\text{dop},\text{b}})+h(\eta)f^{\text{ech}}_{\text{c}}(C_{\text{V}_{\text{O}}^{..},\text{c}},C_{\text{dop},\text{c}})-\frac{\epsilon_{0}\epsilon_{\text{r}}}{2}(\nabla\phi)^{2},$ (23)
-where the subscripts b and c indicate the bulk phase and core, respectively. $h(\eta)$ is the interpolation function and is given by
+where the subscripts b and c indicate the bulk phase and GB core, respectively. $h(\eta)$ is the interpolation function and is given by
 $h(\eta)=\frac{4}{3}\left[1-4\sum_{i=1}^{n}\eta_{i}^{3}+3\left(\sum_{i=1}^{n}\eta_{i}^{2}\right)^{2}\right].$ (24)
@@
 $\epsilon_{0}$ is the vacuum permittivity; $\epsilon_{\text{r}}$ is the relative permittivity.
-For STO, $\epsilon_{\text{r}}=9\times 10^{4}/(T-35)$ [30].
+For STO, $\epsilon_{\text{r}}=9\times 10^{4}/(T-35)$ de2009formation .
 $\epsilon_{\text{r}}$ is only dependent on temperature in the present work.
@@
 $C_{\text{dop}}=[1-h(\eta)]C_{\text{dop},\text{b}}+h(\eta)C_{\text{dop},\text{c}}.$ (26)
-In principle of the thermodynamic model extended by Ref.
-[30], the expressions for the electrochemical free energies $f^{\text{ech}}_{\text{c}}(C_{\text{V}_{\text{O}}^{..},\text{c}},C_{\text{dop},\text{c}})$ and $f^{\text{ech}}_{\text{b}}(C_{\text{V}_{\text{O}}^{..},\text{b}},C_{\text{dop},\text{b}})$ are given by
-$\begin{split}f^{\text{ech}}_{\text{c}}(C_{\text{V}_{\text{O}}^{..},\text{c}},C_{\text{dop},\text{c}})&=\frac{1}{V_{\text{m}}}\left\{g^{0}_{\text{V}_{\text{O}}^{..},\text{c}}C_{\text{V}_{\text{O}}^{..},\text{c}}+g^{0}_{\text{dop},\text{c}}C_{\text{dop},\text{c}}\right.\\ &+RT\left[C_{\text{V}_{\text{O}}^{..},\text{c}}\ln(C_{\text{V}_{\text{O}}^{..},\text{c}})+(\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}-C_{\text{V}_{\text{O}}^{..},\text{c}})\ln(\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}-C_{\text{V}_{\text{O}}^{..},\text{c}})\right.\\ &\left.\left.+C_{\text{dop},\text{c}}\ln(C_{\text{dop},\text{c}})+(\tilde{N}_{\text{dop},\text{c}}-C_{\text{dop},\text{c}})\ln(\tilde{N}_{\text{dop},\text{c}}-C_{\text{dop},\text{c}})\right]\right\}\\ &+\frac{\mathcal{F}}{V_{\text{m}}}\left(z_{\text{V}_{\text{O}}^{..}}C_{\text{V}_{\text{O}}^{..},\text{c}}+z_{\text{dop}}C_{\text{dop},\text{c}}\right)\phi,\end{split}$ (26)
-$\begin{split}f^{\text{ech}}_{\text{b}}(C_{\text{V}_{\text{O}}^{..},\text{b}},C_{\text{dop},\text{b}})&=\frac{1}{V_{\text{m}}}\left\{g^{0}_{\text{V}_{\text{O}}^{..},\text{b}}C_{\text{V}_{\text{O}}^{..},\text{b}}+g^{0}_{\text{dop},\text{b}}C_{\text{dop},\text{b}}\right.\\ &+RT\left[C_{\text{V}_{\text{O}}^{..},\text{b}}\ln(C_{\text{V}_{\text{O}}^{..},\text{b}})+(\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}-C_{\text{V}_{\text{O}}^{..},\text{b}})\ln(\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}-C_{\text{V}_{\text{O}}^{..},\text{b}})\right.\\ &\left.\left.+C_{\text{dop},\text{b}}\ln(C_{\text{dop},\text{b}})+(\tilde{N}_{\text{dop},\text{b}}-C_{\text{dop},\text{b}})\ln(\tilde{N}_{\text{dop},\text{b}}-C_{\text{dop},\text{b}})\right]\right\}\\ &+\frac{\mathcal{F}}{V_{\text{m}}}\left(z_{\text{V}_{\text{O}}^{..}}C_{\text{V}_{\text{O}}^{..},\text{b}}+z_{\text{dop}}C_{\text{dop},\text{b}}\right)\phi,\end{split}$ (27)
+In principle of defect chemistry thermodynamic model as mentioned in Section 2, the electrochemical free energies $f^{\text{ech}}_{\text{c}}(C_{\text{V}_{\text{O}}^{..},\text{c}},C_{\text{dop},\text{c}})$ and $f^{\text{ech}}_{\text{b}}(C_{\text{V}_{\text{O}}^{..},\text{b}},C_{\text{dop},\text{b}})$ deliver the required equilibrium thermodynamic information in bulk and GB core and their expressions are given by
+$\begin{split}f^{\text{ech}}_{\text{c}}(C_{\text{V}_{\text{O}}^{..},\text{c}},C_{\text{dop},\text{c}})&=\frac{1}{V_{\text{m}}}\left\{g^{0}_{\text{V}_{\text{O}}^{..},\text{c}}C_{\text{V}_{\text{O}}^{..},\text{c}}+g^{0}_{\text{dop},\text{c}}C_{\text{dop},\text{c}}\right.\\ &+RT\left[C_{\text{V}_{\text{O}}^{..},\text{c}}\ln\left(\frac{C_{\text{V}_{\text{O}}^{..},\text{c}}}{\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}}\right)+(\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}-C_{\text{V}_{\text{O}}^{..},\text{c}})\ln\left(\frac{\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}-C_{\text{V}_{\text{O}}^{..},\text{c}}}{\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}}\right)\right.\\ &\left.\left.+C_{\text{dop},\text{c}}\ln\left(\frac{C_{\text{dop},\text{c}}}{\tilde{N}_{\text{dop},\text{c}}}\right)+(\tilde{N}_{\text{dop},\text{c}}-C_{\text{dop},\text{c}})\ln\left(\frac{\tilde{N}_{\text{dop},\text{c}}-C_{\text{dop},\text{c}}}{\tilde{N}_{\text{dop},\text{c}}}\right)\right]\right\}\\ &+\frac{\mathcal{F}}{V_{\text{m}}}\left(z_{\text{V}_{\text{O}}^{..}}C_{\text{V}_{\text{O}}^{..},\text{c}}+z_{\text{dop}}C_{\text{dop},\text{c}}\right)\phi,\end{split}$ (27)
+$\begin{split}f^{\text{ech}}_{\text{b}}(C_{\text{V}_{\text{O}}^{..},\text{b}},C_{\text{dop},\text{b}})&=\frac{1}{V_{\text{m}}}\left\{g^{0}_{\text{V}_{\text{O}}^{..},\text{b}}C_{\text{V}_{\text{O}}^{..},\text{b}}+g^{0}_{\text{dop},\text{b}}C_{\text{dop},\text{b}}\right.\\ &+RT\left[C_{\text{V}_{\text{O}}^{..},\text{b}}\ln\left(\frac{C_{\text{V}_{\text{O}}^{..},\text{b}}}{\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}}\right)+(\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}-C_{\text{V}_{\text{O}}^{..},\text{b}})\ln\left(\frac{\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}-C_{\text{V}_{\text{O}}^{..},\text{b}}}{\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}}\right)\right.\\ &\left.\left.+C_{\text{dop},\text{b}}\ln\left(\frac{C_{\text{dop},\text{b}}}{\tilde{N}_{\text{dop},\text{b}}}\right)+(\tilde{N}_{\text{dop},\text{b}}-C_{\text{dop},\text{b}})\ln\left(\frac{\tilde{N}_{\text{dop},\text{b}}-C_{\text{dop},\text{b}}}{\tilde{N}_{\text{dop},\text{b}}}\right)\right]\right\}\\ &+\frac{\mathcal{F}}{V_{\text{m}}}\left(z_{\text{V}_{\text{O}}^{..}}C_{\text{V}_{\text{O}}^{..},\text{b}}+z_{\text{dop}}C_{\text{dop},\text{b}}\right)\phi,\end{split}$ (28)
 where $g^{0}_{\text{V}_{\text{O}}^{..},\text{c}}$, $g^{0}_{\text{V}_{\text{O}}^{..},\text{b}}$, $g^{0}_{\text{dop},\text{c}}$, $g^{0}_{\text{dop},\text{b}}$ are standard formation energies of oxygen vacancy and dopant in bulk and core.
@@
 $\mathcal{F}$, $V_{\text{m}}$, $R$, $T$ are the Faraday constant, molar volume, gas constant and temperature, respectively.
-The electrochemical potentials of oxygen vacancy and dopant in bulk and core are obtained by
-$\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{\text{PF}}=\frac{\delta\mathscr{F}}{\delta C_{\text{V}_{\text{O}}^{..},\text{b}}}=\frac{g_{\text{V}_{\text{O}}^{..},\text{b}}^{0}}{V_{\text{m}}}+\frac{RT}{V_{\text{m}}}\ln{\left[\frac{C_{\text{V}_{\text{O}}^{..},\text{b}}}{\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}-C_{\text{V}_{\text{O}}^{..},\text{b}}}\right]}+\frac{\mathcal{F}}{V_{\text{m}}}z_{\text{V}_{\text{O}}^{..}}\phi,$ (28)
-$\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{\text{PF}}=\frac{\delta\mathscr{F}}{\delta C_{\text{V}_{\text{O}}^{..},\text{c}}}=\frac{g_{\text{V}_{\text{O}}^{..},\text{c}}^{0}}{V_{\text{m}}}+\frac{RT}{V_{\text{m}}}\ln{\left[\frac{C_{\text{V}_{\text{O}}^{..},\text{c}}}{\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}-C_{\text{V}_{\text{O}}^{..},\text{c}}}\right]}+\frac{\mathcal{F}}{V_{\text{m}}}z_{\text{V}_{\text{O}}^{..}}\phi,$ (29)
-$\mu_{\text{dop},\text{b}}^{\text{PF}}=\frac{\delta\mathscr{F}}{\delta C_{\text{dop},\text{b}}}=\frac{g_{\text{dop},\text{b}}^{0}}{V_{\text{m}}}+\frac{RT}{V_{\text{m}}}\ln{\left[\frac{C_{\text{dop},\text{b}}}{\tilde{N}_{\text{dop},\text{b}}-C_{\text{dop},\text{b}}}\right]}+\frac{\mathcal{F}}{V_{\text{m}}}z_{\text{dop}}\phi,$ (30)
-$\mu_{\text{dop},\text{c}}^{\text{PF}}=\frac{\delta\mathscr{F}}{\delta C_{\text{dop},\text{c}}}=\frac{g_{\text{dop},\text{c}}^{0}}{V_{\text{m}}}+\frac{RT}{V_{\text{m}}}\ln{\left[\frac{C_{\text{dop},\text{c}}}{\tilde{N}_{\text{dop},\text{c}}-C_{\text{dop},\text{c}}}\right]}+\frac{\mathcal{F}}{V_{\text{m}}}z_{\text{dop}}\phi,$ (31)
+Based on experimental evidence of abnormal grain growth in 0.2% Fe-doped STO zahler2023grain , where defect-defect interactions are negligible, we limit our consideration to the dilute defect case.
+This simplification ensures that the fundamental characteristics of the system are captured without introducing unnecessary complexities.
+However, the dilute assumption is not valid when high defect concentration is involved, the interactions between different defects should be considered chen2022thermodynamic .
+The electrochemical potentials of oxygen vacancy and dopant in bulk and core are obtained according to
+$\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{\text{PF}}=\frac{\delta\mathscr{F}}{\delta C_{\text{V}_{\text{O}}^{..},\text{b}}}=\frac{g_{\text{V}_{\text{O}}^{..},\text{b}}^{0}}{V_{\text{m}}}+\frac{RT}{V_{\text{m}}}\ln{\left(\frac{C_{\text{V}_{\text{O}}^{..},\text{b}}}{\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}-C_{\text{V}_{\text{O}}^{..},\text{b}}}\right)}+\frac{\mathcal{F}}{V_{\text{m}}}z_{\text{V}_{\text{O}}^{..}}\phi,$ (29)
+$\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{\text{PF}}=\frac{\delta\mathscr{F}}{\delta C_{\text{V}_{\text{O}}^{..},\text{c}}}=\frac{g_{\text{V}_{\text{O}}^{..},\text{c}}^{0}}{V_{\text{m}}}+\frac{RT}{V_{\text{m}}}\ln{\left(\frac{C_{\text{V}_{\text{O}}^{..},\text{c}}}{\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}-C_{\text{V}_{\text{O}}^{..},\text{c}}}\right)}+\frac{\mathcal{F}}{V_{\text{m}}}z_{\text{V}_{\text{O}}^{..}}\phi,$ (30)
+$\mu_{\text{dop},\text{b}}^{\text{PF}}=\frac{\delta\mathscr{F}}{\delta C_{\text{dop},\text{b}}}=\frac{g_{\text{dop},\text{b}}^{0}}{V_{\text{m}}}+\frac{RT}{V_{\text{m}}}\ln{\left(\frac{C_{\text{dop},\text{b}}}{\tilde{N}_{\text{dop},\text{b}}-C_{\text{dop},\text{b}}}\right)}+\frac{\mathcal{F}}{V_{\text{m}}}z_{\text{dop}}\phi,$ (31)
+$\mu_{\text{dop},\text{c}}^{\text{PF}}=\frac{\delta\mathscr{F}}{\delta C_{\text{dop},\text{c}}}=\frac{g_{\text{dop},\text{c}}^{0}}{V_{\text{m}}}+\frac{RT}{V_{\text{m}}}\ln{\left(\frac{C_{\text{dop},\text{c}}}{\tilde{N}_{\text{dop},\text{c}}-C_{\text{dop},\text{c}}}\right)}+\frac{\mathcal{F}}{V_{\text{m}}}z_{\text{dop}}\phi,$ (32)
 which are consistent with the standard formations of electrochemical potentials in defect chemistry, i.e. Eq. 2 to Eq. 5.
 3.2 Governing equations
-In the framework of KKS phase-field model, the evolution of phase-field OPs obeys the Allen-Cahn equation. For grain $i$, the evolution of $\eta_{i}$ in time is
+As mentioned in the introduction, we focus on the solute drag effect, which is the most fundamental and primary factor influencing grain growth patterns.
+Therefore, our phase-field model employs constant GB energy and mobility and uses order parameters to distinguish different grains without explicitly considering grain misorientation.
+Then, the evolution of phase-field OPs for grain $i$ with isotropic GB properties becomes
 $\frac{1}{L}\frac{\partial\eta_{i}}{\partial t}=\frac{\delta\mathscr{F}}{\delta\eta_{i}},$ (33)
@@
 At the quasi-equilibrium state, the core moves with a constant velocity, $v_{c}$, we have $\dot{\eta}=-v_{c}\eta^{\prime}(x)$, thus Eq. (33) becomes
-$\begin{split}\frac{v_{c}}{L}\frac{d\eta}{dx}&=\nabla\cdot\kappa\nabla\eta_{i}-\omega\frac{\partial f^{\text{loc}}(\eta)}{\partial\eta_{i}}\\ &+\frac{\partial h(\eta)}{\partial\eta_{i}}\left[f_{\text{b}}^{\text{ech}}-f_{\text{c}}^{\text{ech}}-\frac{\partial f_{\text{b}}^{\text{ech}}}{\partial C_{\text{V}_{\text{O}}^{..},\text{b}}}(C_{\text{V}_{\text{O}}^{..},\text{b}}-C_{\text{V}_{\text{O}}^{..},\text{c}})-\frac{\partial f_{\text{b}}^{\text{ech}}}{\partial C_{\text{dop},\text{b}}}(C_{\text{dop},\text{b}}-C_{\text{dop},\text{c}})\right],\end{split}$ (34)
-Conserved concentration fields $C_{\text{V}_{\text{O}}^{..}}$ and $C_{\text{dop}}$ follow Fick’s law, which are expressed as
+$\begin{split}\frac{v_{c}}{L}\frac{d\eta}{dx}&=\nabla\cdot\kappa\nabla\eta_{i}-\omega\frac{\partial f^{\text{loc}}(\eta)}{\partial\eta_{i}}\\ &+\frac{\partial h(\eta)}{\partial\eta_{i}}\left[f_{\text{b}}^{\text{ech}}-f_{\text{c}}^{\text{ech}}-\frac{\partial f_{\text{b}}^{\text{ech}}}{\partial C_{\text{V}_{\text{O}}^{..},\text{b}}}(C_{\text{V}_{\text{O}}^{..},\text{b}}-C_{\text{V}_{\text{O}}^{..},\text{c}})-\frac{\partial f_{\text{b}}^{\text{ech}}}{\partial C_{\text{dop},\text{b}}}(C_{\text{dop},\text{b}}-C_{\text{dop},\text{c}})\right].\end{split}$ (35)
+By combining Fick’s first and second laws, the evolution of the conserved concentration fields $C_{\text{V}_{\text{O}}^{..}}$ and $C_{\text{dop}}$ are expressed as
 $\frac{\partial C_{\text{V}_{\text{O}}^{..}}}{\partial t}=\nabla\left(M_{\text{V}_{\text{O}}^{..}}\nabla\frac{\delta\mathscr{F}}{\delta C_{\text{V}_{\text{O}}^{..}}}\right),$ (36)
@@
 Therefore, $\delta\mathscr{F}/\delta C_{\text{V}_{\text{O}}^{..}}$ and $\delta\mathscr{F}/\delta C_{\text{dop}}$ in Eqs.
-35 and 36 can be substituted by $\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{\text{PF}}$, $\mu_{\text{dop},\text{b}}^{\text{PF}}$ or $\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{\text{PF}}$, $\mu_{\text{dop},\text{c}}^{\text{PF}}$.
+36 and 37 can be substituted by $\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{\text{PF}}$, $\mu_{\text{dop},\text{b}}^{\text{PF}}$ or $\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{\text{PF}}$, $\mu_{\text{dop},\text{c}}^{\text{PF}}$.
 In addition, the governing equation of electrostatic potential field is obtained by Poisson’s equation
-$\epsilon_{0}\epsilon_{\text{r}}\nabla^{2}\phi+\frac{\mathcal{F}}{V_{\text{m}}}(z_{\text{V}_{\text{O}}^{..}}C_{\text{V}_{\text{O}}^{..}}+z_{\text{dop}}C_{\text{dop}})=0,$ (39)
+[…]
+Substituting the site fractions $_{\text{V}_{\text{O}}^{..},\text{b}}], [VV$$_{\text{V}_{\text{O}}^{..},\text{c}}$], $V_{\text{dop},\text{b}}]$ and $V_{\text{dop},\text{c}}]$ into Eq. 40, we have
+$\begin{split}\epsilon_{0}\epsilon_{\text{r}}\nabla^{2}\phi=&-\frac{\mathcal{F}}{V_{\text{m}}}\left\{z_{\text{V}_{\text{O}}^{..}}\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}[V_{\text{V}_{\text{O}}^{..},\text{b}}][1-h(\eta)]+z_{\text{V}_{\text{O}}^{..}}\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}[V_{\text{V}_{\text{O}}^{..},\text{c}}]h(\eta)\right.\\ &\left.+z_{\text{dop}}\tilde{N}_{\text{dop},\text{b}}[V_{\text{dop},\text{b}}][1-h(\eta)]+z_{\text{dop}}\tilde{N}_{\text{dop},\text{c}}[V_{\text{dop},\text{c}}]h(\eta)\right\}.\end{split}$ (41)
+We clearly observe the numbers of available sites $\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}$, $\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}$, $\tilde{N}_{\text{dop},\text{b}}$, $\tilde{N}_{\text{dop},\text{c}}$ influence the electrostatic potential distribution in Eq.
+41.
 The parameters utilized in phase-field model are tabulated in table 2.
+3.3 Comparison to other grain growth phase-field models for oxide ceramics
+In this part, we outline differences between our defect-chemistry-informed phase-field model, the model developed by Vikrant et al.
+vikrant2020electrochemical and the grand-potential based phase-field model developed by Aagesen et al.
+aagesen2024electrochemical .
+In accordance with defect chemistry theory maier2023physical , we include the available site densities for the bulk and the grain boundary core explicitly, as shown by the bulk and GB core electrochemical free energy (Eq.
+27 and Eq.
+28) and the oxygen vacancy and dopant electrochemical potentials in bulk and GB core (Eq.
+29-Eq.
+32).
+It is important to do so, because they can impact the space charge layers and GB potential considerably usler2024point .
+As mentioned in the Introduction, in Fe-doped SrTiO3, Fe substitutes for Ti4+ on the B-site, and charge compensation is typically provided by oxygen vacancies on the O-site.
+This results in the number density of available sites for the dopant in the bulk that is one-third of that for an oxygen vacancy.
+Even more important, B-site dopant and O-site oxygen vacancy belong to different sub-lattices and thus have to be accounted for separately, in order to correctly represent the respective configurational entropy.
+In addition, the number of available sites for both species will deviate significantly from their bulk values due to the distinct atomic structure of the GB core.
+Therefore, it is necessary to treat the available site densities for oxygen vacancies and dopants in the GB core as distinct from those in the bulk, and these quantities play a critical role in determining the space charge layer.
+As our simulation results in the following sections demonstrate, the different numbers of available site significantly influence both the grain boundary potential and the defect concentration distributions, consequently, leading to different space charge layers formation.
+Much more results can be found in Figures 2 and 3, validated against De Souza’s work de2009formation .
+Such explicit incorporation of the numbers of available site of oxygen vacancy and dopant in the bulk and GB core is not fully regarded in the existing models.
+For instance, Vikrant et al.
+assumed that the oxygen vacancy and acceptor dopant share the same sublattice in the chemical free energy contribution (see Eq.
+1 in Ref.
+vikrant2020electrochemical ).
+$\begin{split}f^{\text{Vik}}(\eta,[V_{\text{V}_{\text{O}}^{..}}],[V_{\text{dop}}],T)&=\frac{1}{\omega}\left\{f_{\text{V}_{\text{O}}^{..}}(\eta,T)[V_{\text{V}_{\text{O}}^{..}}]+f_{\text{dop}}(\eta,T)[V_{\text{dop}}]+k_{\text{B}}T\left([V_{\text{V}_{\text{O}}^{..}}]\ln[V_{\text{V}_{\text{O}}^{..}}]+[V_{\text{dop}}]\ln[V_{\text{dop}}]\right)\right.\\ &\left.+k_{\text{B}}T\left(1-[V_{\text{V}_{\text{O}}^{..}}]-[V_{\text{dop}}]\right)\ln\left(1-[V_{\text{V}_{\text{O}}^{..}}]-[V_{\text{dop}}]\right)+\Omega_{\text{V}_{\text{O}}^{..}\text{dop}}[V_{\text{V}_{\text{O}}^{..}}][V_{\text{dop}}]\right\},\end{split}$ (42)
+where $f_{\text{V}_{\text{O}}^{..}}(\eta,T)$ and $f_{\text{dop}}(\eta,T)$ are temperature related formation energy of oxygen vacancy and dopant in bulk and GB core.
+$V_{\text{V}_{\text{O}}^{..}}]$ and $V_{\text{dop}}]$ are the site fractions of oxygen vacancy and dopant.
+$\Omega_{\text{V}_{\text{O}}^{..}\text{dop}}$ is the interaction coefficient between oxygen vacancy and dopant.
+At here, we omit $\Omega_{\text{V}_{\text{O}}^{..}\text{dop}}$ for dilute case.
+In Eq.
+42, both oxygen vacancy and dopant are described as occupying the same sublattice, which is reflected in the entropic mixing term $\left(1-[V_{\text{V}_{\text{O}}^{..}}]-[V_{\text{dop}}]\right)\allowbreak\ln\left(1-[V_{\text{V}_{\text{O}}^{..}}]-[V_{\text{dop}}]\right)$.
+Then, we can obtain the electrochemical potential of oxygen vacancy and dopant from Vikrant’s model
+$\mu_{\text{V}_{\text{O}}^{..}}^{\text{Vik}}=\frac{f_{\text{V}_{\text{O}}^{..}}(\eta,T)}{\omega}+\frac{k_{\text{B}}T}{{\omega}}\ln\left(\frac{[V_{\text{V}_{\text{O}}^{..}}]}{1-[V_{\text{V}_{\text{O}}^{..}}]-[V_{\text{dop}}]}\right)+z_{\text{V}_{\text{O}}^{..}}e\phi,$ (43)
+$\mu_{\text{dop}}^{\text{Vik}}=\frac{f_{\text{dop}}(\eta,T)}{\omega}+\frac{k_{\text{B}}T}{\omega}\ln\left(\frac{[V_{\text{dop}}]}{1-[V_{\text{V}_{\text{O}}^{..}}]-[V_{\text{dop}}]}\right)+z_{\text{dop}}e\phi,$ (44)
+which is different from the electrochemical potential obtained from defect-chemistry as shown in Eq.
+1-Eq.
+5 and present work in Eq.
+29 - Eq.
+32.
+In contrast, our model explicitly distinguishes between the numbers of available site of oxygen vacancy and dopant in both bulk and GB core, i.e.
+$\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}$, $\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}$, $\tilde{N}_{\text{dop},\text{b}}$, $\tilde{N}_{\text{dop},\text{c}}$.
+The configurational entropy term in the electrochemical free energy part is $\left(\tilde{N}_{i}-C_{i}\right)\ln\left(\frac{\tilde{N}_{i}-C_{i}}{\tilde{N}_{i}}\right)$, with the subscript $i$ indicates $\text{V}_{\text{O}}^{..},\text{b}$, $\text{V}_{\text{O}}^{..},\text{c}$, $\text{dop},\text{b}$ and $\text{dop},\text{c}$.
+This separation allows us to represent the thermodynamics of defect segregation more accurately at grain boundary core.
+Moreover, the numbers of available site of oxygen vacancy and acceptor dopant in both the bulk and GB core explicitly influence the Poisson’s equation and the electrostatic potential distribution in our model (see Eq.
+41), However, in Vikrant’s model, the Poisson’s equation is $\epsilon_{0}\epsilon_{r}\nabla^{2}\phi=-e(z_{\text{V}_{\text{O}}^{..}}[V_{\text{V}_{\text{O}}^{..}}]+z_{\text{dop}}[V_{\text{dop}}])$, the influence of these parameters, $\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}$, $\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}$, $\tilde{N}_{\text{dop},\text{b}}$, $\tilde{N}_{\text{dop},\text{c}}$, on the electrostatic potential distribution is not considered.
+Aagsen’s model is developed based on the grand potential functional.
+The thermodynamic equivalence between the free energy based KKS model and grand potential based phase-field model developed has been clarified in Ref.
+plapp2011unified (see section VA).
+However, in Aagsen’s model, the numbers of available site (denoted as $n_{\text{Y}}$ and $n_{\text{O}}$) of both cation vacancy and oxygen vacancy in Y2O3 are treated as uniform across the bulk and GB core, i.e.
+$n_{\text{Y,b}}=n_{\text{Y,c}}$ and $n_{\text{O,b}}=n_{\text{O,c}}$, which may overlook the important structural variations at GB core, especially when $n_{\text{Y,b}}\neq n_{\text{Y,c}}$ and $n_{\text{O,b}}\neq n_{\text{O,c}}$, and result in inappropriate space charge layer formation.
+These scenarios have been comprehensively discussed in De Souza’s work de2009formation and demonstrated in the present work (see Fig.
+2 and Fig.
+3) in Section 5.1.
+In addition, such explicit consideration of numbers of available site in our formulation paves the foundation for the subsequent extension towards the treatment of charge state transition of dopants, such as Fe4+/Fe3+ and Fe3+/Fe2+ transitions in Fe-doped SrTiO3 klein2023fermi, bonkowski2024single.
+Vikrant’s work, Aagsen’s work and the present work primarily focus on a single valence state (typically -1), which limits their applicability under varying thermodynamic conditions.
+The dominant charge state of Fe can vary significantly with temperature and oxygen partial pressure wang2016defect .
+These charge states all share the same B-site sublattice in the perovskite structure, which reinforces the importance of correctly defining and distinguishing the numbers of available site between oxygen vacancy and dopant in bulk phase and grain boundary core in the phase-field model.
+Without completely considering the defect chemistry, the extension of Vikrant’s model or Aagsen’s model in this direction is not applicable due to the mixture of the numbers of available site.
+However, our defect-chemistry phase-field is developed to accommodate this level of resolution, enabling us to simulate space charge layer formation with proper consideration of the different valence states of dopant and their impact on different space charge layer formations.
 Table 2: Summary of the parameters in the phase-field model
 DOFs Definition Unit
-$\eta_{i}$ phase-field order parameter of grain $i$, $i=1,2,3,...,n$
-$C_{\text{V}_{\text{O}}^{..}}$ Normalized local concentration of oxygen vacancy ($C_{\text{V}_{\text{O}}^{..}}=a^{3}c_{\text{V}_{\text{O}}^{..}}$, with $a$ being the lattice constant) -
-$C_{\text{dop}}$ Normalized local concentration of dopant, $C_{\text{dop}}=a^{3}c_{\text{dop}}$
-$C_{\text{V}_{\text{O}}^{..},\text{b}}$ Normalized concentration of oxygen vacancy in bulk phase, $C_{\text{V}_{\text{O}}^{..},\text{b}}=a^{3}c_{\text{V}_{\text{O}}^{..},\text{b}}$
-$C_{\text{V}_{\text{O}}^{..},\text{c}}$ Normalized concentration of oxygen vacancy in core region, $C_{\text{V}_{\text{O}}^{..},\text{c}}=a^{3}c_{\text{V}_{\text{O}}^{..},\text{c}}$
-$C_{\text{dop},\text{b}}$ Normalized concentration of dopant in bulk phase, $C_{\text{dop},\text{b}}=a^{3}c_{\text{dop},\text{b}}$
-$C_{\text{dop},\text{c}}$ Normalized concentration of dopant in core region, $C_{\text{dop},\text{c}}=a^{3}c_{\text{dop},\text{c}}$
+$\eta_{i}$ phase-field order parameter of grain $i$, i=1,2,3,,ni=1,2,3,...,n -
+$C_{\text{V}_{\text{O}}^{..}}$
+Normalized local concentration of oxygen vacancy
+($C_{\text{V}_{\text{O}}^{..}}=a^{3}c_{\text{V}_{\text{O}}^{..}}$, with $a$ being the lattice constant)
+$C_{\text{dop}}$ Normalized local concentration of dopant, $C_{\text{dop}}=a^{3}c_{\text{dop}}$ -
+$C_{\text{V}_{\text{O}}^{..},\text{b}}$ Normalized concentration of oxygen vacancy in bulk phase, $C_{\text{V}_{\text{O}}^{..},\text{b}}=a^{3}c_{\text{V}_{\text{O}}^{..},\text{b}}$ -
+$C_{\text{V}_{\text{O}}^{..},\text{c}}$ Normalized concentration of oxygen vacancy in core region, $C_{\text{V}_{\text{O}}^{..},\text{c}}=a^{3}c_{\text{V}_{\text{O}}^{..},\text{c}}$ -
+$C_{\text{dop},\text{b}}$ Normalized concentration of dopant in bulk phase, $C_{\text{dop},\text{b}}=a^{3}c_{\text{dop},\text{b}}$ -
+$C_{\text{dop},\text{c}}$ Normalized concentration of dopant in core region, $C_{\text{dop},\text{c}}=a^{3}c_{\text{dop},\text{c}}$ -
 $\phi$} Electrostatic potential V
@@
 $L$ GB mobility $\frac{\text{m}^{3}}{\text{Js}}$
-$g^{0}_{\text{V}_{\text{O}}^{..},\text{b}}$ Formation energy of oxygen vacancy in the bulk phase.
-$g^{0}_{\text{V}_{\text{O}}^{..},\text{b}}=\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{0}N_{\text{A}}$, with $N_{\text{A}}$ being Avogadro constant $\frac{\text{J}}{\text{mol}}$
-$g_{\text{V}_{\text{O}}^{..},\text{c}}^{0}$ Formation energy of oxygen vacancy in the core region.
-$g^{0}_{\text{V}_{\text{O}}^{..},\text{c}}=\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{0}N_{\text{A}}$ $\frac{\text{J}}{\text{mol}}$
-$g^{0}_{\text{dop},\text{b}}$ Formation energy of dopant in the bulk phase. $g^{0}_{\text{dop},\text{b}}=\mu_{\text{dop},\text{b}}^{0}N_{\text{A}}$ $\frac{\text{J}}{\text{mol}}$
-$g^{0}_{\text{dop},\text{c}}$ Formation energy of dopant in the core region. $g^{0}_{\text{dop},\text{c}}=\mu_{\text{dop},\text{c}}^{0}N_{\text{A}}$ $\frac{\text{J}}{\text{mol}}$
+$g^{0}_{\text{V}_{\text{O}}^{..},\text{b}}$
+Formation energy of oxygen vacancy in the bulk phase.
+$g^{0}_{\text{V}_{\text{O}}^{..},\text{b}}=\mu_{\text{V}_{\text{O}}^{..},\text{b}}^{0}N_{\text{A}}$, with $N_{\text{A}}$ being Avogadro constant
+$\frac{\text{J}}{\text{mol}}$
+$g_{\text{V}_{\text{O}}^{..},\text{c}}^{0}$
+Formation energy of oxygen vacancy in the core region.
+$g^{0}_{\text{V}_{\text{O}}^{..},\text{c}}=\mu_{\text{V}_{\text{O}}^{..},\text{c}}^{0}N_{\text{A}}$
+$\frac{\text{J}}{\text{mol}}$
+$g^{0}_{\text{dop},\text{b}}$
+Formation energy of dopant in the bulk phase.
+$g^{0}_{\text{dop},\text{b}}=\mu_{\text{dop},\text{b}}^{0}N_{\text{A}}$
+$\frac{\text{J}}{\text{mol}}$
+$g^{0}_{\text{dop},\text{c}}$
+Formation energy of dopant in the core region.
+$g^{0}_{\text{dop},\text{c}}=\mu_{\text{dop},\text{c}}^{0}N_{\text{A}}$
+$\frac{\text{J}}{\text{mol}}$
 $V_{\text{m}}$ Molar volume $\frac{\text{m}^{3}}{\text{mol}}$
-$\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}$ Dimensionless available sites of oxygen vacancy in core region.
-$\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}=N_{\text{V}_{\text{O}}^{..},\text{c}}a^{3}$, with $a$ being the lattice constant -
-$\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}$ Dimensionless available sites of oxygen vacancy in bulk phase.
+$\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}$
+Dimensionless available sites of oxygen vacancy in core region.
+$\tilde{N}_{\text{V}_{\text{O}}^{..},\text{c}}=N_{\text{V}_{\text{O}}^{..},\text{c}}a^{3}$, with $a$ being the lattice constant
+$\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}$
+Dimensionless available sites of oxygen vacancy in bulk phase.
 $\tilde{N}_{\text{V}_{\text{O}}^{..},\text{b}}=N_{\text{V}_{\text{O}}^{..},\text{b}}a^{3}$
-$\tilde{N}_{\text{dop},\text{c}}$ Dimensionless available sites of dopant in core region. $\tilde{N}_{\text{dop},\text{c}}=N_{\text{dop},\text{c}}a^{3}$
-$\tilde{N}_{\text{dop},\text{b}}$ Dimensionless available sites of dopant in bulk phase. $\tilde{N}_{\text{dop},\text{b}}=N_{\text{dop},\text{b}}a^{3}$
+$\tilde{N}_{\text{dop},\text{c}}$
+Dimensionless available sites of dopant in core region.
+$\tilde{N}_{\text{dop},\text{c}}=N_{\text{dop},\text{c}}a^{3}$
+$\tilde{N}_{\text{dop},\text{b}}$
+Dimensionless available sites of dopant in bulk phase.
+$\tilde{N}_{\text{dop},\text{b}}=N_{\text{dop},\text{b}}a^{3}$
 $R$ Gas constant $\frac{\text{J}}{\text{K}}$
@@
 Moreover, to solve Eqs.
-(35) and (36), we use the mixed finite element formulation and treat the electro-chemical potentials $\mu_{\text{V}_{\text{O}}^{..}}$ and $\mu_{\text{dop}}$ as additional DOFs [44].
+(36) and (37), we use the mixed finite element formulation and treat the electro-chemical potentials $\mu_{\text{V}_{\text{O}}^{..}}$ and $\mu_{\text{dop}}$ as additional DOFs bai2019two .
 The strong forms of the corresponding governing equations have already been given in Eqs.
 (25), (26) and (34) to (40).
-According to the FEM [45], we formulate the residuals from the weak forms of these governing equations by introducing corresponding test functions $\psi_{C_{\text{V}_{\text{O}}^{..},\text{c}}}$, $\psi_{C_{\text{dop},\text{c}}}$, $\psi_{\eta_{i}}$, $\psi_{C_{\text{V}_{\text{O}}^{..}}}$, $\psi_{C_{\text{dop}}}$, $\psi_{\mu_{\text{V}_{\text{O}}^{..}}}$, $\psi_{\mu_{\text{dop}}}$, $\psi_{C_{\text{V}_{\text{O}}^{..},\text{b}}}$, $\psi_{C_{\text{dop},\text{b}}}$ and $\psi_{\phi}$.
+According to the FEM hughes2012finite , we formulate the residuals from the weak forms of these governing equations by introducing corresponding test functions $\psi_{C_{\text{V}_{\text{O}}^{..},\text{c}}}$, $\psi_{C_{\text{dop},\text{c}}}$, $\psi_{\eta_{i}}$, $\psi_{C_{\text{V}_{\text{O}}^{..}}}$, $\psi_{C_{\text{dop}}}$, $\psi_{\mu_{\text{V}_{\text{O}}^{..}}}$, $\psi_{\mu_{\text{dop}}}$, $\psi_{C_{\text{V}_{\text{O}}^{..},\text{b}}}$, $\psi_{C_{\text{dop},\text{b}}}$ and $\psi_{\phi}$.
 Firstly, the residuals for the weak forms of Eqs.
@@
 $R_{\phi}=\int_{V}\epsilon_{0}\epsilon_{\text{r}}\phi_{,i}\psi_{\phi,i}\text{d}V-\int_{\Gamma}\epsilon_{0}\epsilon_{\text{r}}\phi_{,i}\psi_{\phi}\hat{n}_{i}\text{d}\Gamma-\int_{V}\psi_{\phi}\frac{\mathcal{F}}{V_{\text{m}}}(z_{\text{V}_{\text{O}}^{..}}C_{\text{V}_{\text{O}}^{..}}+z_{\text{dop}}C_{\text{dop}})\text{d}V.$ (54)
-Following the Galerkin approach [46], the test functions are discretized as $\psi_{C_{\text{V}_{\text{O}}^{..},\text{c}}}=N_{C_{\text{V}_{\text{O}}^{..},\text{c}}}^{I}\psi_{C_{\text{V}_{\text{O}}^{..},\text{c}}}^{I}$, $\psi_{C_{\text{dop},\text{c}}}=N_{C_{\text{dop},\text{c}}}^{I}\psi_{C_{\text{dop},\text{c}}}^{I}$, $\psi_{\eta_{i}}=N_{\eta_{i}}^{I}\psi_{\eta_{i}}^{I}$, $\psi_{C_{\text{V}_{\text{O}}^{..}}}=N_{C_{\text{V}_{\text{O}}^{..}}}^{I}\psi_{C_{\text{V}_{\text{O}}^{..}}}^{I}$, $\psi_{C_{\text{dop}}}=N_{C_{\text{dop}}}\psi_{C_{\text{dop}}}^{I}$, $\psi_{\mu_{\text{V}_{\text{O}}^{..}}}=N_{\mu_{\text{V}_{\text{O}}^{..}}}\psi_{\mu_{\text{V}_{\text{O}}^{..}}}^{I}$, $\psi_{\mu_{\text{dop}}}=N_{\mu_{\text{dop}}}\psi_{\mu_{\text{dop}}}^{I}$, $\psi_{C_{\text{V}_{\text{O}}^{..},\text{b}}}=N_{C_{\text{V}_{\text{O}}^{..},\text{b}}}\psi_{C_{\text{V}_{\text{O}}^{..},\text{b}}}^{I}$, $\psi_{C_{\text{dop},\text{b}}}=N_{C_{\text{dop},\text{b}}}\psi_{C_{\text{dop},\text{b}}}^{I}$ and $\psi_{\phi}=N_{\phi}\psi_{\phi}^{I}$, where $I$ denotes the node index and Einstein summation convention is used.
+Following the Galerkin approach brenner2008mathematical , the test functions are discretized as $\psi_{C_{\text{V}_{\text{O}}^{..},\text{c}}}=N_{C_{\text{V}_{\text{O}}^{..},\text{c}}}^{I}\psi_{C_{\text{V}_{\text{O}}^{..},\text{c}}}^{I}$, $\psi_{C_{\text{dop},\text{c}}}=N_{C_{\text{dop},\text{c}}}^{I}\psi_{C_{\text{dop},\text{c}}}^{I}$, $\psi_{\eta_{i}}=N_{\eta_{i}}^{I}\psi_{\eta_{i}}^{I}$, $\psi_{C_{\text{V}_{\text{O}}^{..}}}=N_{C_{\text{V}_{\text{O}}^{..}}}^{I}\psi_{C_{\text{V}_{\text{O}}^{..}}}^{I}$, $\psi_{C_{\text{dop}}}=N_{C_{\text{dop}}}\psi_{C_{\text{dop}}}^{I}$, $\psi_{\mu_{\text{V}_{\text{O}}^{..}}}=N_{\mu_{\text{V}_{\text{O}}^{..}}}\psi_{\mu_{\text{V}_{\text{O}}^{..}}}^{I}$, $\psi_{\mu_{\text{dop}}}=N_{\mu_{\text{dop}}}\psi_{\mu_{\text{dop}}}^{I}$, $\psi_{C_{\text{V}_{\text{O}}^{..},\text{b}}}=N_{C_{\text{V}_{\text{O}}^{..},\text{b}}}\psi_{C_{\text{V}_{\text{O}}^{..},\text{b}}}^{I}$, $\psi_{C_{\text{dop},\text{b}}}=N_{C_{\text{dop},\text{b}}}\psi_{C_{\text{dop},\text{b}}}^{I}$ and $\psi_{\phi}=N_{\phi}\psi_{\phi}^{I}$, where $I$ denotes the node index and Einstein summation convention is used.
 $\psi_{C_{\text{V}_{\text{O}}^{..},\text{c}}}^{I}$, $\psi_{C_{\text{dop},\text{c}}}^{I}$, $\psi_{\eta_{i}}^{I}$, $\psi_{C_{\text{V}_{\text{O}}^{..}}}^{I}$, $\psi_{C_{\text{dop}}}^{I}$, $\psi_{\mu_{\text{V}_{\text{O}}^{..}}}^{I}$, $\psi_{\mu_{\text{dop}}}^{I}$, $\psi_{C_{\text{V}_{\text{O}}^{..},\text{b}}}^{I}$, $\psi_{C_{\text{dop},\text{b}}}^{I}$ and $\psi_{\phi}^{I}$ are the nodal weights, while $N_{C_{\text{V}_{\text{O}}^{..},\text{c}}}^{I}$, $N_{C_{\text{dop},\text{c}}}^{I}$, $N_{\eta_{i}}^{I}$, $N_{C_{\text{V}_{\text{O}}^{..}}}^{I}$, $N_{C_{\text{dop}}}^{I}$, $N_{\mu_{\text{V}_{\text{O}}^{..}}}^{I}$, $N_{\mu_{\text{dop}}}^{I}$, $N_{C_{\text{V}_{\text{O}}^{..},\text{b}}}^{I}$, $N_{C_{\text{dop},\text{b}}}^{I}$ and $N_{\phi}^{I}$ are the shape functions for the corresponding variables.
@@
 $R_{C_{\text{dop},\text{c}}}^{I}=\int_{V}N_{C_{\text{dop},\text{c}}}^{I}\{C_{\text{dop}}-[1-h(\eta)]C_{\text{dop},\text{b}}-h(\eta)C_{\text{dop},\text{c}}\}\text{d}V$ (56)
-$R_{\eta_{i}}^{I}&=\int_{V}\frac{N_{\eta_{i}}^{I}}{\Delta t}\text{d}V+L\int_{V}N_{\eta_{i}}^{I}\omega\frac{\partial f^{\text{loc}}(\eta)}{\partial\eta_{i}}\text{d}V+L\int_{V}\kappa\eta_{i,j}N_{\eta_{i},j}^{I}\text{d}V-L\int_{\Gamma}\kappa\eta_{i,j}N_{\eta_{i}}^{I}\hat{n}_{j}\text{d}\Gamma\\ &-L\int_{V}N_{\eta_{i}}^{I}\frac{\partial h(\eta)}{\partial\eta_{i}}\left[f_{\text{b}}^{\text{ech}}-f_{\text{c}}^{\text{ech}}-\frac{\partial f_{\text{b}}^{\text{ech}}}{\partial C_{\text{V}_{\text{O}}^{..},\text{b}}}(C_{\text{V}_{\text{O}}^{..},\text{b}}-C_{\text{V}_{\text{O}}^{..},\text{c}})-\frac{\partial f_{\text{b}}^{\text{ech}}}{\partial C_{\text{dop},\text{b}}}(C_{\text{dop},\text{b}}-C_{\text{dop},\text{c}})\right]\text{d}V\end{aligned}$ (52)
+$\begin{aligned} R_{\eta_{i}}^{I}&=\int_{V}\frac{N_{\eta_{i}}^{I}}{\Delta t}\text{d}V+L\int_{V}N_{\eta_{i}}^{I}\omega\frac{\partial f^{\text{loc}}(\eta)}{\partial\eta_{i}}\text{d}V+L\int_{V}\kappa\eta_{i,j}N_{\eta_{i},j}^{I}\text{d}V-L\int_{\Gamma}\kappa\eta_{i,j}N_{\eta_{i}}^{I}\hat{n}_{j}\text{d}\Gamma\\ &-L\int_{V}N_{\eta_{i}}^{I}\frac{\partial h(\eta)}{\partial\eta_{i}}\left[f_{\text{b}}^{\text{ech}}-f_{\text{c}}^{\text{ech}}-\frac{\partial f_{\text{b}}^{\text{ech}}}{\partial C_{\text{V}_{\text{O}}^{..},\text{b}}}(C_{\text{V}_{\text{O}}^{..},\text{b}}-C_{\text{V}_{\text{O}}^{..},\text{c}})-\frac{\partial f_{\text{b}}^{\text{ech}}}{\partial C_{\text{dop},\text{b}}}(C_{\text{dop},\text{b}}-C_{\text{dop},\text{c}})\right]\text{d}V\end{aligned}$ (57)
 $R_{C_{\text{V}_{\text{O}}^{..}}}^{I}=\int_{V}N_{C_{\text{V}_{\text{O}}^{..}}}^{I}\left(\mu_{\text{V}_{\text{O}}^{..}}-\frac{\delta\mathscr{F}}{\delta C_{\text{V}_{\text{O}}^{..},\text{b}}}\right)\text{d}V$ (58)
@@
 The non-zero terms in the tangent matrix are presented in supplementary.
-The numerical implementation of the proposed phase-field model is carried out in Multiphysics Object-Oriented Simulation Environment (MOOSE) framework [47, 48].
+The numerical implementation of the proposed phase-field model is carried out in Multiphysics Object-Oriented Simulation Environment (MOOSE) framework schwen2023phasefield , giudicelli2024moose .
+Spatial discretization was performed with linear elements for one-dimensional (1D) simulations (Section 5.1 and 5.2) and quadrilatera elements for two-dimensional (2D) simulations (Section 5.3), with linear Lagrange shape functions in each case aagesen2024electrochemical .
+The same basis functions were used for both trial (shape) and test functions, which is standard in the Galerkin finite element method brenner2008mathematical .
+Different mesh resolutions were adopted based on the specific requirements of each case.
+To ensure numerical accuracy, the relative convergence tolerance is set to $1\times 10^{-8}$.
+All phase-field simulations were conducted on CPU nodes equipped with Intel Xeon Platinum 8470Q processors and were accelerated using the Message Passing Interface (MPI) on the Lichtenberg high-performance computing cluster at Technische Universität Darmstadt.
 5 Phase-field simulation results of Fe-doped SrTiO3
-We present first in Section 5.1 one-dimensional phase-field simulation results to demonstrate its capability of reproducing the formation of a SCL at equilibrium state.
+We present first in Section 5.1 one-dimensional (1D) phase-field simulation results to reproduce the formation of a SCL at equilibrium state under the influence of different defect chemistry parameters.
 Subsequently, we investigate the influence of solute drag effects induced by the segregation of point defects in the core on the formation of an asymmetric SCL under quasi-equilibrium state in Section 5.2.
 Finally, we conduct two-dimensional (2D) phase-field simulations to study the microstructure evolution process in STO polycrystalline material during sintering and analyze the resulting grain size distribution in Section 5.3.
-Note that in the paper we ignore the GB orientation dependency or other anisotropy of the segregation energies $\Delta g_{\text{V}_{\text{O}}^{..}}$ and $\Delta g_{\text{dop}}$, for we intend to reveal the most fundamental mechanisms of SCL formation and abnormal grain growth.
+Note that in the paper we ignore the GB orientation dependency or other anisotropy of the segregation energies $\Delta g_{\text{V}_{\text{O}}^{..}}$ and $\Delta g_{\text{dop}}$, for we intend to reveal the most fundamental mechanisms of SCL formation and the grain growth patterns.
 Without loss of generality, the proposed phase-field model can be extended accordingly to include such anisotropy aspects.
 5.1 GB space charge layers at equilibrium
-To demonstrate the capability of the phase-field model at thermodynamically equilibrium state, one-dimensional phase-field simulations are conducted and compared with the sharp interface predictions [30].
-The semi-finite simulation domain has the size of 780 nm, with the core located at the left boundary.
-The grain boundary thickness is prescribed to be 0.78 nm [30].
-The domain is discretized into a mesh consisting of 2000 elements.
-The simulation temperature is set at 600 K, and the relative permittivity $\epsilon_{\text{r}}$ of strontium titanate (STO) is specified as 160, according to [30], which is calculated via $\epsilon_{\text{r}}=9\times 10^{4}/(T-35)$ as mentioned in section 3.1.
-The interface energy, molar volume and interface mobility of STO are 0.6 $\text{J}\text{m}^{-2}$, 3.6×$\times 10^{-5}$ $\text{m}^{3}\text{mol}^{-1}$ and 1×$\times 10^{-7}$ $\text{m}^{3}\text{J}^{-1}\text{s}^{-1}$, respectively [38].
-The relative tolerance for convergence is set to $1\times 10^{-8}$.
-5.1.1 Benchmark case 1: The MS Model
-By fixing a uniform dopant concentration, the proposed phase-field model can be used to simulate the MS case.
-The results are given in Fig.
+To examine the influence of various defect parameters on SCL formation, 1D phase-field simulations are conducted for a bicrystal configuration at temperatures of $T=600$ K and $T=1623$ K.
+At 600 K, the distribution of acceptor dopant is frozen, while that of the oxygen vacancy achieves chemical equilibrium (MS model).
+At 1623 K, the diffusion of acceptor dopant becomes active, enabling both the acceptor dopant and oxygen vacancy to achieve electrochemical equilibrium, referred to as the GC model.
+Furthermore, we chose the physical GB width (0.78 nm de2009formation ) in the phase-field simulations.
+This choice can decrease the abnormal interface effects raised by an artificially enlarged interface thickness wang2020modeling , wang2021quantitative .
+The defect chemistry parameters used for the phase-field simulations of SCLs are summarized in Table 3.
+We summarize the numerical settings employed in the FEM implementation.
+The simulation domain is one-dimensional and symmetric, spanning from x=390x=-390 nm to $x=390$ nm, with the GB located at $x=0$ nm.
+The domain is discretized into 10,000 elements with a uniform element size of 0.078 nm.
+For the electrostatic potential, a Neumann boundary condition is applied at the right boundary, while the left boundary is grounded.
+Neumann boundary conditions are also imposed on the concentration field at both boundaries.
+Each simulation is executed on a single CPU node with 16 cores and 8 gigabyte of memory, requiring approximately 5 hours of computation time.
+Table 3: The defect chemistry parameters for the 1D phase-field simulations in MS and GC models to reproduce equilibrium SCLs. The definition of these parameters are described in Table 1.
+Parameters Value in MS model de2009formation Value in GC model de2009formation
+TT/K 600 1623
+aa/nm 0.39 0.39
+$\omega_{c}$/nm 0.78 0.78
+$\epsilon_{r}$ 160 56
+$c_{\text{dop},\text{b}}(\infty)$/ $\text{m}^{-3}$ $7\times 10^{25}$ $7\times 10^{25}$
+$c_{\text{V}_{\text{O}}^{..},\text{b}}(\infty)$/ $\text{m}^{-3}$ $3.5\times 10^{25}$ $3.5\times 10^{25}$
+$l_{\text{D}}$/nm 1.81 1.76
+$\Delta g_{\text{V}_{\text{O}}^{..}}$/eV -1.5 $\sim$ 0 -1.5
+$\Delta g_{\text{dop}}$/eV - -1.0 $\sim$ 0.5
+$N_{\text{V}_{\text{O}}^{..},\text{b}}$/ $\text{m}^{-3}$ $5.1\times 10^{28}$ $5.1\times 10^{28}$
+$N_{\text{dop},\text{b}}$/ $\text{m}^{-3}$ - $1.68\times 10^{28}$
+$N_{\text{V}_{\text{O}}^{..},\text{c}}$/ $\text{m}^{-3}$ $5\times 10^{26}$ $\sim$ $1\times 10^{27}$ $5\times 10^{26}$ $\sim$ $1\times 10^{27}$
+$N_{\text{dop},\text{c}}$/ $\text{m}^{-3}$ - $1.68\times 10^{26}$ $\sim$ $1.68\times 10^{27}$ de2009formation
+5.1.1 Benchmark case 1: MS Model
+By prescribing a uniform dopant concentration, the proposed phase-field model is employed to simulate the MS case.
+Phase-field simulations of SCLs with varying available number densities in the MS model are systematically benchmarked against the sharp interface (SI) calculations of GB||SCL configurations, as described in Section 2.
+The results of this comparison are presented in Fig.
 2.
-The initial concentrations of oxygen vacancies and dopant far from the core (denoted as $c_{\text{dop},\text{b}}(\infty)$ and $c_{\text{V}_{\text{O}}^{..},\text{b}}(\infty)$, respectively) are $7\times 10^{24}\text{ m}^{-3}$ and $3.5\times 10^{24}\text{ m}^{-3}$ [30].
-The number of available sites of oxygen vacancy in bulk phase and GB core are $N_{\text{V}_{\text{O}}^{..},\text{b}}=5.12\times 10^{28}$ $\text{m}^{-3}$ and $N_{\text{V}_{\text{O}}^{..},\text{c}}=3\times 10^{26}$ $\text{m}^{-3}$, respectively[30, 49].
-The grain boundary potential $\Phi_{0}$ is plotted as a function of $\Delta g_{\text{V}_{\text{O}}^{..}}$ in Fig.
-2(a).
-Due to the right boundary being grounded, the grain boundary potential is directly obtained at the interface position when $\eta=0.5$, as defined in the phase-field method.
-For $\Delta g_{\text{V}_{\text{O}}^{..}}$ larger than -1.25 eV, the phase-field results exhibit good agreement with analytical calculations.
-As $\Delta g_{\text{V}_{\text{O}}^{..}}$ becomes more negative, particularly at -1.5 eV and -2.0 eV, the available oxygen vacancy sites are completely exhausted.
-Therefore, the concentrations of oxygen vacancies in the core are close to $N_{\text{V}_{\text{O}}^{..},\text{c}}$.
-Although the phase-field results demonstrate convergent behavior, a discrepancy between the phase-field results and the analytical solution emerges.
-To explain the aforementioned discrepancy, we plot the concentration of oxygen vacancies in the core as a function of $\Delta g_{\text{V}_{\text{O}}^{..}}$ in Fig.
+The phase-field results are represented by solid lines, while the SI calculations are denoted by cross symbols.
+The orange, blue, and purple colors correspond to different available site densities: orange for $N_{\text{V}_{\text{O}}^{..},\text{c}}=1\times 10^{27}\text{m}^{-3}$, blue for $N_{\text{V}_{\text{O}}^{..},\text{c}}=7.5\times 10^{26}\text{m}^{-3}$, and purple for $N_{\text{V}_{\text{O}}^{..},\text{c}}=5\times 10^{26}\text{m}^{-3}$.
+Fig.
+2(a) and (b) illustrate the GB potentials ($\Phi_{0}$) and oxygen vacancy concentrations at the GB core ($c_{\text{V}_{\text{O}}^{..},\text{c}}$) for various segregation energies and available oxygen vacancy number densities.
+It should be clarified that $\Phi_{0}$ and $c_{\text{V}_{\text{O}}^{..},\text{c}}$ are directly obtained at the interface position, defined by $\eta=0.5$ in the phase-field method.
+An increase in $|\Delta g_{\text{V}_{\text{O}}^{..}}|$ ($\Delta g_{\text{V}_{\text{O}}^{..}}<0$) promotes greater oxygen vacancy segregation at the GB core, thereby enhancing the GB potential.
+However, the GB potential does not increase indefinitely.
+At sufficiently large $|\Delta g_{\text{V}_{\text{O}}^{..}}|$, when all available sites for oxygen vacancies are fully occupied, the GB potential reaches a plateau, as does the oxygen vacancy concentration in the GB core ($c_{\text{V}_{\text{O}}^{..},\text{c}}\approx N_{\text{V}_{\text{O}}^{..},\text{c}}$, see Fig.
+2(b)).
+The available number density of oxygen vacancy sites ($N_{\text{V}_{\text{O}}^{..},\text{c}}$) plays a critical role in determining the plateau value.
+For instance, when $N_{\text{V}_{\text{O}}^{..},\text{c}}$ increases from $5\times 10^{26}\text{ m}^{-3}$ to $7.5\times 10^{26}\text{ m}^{-3}$, the plateau GB potential rises from 0.12 V to 0.24 V in the phase-field results and from 0.13 V to 0.27 V in the sharp interface (SI) calculations.
+However, when $N_{\text{V}_{\text{O}}^{..},\text{c}}$ is as high as $1\times 10^{27}\text{ m}^{-3}$, this plateau behavior becomes less pronounced, as no clear convergence of the GB potential is observed in both the SI calculations and phase-field results.
+Although phase-field results exhibit the same trend as predicted by SI calculations, discrepancies in GB potentials are still observed in the large $|\Delta g_{\text{V}_{\text{O}}^{..}}|$ region.
+To further investigate the GB core, we plot the area density of charge in the GB core, $Q_{\text{c}}=\int_{-w_{c}/2}^{w_{c}/2}e(z_{\text{dop}}c_{\text{dop},\text{c}}+z_{\text{V}_{\text{O}}^{..}}c_{\text{V}_{\text{O}}^{..},\text{c}})\text{d}x$, obtained from phase-field results and SI calculations, as a function of $\Delta g_{\text{V}_{\text{O}}^{..}}$ for different $N_{\text{V}_{\text{O}}^{..},\text{c}}$ in Fig.
+2(c).
+Here, $Q_{\text{c}}$ represents the boundary condition in the SI calculations, as defined in Eq.
+17, which plays a critical role in determining the GB potential and defect concentration in GB core.
+When $|\Delta g_{\text{V}_{\text{O}}^{..}}|$ is low, the phase-field results align well with the SI calculations.
+However, discrepancies emerge in the high $|\Delta g_{\text{V}_{\text{O}}^{..}}|$ region, where phase-field simulations reproduce lower area charge densities.
+These discrepancies are likely due to the diffuse interface in phase-field simulations.
+In the SI model, the defect concentrations between the GB core and the bulk phase are discontinuous, with defect concentrations assumed to be constant across the GB core.
+Conversely, the smooth interface characteristics inherent to phase-field simulations result in a less positively charged GB core compared to the SI calculations when the same GB width, $w_{c}=0.78$ nm, is employed.
+In addition to the GB core, we also compare the distributions of electrostatic potentials and oxygen vacancy concentrations in the bulk phase for different $\Delta g_{\text{V}_{\text{O}}^{..}}$ values, as shown in Fig.
+2(d) to (f).
+Due to the symmetric bicrystal configuration, only half of the interface region is included in the analysis.
+In Fig.
+2 (d1), (e1), and (f1), depletion zones of oxygen vacancy are observed near the GB core.
+Larger $N_{\text{V}_{\text{O}}^{..},\text{c}}$ values result in more pronounced depletion of oxygen vacancy.
+Overall, the phase-field results show good agreement with the SI calculations.
+n Fig.
+2(d1), (e1), and (f1), depletion zones of oxygen vacancy are observed near the GB core.
+Larger $N_{\text{V}_{\text{O}}^{..},\text{c}}$ values lead to more pronounced depletion of oxygen vacancy.
+Overall, the phase-field results align well with the SI calculations.
+However, discrepancies are observed near the GB core, where oxygen vacancy concentrations decrease in the SI calculations but increase in the phase-field results (see Fig.
+2(e1) and (e2) in the range of [0, 2 nm]).
+These differences stem from the contrasting assumptions used to connect the GB core and bulk phase in the phase-field model and the discontinuous SI model, as discussed earlier.
+In Fig.
+2 (d2), (e2), and (f2), larger $N_{\text{V}_{\text{O}}^{..},\text{c}}$ valules lead to higher electrostatic potential and longer electrostatic decay lengths.
+The phase-field simulations is capable of capturing these trends and nicely reproducing the SI predictions.
+Figure 2: (a) Comparison of phase-field (PF) simulation outcomes with those of the abrupt GB||SCL sharp interface (SI) predictions with different available number density of oxygen vacancy, such as $N_{\text{V}_{\text{O}}^{..},\text{b}}=5\times 10^{26}\text{m}^{-3}$, $7.5\times 10^{26}\text{m}^{-3}$ and $1\times 10^{27}\text{m}^{-3}$.
+Figure (a), (b) and (c) illustrates the GB potential, GB concentration of oxygen vacancy and area density of charge as a function of $\Delta g_{\text{V}_{\text{O}}^{..}}$.
+Figure (d), (e) and (f) depicts the spatial distribution of oxygen vacancy concentration and electrostatic potential in bulk phase.
+5.1.2 Benchmark case 2: GC Model
+By fixing the electrochemical potentials of the acceptor dopant and oxygen vacancies as constants, the proposed phase-field model reduces to the GC model, enabling the reproduction of equilibrium SCL formation in the GC case.
+In this framework, the effects of the segregation energy of the acceptor dopant ($\Delta g_{\text{dop}}$) and the available site densities of oxygen vacancies and acceptor dopants in the GB core ($N_{\text{dop},\text{c}}$ and $N_{\text{V}_{\text{O}}^{..},\text{c}}$) on SCL formation are investigated and compared with SI calculations, as illustrated in Fig.
+3.
+In both phase-field and SI results, the segregation energy of oxygen vacancy $\Delta g_{\text{V}_{\text{O}}^{..}}$ is set to 1.5 eV.
+For $N_{\text{V}_{\text{O}}^{..},\text{c}}=1\times 10^{27}\text{m}^{-3}$, phase-field results and SI predictions are shown using solid lines and cross symbols, respectively.
+For $N_{\text{V}_{\text{O}}^{..},\text{c}}=5\times 10^{26}\text{m}^{-3}$, dashed lines and triangle symbols are used.
+Additionally, different colors are employed to represent varying $N_{\text{dop},\text{c}}$: red for $N_{\text{dop},\text{c}}=1.68\times 10^{26}\text{m}^{-3}$, blue for $N_{\text{dop},\text{c}}=8.4\times 10^{26}\text{m}^{-3}$, and orange for $N_{\text{dop},\text{c}}=1.68\times 10^{27}\text{m}^{-3}$.
+In Fig.
+3(a), (b), and (c), the GB potential ($\Phi_{0}$), GB concentrations of oxygen vacancy ($c_{\text{V}_{\text{O}}^{..},\text{c}}$), and acceptor dopant ($c_{\text{dop},\text{c}}$) are presented as functions of $\Delta g_{\text{dop}}$.
+In Fig.
+3(a), the GB potential curves exhibit two distinct plateaus.
+For positive $\Delta g_{\text{dop}}$, the GB potential reaches an upper plateau due to minimal dopant segregation in the GB core.
+In this case, the dopant concentration in the GB core can be 2 to 3 orders of magnitude lower than in the bulk phase (see Fig.
+3(b) at positive $\Delta g_{\text{dop}}$ region), making it insufficient to compensate for the charge of the GB core induced by oxygen vacancy.
+As $\Delta g_{\text{dop}}$ becomes negative, the positively charged GB core is gradually compensated by the segregating dopant (see Fig.
+3(b) at negative $\Delta g_{\text{dop}}$ region), leading to a reduction in GB potential.
+When all available sites for acceptor dopant are fully occupied, the GB potential reaches its lower plateau.
+Increasing $N_{\text{V}_{\text{O}}^{..},\text{c}}$ raises both the upper and lower plateaus simultaneously, with little impact on the shape of the GB potential curve.
+The difference between the upper and lower plateaus is primarily determined by $N_{\text{dop},\text{c}}$, with larger $N_{\text{dop},\text{c}}$ resulting in a greater plateau difference.
+For example, when $N_{\text{V}_{\text{O}}^{..},\text{c}}=5\times 10^{26}\text{m}^{-3}$, the plateau difference is 0.17 V for $N_{\text{dop},\text{c}}=1.68\times 10^{26}\text{m}^{-3}$ and increases to 1.5 V for $N_{\text{dop},\text{c}}=1.68\times 10^{27}\text{m}^{-3}$.
+In addition, very large $N_{\text{dop},\text{c}}$ might lead to negative GB potential as presented by the orange dashed line and triangle symbols when $\Delta g_{\text{dop}}$ is smaller than -0.8 eV in Fig.
+3(a).
+Adjusting $N_{\text{V}_{\text{O}}^{..},\text{c}}$ and $N_{\text{dop},\text{c}}$ has a pronounced effect on the GB concentrations of oxygen vacancy and acceptor dopant.
+When the available sites are not fully occupied, increasing $N_{\text{V}_{\text{O}}^{..},\text{c}}$ and $N_{\text{dop},\text{c}}$ results in greater dopant accumulation within the GB core for a given $\Delta g_{\text{dop}}$, as shown in Fig.
 2(b).
-Similar to the grain boundary potential, $c_{\text{V}_{\text{O}}^{..}}$ is also extracted at $\eta=0.5$.
-We observe that the phase-field results coincide with the analytical predictions, even when $\Delta g_{\text{V}_{\text{O}}^{..}}=-2.0$ eV.
-Next, we examine the concentration of oxygen vacancy in the bulk in terms of $x$.
-Here, we plot the phase-field results for $\Delta g_{\text{V}_{\text{O}}^{..}}$=-0.5 eV, -1.0 eV and -1.5 eV in Fig.
-2 (c), (d) and (e) in the range of 0 to 50 nm, which includes half of the interface region.
-For the analytical solution, we shift it to the right by a distance of $w_{\text{c}}/2$ to match the bulk phase in the phase-field results.
-When ${\Delta g_{\text{V}_{\text{O}}^{..}}}=-0.5$ eV, $c_{\text{V}_{\text{O}}^{..},\text{b}}$ obtained by phase-field simulations agrees well with analytical predictions as shown in Fig.
-2(c1).
-However, in Fig.
-2(d1) and (e1), for $\Delta g_{\text{V}_{\text{O}}^{..}}=-1.5$ and $-2.0$ eV, the phase-field results produce higher $c_{\text{V}_{\text{O}}^{..},\text{b}}$.
-This comparison indicates that less oxygen vacancy is segregated in the core in the phase-field simulations when choosing more negative $\Delta g_{\text{V}_{\text{O}}^{..}}$.
-Additionally, we plot the area density of the net charge, $Q_{\mathrm{net}}=\int_{-w_{\text{c}}/2}^{w_{\text{c}}/2}(z{{}_{\text{dop}}}c_{{}_{\text{dop},\text{c}}}+z_{{}_{\text{V}_{\text{O}}^{..}}}c_{{}_{\text{V}_{\text{O}}^{..},\text{c}}})\text{d}x$, in the core in Fig.
-2(f).
-Less net charge is reproduced via phase-field simulation when $\Delta g_{\text{V}_{\text{O}}^{..}}$ is smaller.
-In short, this discrepancy is attributed to the diffuse interface in the phase-field simulations.
-Due to its finite interface thickness and smooth interface characteristics, the core is less positively charged compared to analytical solutions.
-Consequently, the electrostatic potential profile obtained via phase-field simulations decays faster than analytical calculations as shown in Fig.
-2(c2), (d2) and (e2).
-Up to this point, we clearly observe a good agreement between the oxygen vacancy concentrations in the GB core and bulk phase as predicted by the phase-field results and the sharp interface predictions.
-The different approaches to handling interfaces in the phase-field model and the sharp interface method can result in discrepancies when the available densities of point defects are fully occupied.
-Figure 2: Comparison of phase-field (PF) simulation outcomes with those of the sharp interface (SI) model predictions obtained via FEM and taken from Ref.
-[30] in the MS model.
-Figure (a) illustrates the grain boundary potential as a function of $\Delta g_{\text{V}_{\text{O}}^{..}}$.
-Figure (b) depicts the relationship between the oxygen vacancy concentration and $\Delta g_{\text{V}_{\text{O}}^{..}}$.
-Figures (c), (d), and (e) present the concentration distribution of dopant and the electrostatic potential distribution for different oxygen vacancy segregation energy.
-Figure (f) demonstrates the net area density of charge in the core as a function of $\Delta g_{\text{V}_{\text{O}}^{..}}$.
-5.1.2 Benchmark case 2: The GC Model
-By setting the electrochemical potentials of the acceptor dopant and of the oxygen vacancies to be constant, the presented phase-field model is then reduced to the GC model, in order to reproduce the equilibrium SCL formation in the GC case.
-The number of available sites of dopant and oxygen vacancy in bulk and GB core are $N_{\text{V}_{\text{O}}^{..},\text{c}}=1\times 10^{27}$ $\text{m}^{-3}$, $N_{\text{V}_{\text{O}}^{..},\text{b}}=5.12\times 10^{28}$ $\text{m}^{-3}$, $N_{\text{dop},\text{c}}=1.68\times 10^{27}$ $\text{m}^{-3}$ and $N_{\text{dop},\text{b}}=1.69\times 10^{28}$ $\text{m}^{-3}$, respectively [30, 49].
-The Debye lengths $l_{\text{D}}=\sqrt{\frac{\epsilon_{0}\epsilon_{\text{r}}k_{\text{B}}T}{2e^{2}c_{\text{dop},\text{b}}(\infty)}}$ in both MS and GC models are 5.71 nm.
+The segregation of dopant also influences the distribution of oxygen vacancy.
 In Fig.
-3(a), the grain boundary potentials obtained via the phase-field simulations are compared with the sharp interface predictions.
-$\Delta g_{\text{V}_{\text{O}}^{..}}$ is fixed at -1.0 eV here.
-When $\Delta g_{\text{dop}}$ is positive, dopant cations are rejected from the core, thereby causing a more pronounced grain boundary potential.
-Conversely, as $\Delta g_{\text{dop}}$ becomes negative, dopant segregation to the core occurs, reducing the net charge and consequently decreasing the grain boundary potential.
-Moreover, as $\Delta g_{\text{dop}}$ becomes sufficiently negative, the available number sites for dopant segregating within the core become scarce.
-Therefore, the grain boundary potential profile is flat when $\Delta g_{\text{dop}}$ is positive or sufficiently negative.
-The comparison indicates the phase-field results agree with the predictions of analytical solution.
-The discrepancy is smaller than 7%.
-The concentrations of vacancies and acceptor dopants in the core as a function of $\Delta g_{\text{dop}}$ are plotted in Fig.
-3(b).
+2(c), a more negative $\Delta g_{\text{dop}}$ leads to a higher GB concentration of oxygen vacancy.
+Similar to the GB potential curves, the oxygen vacancy concentration curves exhibit two plateaus.
+An increase in $N_{\text{dop},\text{c}}$ amplifies the difference between these plateaus.
+We also present the distributions of electrostatic potential, oxygen vacancy concentration, and acceptor dopant concentration in the bulk phase for different $N_{\text{dop},\text{c}}$ and $\Delta g_{\text{dop}}$ in Fig.
+3(d), (e), and (f).
+Here, $N_{\text{V}_{\text{O}}^{..},\text{c}}$ is fixed at $5\times 10^{26}\text{m}^{-3}$.
+First, we observe a depletion zone for oxygen vacancy and an accumulation zone for acceptor dopant in the SCLs when the GB potential is positive, while the opposite behavior is observed for negative GB potential (see the orange dashed lines and triangle symbols in Fig.
+3(f1), (f2), and (f3)).
+For $\Delta g_{\text{dop}}=0$ eV, increasing $N_{\text{dop},\text{c}}$ has minimal impact on the distributions of oxygen vacancy concentration, acceptor dopant concentration, and electrostatic potential in the bulk phase.
+As a result, the curves for $N_{\text{dop},\text{c}}=1.68\times 10^{26}\text{m}^{-3}$, $8.4\times 10^{26}\text{m}^{-3}$, and $1.68\times 10^{27}\text{m}^{-3}$ are nearly indistinguishable, as shown in Fig.
+3(d1), (d2), and (d3).
+For $\Delta g_{\text{dop}}=-0.5$ eV and 1.0-1.0 eV, increasing $N_{\text{dop},\text{c}}$ raises the oxygen vacancy concentration in the depletion zone, reduces the acceptor dopant concentration in the accumulation zone, and decreases the electrostatic potential in SCLs.
+However, we still observe discrepancies between phase-field results and SI predictions near the GB core, which has been clarified in the MS case.
+We compare the equilibrium SCL formation obtained from phase-field simulations with the results of Vikrant’s simulations for Fe-doped SrTiO3 (see Figure 1 in Ref.
+vikrant2020electrochemical ).
+In Vikrant’s study, a higher dopant concentration is used, i.e.
+2% (in present work, dopant concentration is 0.42%), which is considered as non-dilute case and different grain boundary misorientation is considered.
+Their simulation temperature is set at 1623 K, and the acceptor dopant is considered mobile.
+They employed a negative segregation energy for oxygen vacancy (0.8-0.8 eV) and a zero segregation energy for dopant.
+The major differences between the two studies are as follows: (i) In Vikrant’s results, the dopant concentration in the GB core is consistently higher than in the bulk phase for all GB misorientations.
+(ii) Their simulations also reveal a depletion zone for acceptor dopant near the GB core within the SCL.
+In the present work, a zero segregation energy for dopant results in significantly lower dopant concentrations in the GB core.
+As shown in Fig.
+3(b), the dopant concentration in the GB core, $c_{\text{dop},\text{c}}$, is consistently smaller than the bulk concentration, $c_{\text{dop},\text{b}}(\infty)=7\times 10^{25}$ $\text{m}^{-3}$, for all cases where $\Delta g_{\text{dop}}=0$ eV.
+Furthermore, we observe an accumulation zone for acceptor dopant in the SCL region under positive electrostatic potentials, as depicted in Fig.
+3(d2), (e2) and (f2).
+We note that defect–defect interactions were included in Vikrant’s work and can influence space charge layer formation, as discussed in Refs.
+lund2021thermodynamically , mebane2015generalised.
+However, their model assumes that oxygen vacancies and dopants share the same sublattice, which affects the configurational entropy and consequently the electrochemical potential (see Section 3.3), particularly in the non-dilute regime where these quantities are sensitive to the underlying sublattice occupancy assumptions.
+Thus far, we have successfully benchmarked the capabilities of the proposed phase-field model against the abrupt GB||SCL model in both the MS and GC cases.
+By comprehensively accounting for the segregation energies of various defects and partitioning the available site densities among different defects, diverse SCL formation behaviors have been accurately reproduced.
+As a result, the phase-field results show excellent agreement with SI predictions, not only within the GB core but also in the bulk phase.
+However, due to the assumptions of discontinuous GB and SCL in the SI model, two key discrepancies are observed.
+First, the GB potential in the GB core obtained from the phase-field model is smaller than that predicted by the SI model, particularly when the segregation energy is sufficiently large.
+Second, differences in defect concentrations are confined to the immediate vicinity of the GB core.
+Figure 3: Comparison of phase-field (PF) simulation results with the abrupt GB||SCL sharp interface (SI) results via FEM method for the GC case.
+Figures (a), (b) and (c) showcase the variations in grain boundary potential, dopant concentration, and oxygen vacancy concentration in the core as functions of $\Delta g_{\text{dop}}$.
+Figures (d), (e), and (f) detail the spatial distributions of dopant and oxygen vacancy concentrations in the bulk phase and the electrostatic potential.
+$\Delta g_{\text{V}_{\text{O}}^{..}}$ was kept at -1.5 eV for all calculations.
+5.2 GB space charge layers at quasi-equilibrium
+In this section, we systematically investigate the solute drag effect induced by an asymmetric SCL at a quasi-statically moving GB in a STO bicrystal case.
+This analysis is conducted using the defect-chemistry-consistent phase-field model, which has been validated against sharp interface predictions in previous section.
+While the impacts of GB misorientation have been extensively studied in Ref.
+vikrant2020electrochemical , our focus here is on the influence of defect chemistry parameters, such as the segregation energies of oxygen vacancy and dopant, on SCL formation and solute drag effects at the quasi-equilibrium state in a bicrystal case.
+The phase-field simulation parameters to reproduce the SCLs at quasi-equilibrium state are listed in Table 4.
+For FEM implementation, the simulation domain is a 1D space extending from x=780x=-780 nm to $x=780$ nm, with the GB core initially positioned at x=390x=-390 nm ($t=0$).
+The domain size is chosen to be sufficiently large to ensure the movement of SCL is not hindered by the boundary of simulation box.
+The domain is discretized into 20,000 elements, each with a uniform size of 0.078 nm.
+For the electrostatic potential, a Neumann boundary condition is applied at the right boundary, while the left boundary is grounded.
+Similarly, Neumann boundary conditions are imposed on the concentration field at both boundaries.
+Each simulation is performed on a single CPU node with 16 cores and 8 GB of memory, requiring approximately 12 hours to reach the quasi-equilibrium state.
+Table 4: The defect chemistry parameters for the 1D phase-field simulations to reproduce quasi-equilibrium SCLs.
+Parameters Value
+TT/K 1623
+$D_{\text{V}_{\text{O}}^{..}}$$\text{m}^{-2}\text{s}^{-1}$ $1.35\times 10^{-9}$
+$D_{\text{dop}}$$\text{m}^{-2}\text{s}^{-1}$ $1.3\times 10^{-11}$
+aa/nm 0.39
+$\omega_{c}$/nm 0.78
+$\epsilon_{r}$ 56
+$c_{\text{dop},\text{b}}(\infty)$/ $\text{m}^{-3}$ $2\times 10^{25}$ $\sim$ $2\times 10^{26}$
+$c_{\text{V}_{\text{O}}^{..},\text{b}}(\infty)$/ $\text{m}^{-3}$ $0.5\times c_{\text{dop},\text{b}}(\infty)$
+$\Delta g_{\text{V}_{\text{O}}^{..}}$/eV -0.5 $\sim$ -1.75
+$\Delta g_{\text{dop}}$/eV -0.2 $\sim$ -0.7
+$N_{\text{V}_{\text{O}}^{..},\text{b}}$/ $\text{m}^{-3}$ $5.1\times 10^{28}$
+$N_{\text{dop},\text{b}}$/ $\text{m}^{-3}$ $1.68\times 10^{28}$
+$N_{\text{V}_{\text{O}}^{..},\text{c}}$/ $\text{m}^{-3}$ $1\times 10^{27}$
+$N_{\text{dop},\text{c}}$/ $\text{m}^{-3}$ $1.68\times 10^{27}$
+To investigate the formation of asymmetric SCLs induced by solute drag effects under quasi-equilibrium conditions, a constant core velocity (${v}_{\text{c}}$) is prescribed as a parameter in the simulations.
+The core velocity ranges from 0.01 $\text{nm}\,\text{s}^{-1}$ to 100 $\text{nm}\,\text{s}^{-1}$ to study its influence on the SCL.
+The simulation results, shown in Fig.
+4, illustrate the distributions of acceptor dopant concentration, oxygen vacancy concentration, charge density, and electrostatic potential as functions of $x$.
+The charge density is calculated as $\rho(x)=e(z_{\text{V}_{\text{O}}^{..}}c_{\text{V}_{\text{O}}^{..}}+z_{\text{dop}}c_{\text{dop}})$.
+The GB core moves from left to right along the $x$-axis.
+The acceptor dopant plays a crucial role in solute drag effects due to its significantly lower diffusivity compared to oxygen vacancy.
+Adjusting the segregation energy of the dopant significantly alters its concentration distribution.
+To examine the impact of segregation behavior, two segregation energies for the acceptor dopant, 0 eV and 0.5-0.5 eV, are selected to generate different SCLs.
+The segregation energy of oxygen vacancy is fixed at 1.5-1.5 eV.
+In the following discussion, the region to the left of the GB core is referred to as the growing grain, while the region to the right is referred to as the shrinking grain.
+The SCLs at the equilibrium state (i.e., $v_{\text{c}}$ = 0 $\text{nm}\,\text{s}^{-1}$) are shown in Fig.
+4(a) as a reference.
+The two different segregation energies for the acceptor dopant lead to distinct concentration distributions, as illustrated in Fig.
+4(a1).
+For $\Delta g_{\text{dop}}=0$ eV, the acceptor dopant concentration exhibits a depression at the GB core, while $\Delta g_{\text{dop}}=-0.5$ eV results in segregation.
+Moreover, the charge density in the GB core is higher for $\Delta g_{\text{dop}}=0$ eV than for $\Delta g_{\text{dop}}=-0.5$ eV.
+Consequently, the GB potential for $\Delta g_{\text{dop}}=-0.5$ eV is smaller than that for $\Delta g_{\text{dop}}=0$ eV.
+Compared with symmetric SCLs, a moving GB core introduces notable differences.
+First, the concentration profiles of the dopant are highly sensitive to GB velocity.
+The symmetric pattern begins to deteriorate even at very low velocities, such as 0.01 $\text{nm}\,\text{s}^{-1}$, as shown in Fig.
+4(b1).
+As a result, the dopant accumulation zone increases in the growing grain but decreases in the shrinking grain.
+At higher core velocities, dopant cations fail to keep up with the movement of the GB core, leading to a significant decrease in dopant concentration within the core, as illustrated in Fig.
+4(c1), (d1), and (e1).
+When the core velocity reaches $v_{\text{c}}$ = 100 $\text{nm}\,\text{s}^{-1}$, the dopant concentration profile becomes nearly flat, as depicted in Fig.
+4(e1).
+Second, significant asymmetry in the oxygen vacancy concentration arises only at very high GB velocities, due to its high diffusivity.
+As depicted in Fig.
+4(b2), (c2), and (d2), the oxygen vacancy concentration profiles exhibit only slight asymmetry at lower velocities.
+However, when $v_{\text{c}}$ = 100 $\text{nm}\,\text{s}^{-1}$, the concentration of oxygen vacancy in the depletion zone of the growing grain increases notably, as shown in Fig.
+4 (e2).
+Simultaneously, the depletion zone in the growing grain becomes substantially elongated, whereas it contracts in the shrinking grain.
+Third, as the GB velocity increases, not only does the GB potential rise, but a difference in electrostatic potential between the growing and shrinking grains also emerges, as shown in Fig.
+4(b4)–(e4).
+Additionally, the decay length of the electrostatic potential in the growing grain becomes significantly longer when GB velocity increases, resulting in an elongated SCL length in the growing grain.
+The segregation energy of the acceptor dopant plays a crucial role in SCL formation when the GB core is moving.
+Different values of $\Delta g_{\text{dop}}$, namely 0.5-0.5 eV and 0 eV, lead to distinct dopant concentration distributions.
 In Fig.
-3(c), (d), and (e), we present the concentration profiles of oxygen vacancies and dopant, as well as the distribution of electrostatic potential in the bulk phase for $\Delta g_{\text{dop}}$ = 0 eV, -0.125 eV and -0.375 eV, respectively.
-Due to symmetry, we only depict half of the SCL in one bulk phase.
-At $\Delta g_{\text{dop}}$ = 0 eV, phase-field results predict the formation of a dopant accumulation zone near the core.
-The dopant concentration in the core can be much lower than in the accumulation zone, leading to the formation of a nonmonotonic curve as shown in Fig.
-3(c1).
-Contrarily, in Virkant’s findings [38], the expected accumulation zone at $\Delta g_{\text{dop}}$=0 eV is replaced by a depletion zone.
-As $\Delta g_{\text{dop}}$ decreases to -0.125 eV, dopant concentration in the core increases significantly, causing the formation of monotonic concentration profile.
-Further decreasing $\Delta g_{\text{dop}}$ to -0.375 eV, the dopant concentration in the core increases to $N_{\text{dop},\text{b}}$.
-It is evident that the accumulation zone becomes lower than 0 eV, more exactly -0.125 eV.
-Additionally, near the core, the concentration profile of dopant exhibits a distinct kink, indicating a notable change in the gradient.
-The oxygen vacancy concentration profiles of are plotted in Fig.
-3(c2), (d2) and (e2).
-Depletion appears for all $\Delta g_{\text{dop}}$.
-Notably, as $\Delta g_{\text{dop}}$ decreases, there is an increase in the concentrations within these depletion zones.
-In Fig.
-3(c3), (d3) and (e3), we present a comparison of the electrostatic potential across the SCL with analytical solutions.
-The phase-field simulation results agree well with these analytical predictions.
-Decreasing $\Delta g_{\text{dop}}$ causes lower electrostatic potential distribution near the core.
-Moreover, we also present the net charge area density $Q_{\mathrm{net}}$ in Fig.
-3(f).
-The trend of the curve obtained via phase-field results agrees with the analytical predictions.
-Analogous to the MS case, phase-field results are smaller than analytical predictions.
-Figure 3: Comparison of phase-field (PF) simulation results with the sharp interface (SI) results via FEM method for the GC case.
-Figures (a) and (b) showcase the variations in grain boundary potential, dopant concentration, and oxygen vacancy concentration in the core as functions of $\Delta g_{\text{dop}}$.
-Figures (c), (d), and (e) detail the spatial distributions of dopant and oxygen vacancy concentrations in the bulk phase and the electrostatic potential.
-Figure (f) depicts the area density of net charge in the core, also as a function of $\Delta g_{\text{dop}}$.
-$\Delta g_{\text{V}_{\text{O}}^{..}}$ was kept at -1.0 eV for all calculations.
-5.2 GB space charge layers at quasi-equilibrium
-In this section, we investigate systematically the solute drag effect caused by an asymmetric SCL at a quasi-statically moving GB in a STO bicrystal.
-The solute drag phenomenon arises when point defects, segregated within a moving core in a quasi-equilibrium state, lead to additional free energy dissipation to maintain motion.
-At low core velocities, these defects can keep pace with the core, significantly increasing the necessary driving force.
-However, beyond a critical velocity, the defects lag behind, remaining in the bulk phases and causing the breakdown of symmetric concentration and electrostatic potential profiles across the SCL.
-To investigate the formation of asymmetric SCL due to solute drag effects under quasi-equilibrium conditions, a constant core velocity (${v}_{\text{c}}$) along the simulation domain is assigned as a parameter in the simulations.
-The velocity of the core ranges from 0.01 $\text{nm}\,\text{s}^{-1}$ to 100 $\text{nm}\,\text{s}^{-1}$ to study its influence on the SCL.
-The defect chemistry parameters $N_{\text{V}_{\text{O}}^{..},\text{c}}$, $N_{\text{V}_{\text{O}}^{..},\text{b}}$, $N_{\text{dop},\text{c}}$ and $N_{\text{dop},\text{b}}$ in this case are $1\times 10^{27}$ $\text{m}^{-3}$, $5.12\times 10^{28}$ $\text{m}^{-3}$, $1.68\times 10^{27}$ $\text{m}^{-3}$ and $1.69\times 10^{28}$ $\text{m}^{-3}$, respectively.
-Simulations are conducted at 1623 K, with a relative permittivity ($\epsilon_{\text{r}}$) of STO set to 56, based on [38].
-The diffusivities of oxygen vacancies and dopant in the bulk phase are specified as $1.35\times 10^{-9}$ $\text{m}^{2}\text{s}^{-1}$ and $1.3\times 10^{-11}$ $\text{m}^{2}\text{s}^{-1}$, respectively [38].
-The simulation domain and boundary conditions remain consistent with the previous case.
-In alignment with Vikrant [38], the total driving force $F_{\text{T}}$ for the core motion is given by
-$F_{\text{T}}=\int_{-\infty}^{\infty}\frac{v_{c}}{L}\left(\frac{\partial\eta}{\partial x}\right)^{2}\text{d}x+\int_{-\infty}^{\infty}{\mu_{\text{V}_{\text{O}}^{..}}}\frac{\partial C_{\text{V}_{\text{O}}^{..}}}{\partial x}\text{d}x+\int_{-\infty}^{\infty}{\mu_{\text{dop}}}\frac{\partial C_{\text{dop}}}{\partial x}\text{d}x.$ (60)
-The second and third terms in Eq. 60 denote as the electrochemical drag force, which accounts for the electrostatic and chemical contributions. The integration in Eq. 60 are solved numerically.
-In Fig.
-4, we illustrate the concentration profiles of oxygen vacancies and dopant, as well as the charge density and electrostatic potential as functions of $x$.
-From Fig.
-4(a) to Fig.
-4(e), we progressively increase the velocity of the core from 0 $\text{nm}\,\text{s}^{-1}$ to 100 $\text{nm}\,\text{s}^{-1}$.
-The direction of the core motion is from left to right along $x$-axis.
-In the following context, the left side of the core is named as the growing grain, while the right side is called the shrinking grain.
-In Fig.
-4(a), the symmetric SCL formation is presented at $v_{c}=0$ $\text{nm}\,\text{s}^{-1}$.
-However, the symmetric pattern begins to deteriorate even at very low velocity, e.g.
-0.01 $\text{nm}\,\text{s}^{-1}$, as shown in Fig.
-4(b1), (b2) and (b3).
-Compared with the concentration profiles at an equilibrium state, the accumulation zone of dopant in the growing grain is reduced, while that of the shrinking grain increases.
-This asymmetry arises, since during the core motion, dopant cations in the shrinking grain are captured by core due to their low diffusivity.
-Driving the motion of extra dopant cations within the core demands extra free energy dissipation.
-Therefore, the dopant cations flowing from the core region into the growing grain lead to the enhancement of the accumulation zone in the growing grain.
-However, the concentration of dopant in the core decreases from $7.8\times 10^{25}$ $\text{m}^{-3}$ ($v_{c}=0$ in Fig.
-4(a1)) to $6.8\times 10^{28}$ $\text{m}^{-3}$ at $v_{c}=0.01$ $\text{nm}\,\text{s}^{-1}$.
-In the growing grain far from the core, dopant concentration reaches $7.2\times 10^{25}$ $\text{m}^{-3}$, which is higher than equilibrium concentration, i.e.
-$c_{\text{dop},\text{b}}(\infty)$, as shown in Fig.
-4(a1).
-In Fig.
-4(b2), we plot the calculated charge density $\rho=e(z_{\text{dop}}c_{\text{dop}}+z_{\text{V}_{\text{O}}^{..}}c_{\text{V}_{\text{O}}^{..}})$.
-The charge density in the core decreases.
-Compared with the equilibrium case, we also observe the formation of less (more) negative charged zone in the shrinking (growing) grain, respectively.
-In Fig.
-4(b3), the grain boundary potential increases from 0.248 V ($v_{c}=0$ $\text{nm}\,\text{s}^{-1}$) to 0.265 V at $v_{c}=0.01$ $\text{nm}\,\text{s}^{-1}$.
-Moreover, an electrostatic potential different of 0.011 V develops between the shrinking and growing grains.
-When we increase the velocity of the core to 0.5 $\text{nm}\,\text{s}^{-1}$, dopant cations are not capable of catching the movement of core, as depicted in Fig.
-4(c1).
-Consequently, the dopant concentration within the core decreases substantially, resulting in a smaller depletion zone ahead of the core and a reduced accumulation zone behind it.
-Similarly, the concentration of oxygen vacancies within the core also decreases significantly.
-Correspondingly, as shown in Fig.
-4(c2), the charge density in the core drops to $8.6\times 10^{-7}$ $\text{C}\text{m}^{-3}$, which contrasts with the increase in charge density in the core observed in Virkant’s work [38] at intermediate velocities.
-We attribute the difference to the defect chemistry incorporated in the present phase field model.
-When the core velocity increases to 2 $\text{nm}\,\text{s}^{-1}$, the segregation of dopant in the core becomes negligible, as shown in Fig.
-4(d1).
-The concentration of dopant on both the left and right sides of the bulk phase returns to the equilibrium value, i.e.
-$c_{\text{V}_{\text{O}}^{..},\text{b}}(\infty)$.
-In Fig.
-4(d2), due to the continuous decrease in charge density within the core, the negatively charged zone is diminished.
-Furthermore, the grain boundary potential increases to 0.446 V and the electrostatic potential difference between the left and right sides reaches 0.122 V.
-At a high core velocity, such as 100 $\text{nm}\,\text{s}^{-1}$, the dopant concentration profile becomes completely flat, while the oxygen vacancy concentration profile exhibits strong asymmetry, as depicted in Fig.
-4(e1).
-The concentration of oxygen vacancies within the core decreases significantly to $2.46\times 10^{26}$ $\text{m}^{-3}$.
-In the shrinking grain, the depletion zone of oxygen vacancies remains sharp, about 2.5 nm, but it extends to approximately 20 nm in the growing grain.
-Consequently, as shown in Fig.
-4(e2), the negatively charged zone in the shrinking grain is much smaller, and the negative charge region in the growing grain is negligible.
-In Fig.
-4(e3), the electrostatic potential is illustrated.
-The grain boundary potential increases to 1.02 V and decays very slowly in the growing grain.
-In addition, the electrostatic potential difference becomes more pronounced, reaching 0.995 V.
-In summary, at small core velocity, the asymmetric formation of SCL is mainly caused by dopant coming apart from the core.
-The charge density in the negatively charged SCL and in the positively charged core both decrease.
-As a result, the electrostatic potential in the shrinking grain increases to a positive value.
-At high core velocity, dopant cations do not segregate to the core.
-The oxygen vacancy concentration profile is strongly asymmetric.
-This asymmetry causes an extended diffusion layer in the growing grain and elevates the electrostatic potentials in both the core and the shrinking grain.
-Figure 4: Asymmetric SCL formation in a STO bicrystal system with $c_{\text{dop},\text{b}}(\infty)=7.0\times 10^{25}$ $\text{m}^{-3}$, $\Delta g_{\text{V}_{\text{O}}^{..}}=-1.5$ eV and $\Delta g_{\text{dop}}=-0.5$ eV at quasi-equilibrium state under different constant core velocity (a) $v_{\text{c}}=0$ $\text{nm}\,\text{s}^{-1}$, (b) $v_{\text{c}}=0.01$ $\text{nm}\,\text{s}^{-1}$, (c) $v_{\text{c}}=0.5$ $\text{nm}\,\text{s}^{-1}$, (d) $v_{\text{c}}=2$ $\text{nm}\,\text{s}^{-1}$, (e) $v_{\text{c}}=100$ $\text{nm}\,\text{s}^{-1}$.
-Each row has three sub-figures, depicting the distribution of dopant and oxygen vacancy concentrations, total charge density and electrostatic potential, respectively.
-In order to further investigate solute drag effects and identify the critical core velocity at which a velocity jump occurs during sintering process, we present the core velocity as a function of the total driving force calculated via Eq.
-60 in Fig.
-5.
+4(b1), when the GB velocity $v_{\text{c}}$ is $0.01$ $\text{nm}\,\text{s}^{-1}$, the accumulation zones in the SCL for both segregation energies exhibit the same trend.
+However, as $v_{\text{c}}$ increases to $0.5$ $\text{nm}\,\text{s}^{-1}$ and 2 $\text{nm}\,\text{s}^{-1}$, notable differences emerge as shown in Fig.
+4(c1) and (d1).
+For $\Delta g_{\text{dop}}=-0.5$ eV, the accumulation zone in the shrinking grain transforms into a depletion zone, where the dopant concentration becomes lower than in the bulk.
+Meanwhile, the accumulation zone in the growing grain persists.
+In contrast, for $\Delta g_{\text{dop}}=0$ eV, the opposite behavior is observed.
+The accumulation zone remains in the shrinking grain, while in the growing grain, it diminishes into a depletion zone.
+When $v_{\text{c}}$ = 100 $\text{nm}\,\text{s}^{-1}$, the dopant concentration becomes flat for both cases.
+The influence of the segregation energy of the acceptor dopant on the electrostatic potential is significant only at low GB velocities.
+A more negative $\Delta g_{\text{dop}}$ leads to a smaller GB potential.
+At high GB velocities, where the dopant concentration profiles flatten, their effect on the electrostatic potential distribution diminishes, as shown in Fig.
+4(c4), (d4), and (e4).
+Figure 4: Asymmetric SCL formation in a STO bicrystal system with $c_{\text{dop},\text{b}}(\infty)=7.0\times 10^{25}$ $\text{m}^{-3}$, $\Delta g_{\text{V}_{\text{O}}^{..}}=-1.5$ eV and two $\Delta g_{\text{dop}}=$ 0 and -0.5 eV at quasi-equilibrium state under different constant core velocity (a) $v_{\text{c}}=0$ $\text{nm}\,\text{s}^{-1}$, (b) $v_{\text{c}}=0.01$ $\text{nm}\,\text{s}^{-1}$, (c) $v_{\text{c}}=0.5$ $\text{nm}\,\text{s}^{-1}$, (d) $v_{\text{c}}=2$ $\text{nm}\,\text{s}^{-1}$, (e) $v_{\text{c}}=100$ $\text{nm}\,\text{s}^{-1}$.
+Each row has four sub-figures, depicting the distribution of dopant concentration, oxygen vacancy concentration, total charge density and electrostatic potential, respectively.
+In this figure, blue line indicates $\Delta g_{\text{dop}}$ = -0.5 eV, while red line is $\Delta g_{\text{dop}}$ = 0 eV.
+The influence of defect chemistry parameters, especially segregation energy of dopant, on the formation behavior of different SCLs has been demonstrated.
+In order to further investigate solute drag effects and identify the critical core velocity at which a velocity jump occurs during sintering process, we present the core velocity as a function of the total driving force calculated via the following equation taken from Ref.
+vikrant2020electrochemical
+$\begin{split}F_{\text{T}}&=\int_{-\infty}^{\infty}\frac{v_{c}}{L}\left(\frac{\partial\eta}{\partial x}\right)^{2}\text{d}x+\int_{-\infty}^{\infty}{\mu_{\text{V}_{\text{O}}^{..}}}\frac{\partial C_{\text{V}_{\text{O}}^{..}}}{\partial x}\text{d}x+\int_{-\infty}^{\infty}{\mu_{\text{dop}}}\frac{\partial C_{\text{dop}}}{\partial x}\text{d}x\\ &=\int_{-\infty}^{\infty}\frac{v_{c}}{L}\left(\frac{\partial\eta}{\partial x}\right)^{2}\text{d}x-\int_{-\infty}^{\infty}\left[C_{\text{V}_{\text{O}}^{..}}-C_{\text{V}_{\text{O}}^{..},\text{b}}(\infty)\right]\frac{\partial\mu_{\text{V}_{\text{O}}^{..}}}{\partial x}\text{d}x-\int_{-\infty}^{\infty}\left[C_{\text{dop}}-C_{\text{dop},\text{b}}(\infty)\right]\frac{\partial\mu_{\text{dop}}}{\partial x}\text{d}x.\end{split}$ (65)
+The second and third terms in Eq.
+65 denote as the electrochemical drag force, which accounts for the electrostatic and chemical contributions when the expressions of electrochemical potential is employed.
+The integration in Eq.
+65 are solved numerically.
 The segregation energy difference, $\Delta g_{\text{dop}}$ and $\Delta g_{\text{V}_{\text{O}}^{..}}$, allow direct tuning of the concentrations of oxygen vacancy and dopant in the core, thereby impacting the critical total driving force required for a specific core velocity.
@@
 The solute drag effect therefore becomes negligible.
+Therefore, decreasing the segregation energy of oxygen vacancy could be an effective way to eliminate the solute drag effects.
 Fig.
@@
 5(a), we can evidently observe velocity jumps for all $\Delta g_{\text{dop}}$.
-As we decrease $\Delta g_{\text{dop}}$ to less negative value, such as from -0.2 eV to -0.3 eV, the enhancement of the solute drag effect is not significant.
+When the segregation energy of dopant decrease to -0.2 eV, we can still observe the solute drag effects.
 In Fig.
@@
 We fix $\Delta g_{\text{V}_{\text{O}}^{..}}$ and $\Delta g_{\text{dop}}$ at -1.5 eV and -0.5 eV.
+We clearly find the solute drag effect is strongly dependent on dopant concentration as evidenced in Ref.
+zahler2023grain .
+When $c_{\text{dop},\text{b}}=2\times 10^{26}\text{m}^{-3}$ (corresponding to 0.12% site fraction), solute drag effect still remains.
 As $c_{\text{dop},\text{b}}(\infty)$ increases, a larger driving force is necessary to counterbalance the solute drag effect to achieve the same core velocity.
@@
 In Fig.
-5(a), we observe GB core velocity jumps when $\Delta g_{\text{V}_{\text{O}}^{..}}=$1.5 eV and -1.75 eV.
+5(a), we observe GB core velocity jumps when $\Delta g_{\text{V}_{\text{O}}^{..}}=$ -1.5 eV and -1.75 eV.
 The critical velocities for different $\Delta g_{\text{V}_{\text{O}}^{..}}$ are approximately 0.35 $\text{nm}\,\text{s}^{-1}$.
@@
 5(c), a larger $c_{\text{dop},\text{b}}(\infty)$ leads to a higher critical velocity.
+In summary, the formation of SCLs at the quasi-equilibrium state has been systematically investigated under varying GB velocities and defect chemistry parameters.
+We conclude that (i) higher GB velocities lead to reduced dopant segregation in the GB core, more pronounced asymmetry in the oxygen vacancy concentration profiles, increased GB potential, and an elongated SCL length in the growing grain.
+(ii) the segregation energy of the dopant influences the concentration distribution pattern within the SCL and the GB potential, particularly at intermediate GB velocities.
+However, at extremely high GB velocities, its impact becomes negligible.
+(iii) the fundamental relationship between defect chemistry parameters and solute drag effects has been elucidated.
+When $\lvert\Delta g_{\text{V}_{\text{O}}^{..}}\rvert$ is smaller than 0.5 eV, the solute drag effect is negligible.
+However, at low $\lvert\Delta g_{\text{dop}}\rvert$ and dopant concentrations, minimal solute drag effects can still be observed.
+Thus, reducing the segregation energy of oxygen vacancy emerges as an effective strategy to minimize solute drag effects.
 Figure 5: The total driving force on a GB with a constant velocity with (a) $\Delta g_{\text{dop}}=-0.5$ eV, $c_{\text{dop},\text{b}}(\infty)=7\times 10^{25}$ $\text{m}^{-3}$ and different $\Delta g_{\text{V}_{\text{O}}^{..}}$, (b) $\Delta g_{\text{V}_{\text{O}}^{..}}=-1.5$ eV, $c_{\text{dop},\text{b}}(\infty)=7\times 10^{25}$ $\text{m}^{-3}$ and different $\Delta g_{\text{dop}}$, (c) $\Delta g_{\text{V}_{\text{O}}^{..}}=-1.5$ eV, $\Delta g_{\text{dop}}=-0.5$ eV and different $c_{\text{dop},\text{b}}(\infty)$.
@@
 4 e) are not marked.
-5.3 Abnormal grain growth
-After systematically benchmarks and knowledge gained from quasi-static single GB cases, we investigate different abnormal grain growth patterns of materials with different defect segregation energies $\Delta g_{\text{dop}}$, $\Delta g_{\text{V}_{\text{O}}^{..}}$ and bulk defect concentrations such as $c_{\text{dop},\text{b}}$.
+5.3 Polycrystalline simulations during sintering
+After systematically benchmarks and knowledge gained from quasi-static single GB cases, we investigate different grain growth patterns of materials with only consideration of solute drag effects.
+with different defect segregation energies $\Delta g_{\text{dop}}$, $\Delta g_{\text{V}_{\text{O}}^{..}}$ and bulk defect concentrations such as $c_{\text{dop},\text{b}}$.
 The segregation of point defects in the core plays a significant role in influencing grain growth kinetics.
 To investigate the grain size distribution of STO polycrystalline material after sintering, we conduct 2D phase-field simulations.
-Note that three-dimensional phase-field simulations of the proposed model are achievable but require significantly more computational resources, which will be considered in the near future.
+Note that three-dimensional phase-field simulations of the proposed model are achievable but require significantly more computational resources.
 In fact, to reveal the fundamental mechanisms, 2D simulations should be sufficient and allow for a comprehensive parameter study given the limited computational resources.
-The simulation box size is set to 390 nm $\times$ 390 nm with a grid resolution of 1000 $\times$ 1000 points.
-The polycrystalline microstructure is initialized using a Voronoi tessellation method, where each grain was assigned a unique grain index.
-We started with an initial number of 1000 grains, and the random seed used for generating the polycrystalline microstructure is set to 42.
-To optimize memory usage during simulations, we also utilize grain tracker algorithm [34].
-In the simulations, we set $N_{\text{V}_{\text{O}}^{..},\text{c}}$, $N_{\text{V}_{\text{O}}^{..},\text{b}}$, $N_{\text{dop},\text{c}}$ and $N_{\text{dop},\text{b}}$ as $1\times 10^{27}$ $\text{m}^{-3}$, $5.12\times 10^{28}\text{m}^{-3}$, $1.68\times 10^{27}$ $\text{m}^{-3}$ and $1.69\times 10^{28}$ $\text{m}^{-3}$.
-The simulations are conducted at a temperature of 1623 K, where the non-Arrhenius behavior of the grain growth is found [28].
-Periodic boundary conditions are applied to all edges of the simulation domain for the phase-fields and concentration fields.
-For the electrostatic potential field, the bottom-left corner is grounded, while the other edges are set with the periodic boundary conditions.
-Furthermore, all the parameters used in the simulation are normalized.
-The lattice constant $a$ = $3.9\times 10^{-10}$ m is used for the length normalization.
-Other normalization factors include e.g.
-$f_{0}=\kappa/a^{2}$=$2.3\times 10^{9}$ $\text{J}\text{m}^{-3}$ for the energy density, $t_{0}=1/(Lf_{0})=0.0043$ s for the interface mobility, and $\phi_{0}=f_{0}V_{\text{m}}/\mathcal{F}$= 0.858 V for the electrostatic potential.
-We present first the microstructure evolution of the grain growth process under specific conditions $\Delta g_{\text{V}_{\text{O}}^{..}}$ = -1.5 eV, $\Delta g_{\text{dop}}$=-0.4 eV and $c_{\text{dop},\text{b}}=7\times 10^{25}$ $\text{m}^{-3}$ when the sintering time is $t/\text{h}=\{0.067,0.33,0.67,1.0\}$.
-The average area of grain at time $t=0$ is <S>, i.e., 152.1 $\text{nm}^{2}$.
-When $t$=0.067 h, 0.33 h and 0.67 h, the grain size increases while the area fraction of small grain decreases.
-The grain size distribution starts to transition from a mono-modal to a bimodal pattern.
-When $t$=1 h, the largest grain area is about 32 times larger than <S>.
-The bimodal grain size distribution is clearly visible.
-Figure 6: The microstructure evolution of grain growth process at (a) 0.067 h, (b) 0.33 h, (c) 0.67 h and (d) 1 h, with defect chemistry parameters set at $\Delta g_{\text{V}_{\text{O}}^{..}}$ = -1.5 eV, $\Delta g_{\text{dop}}$=-0.4 eV and $c_{\text{dop},\text{b}}=7\times 10^{25}$ $\text{m}^{-3}$.
+The defect chemistry parameters used in this section are tabulated in Table 4.
+Then, we describe the FEM settings for the 2D phase-field simulations as follows.
+The simulation domain is 390 nm $\times$ 390 nm, with a uniform mesh spacing of $\Delta x=\Delta y=0.39$ nm.
+The GB core width is set to 0.78 nm, consistent with the value used in the 1D simulations to ensure accurate reproduction of the SCLs.
+The polycrystalline microstructure is initialized using a Voronoi tessellation method, where each grain is assigned a unique grain index.
+The initial configuration consists of 1000 grains, and the random seed for generating the polycrystalline microstructure is set to 42.
+To optimize memory usage during the simulations, the grain tracker algorithm is employed yang20193d .
+Periodic boundary conditions are applied to all edges of the simulation domain for both the phase-field and concentration fields.
+For the electrostatic potential field, the bottom-left corner is grounded, while the remaining edges are set with periodic boundary conditions.
+Each simulation is performed on 8 nodes (a total of 768 cores) with 3072 gigabytes of memory, requiring approximately 30 days to complete.
+5.3.1 Skewed grain size distribution
+Compared with normal grain growth pattern, the grain size distribution is skewed when the defect segregation in the GB core is considered as shown in Fig.
+6.
+To clearly distinguish the grain size distribution, a histogram of grain area fraction is constructed.
+The bin width, denoted as $\xi_{\text{bin}}$, is determined using the Freedman-Diaconis rule freedman1981histogram .
+Subsequently, the grain size distribution is fitted as a sum of two weighted lognormal distribution functions vikrant2020electrochemicallyActa , given by $\text{PDF}(\hat{S})=\hat{k}_{1}\mathcal{N}_{1}(\hat{S},\hat{\mu}_{1},\hat{\sigma}_{1})+\hat{k}_{2}\mathcal{N}_{2}(\hat{S},\hat{\mu}_{2},\hat{\sigma}_{2})$, where $\hat{S}$ is the normalized grain area and can be expressed as $\hat{S}=S/\langle S\rangle/\xi_{\text{bin}}$.
+$\langle S\rangle=152.1$ $\text{nm}^{2}$ is the average area of grain at time $t=0$.
+$\hat{k}_{1}$ and $\hat{k}_{2}$ are the proportionality coefficients.
+$\hat{\mu}_{1}$ and $\hat{\mu}_{2}$ are the mean grain areas.
+$\hat{\sigma}_{1}$ and $\hat{\sigma}_{2}$ are the standard deviations.
+the lognormal distributions $\mathcal{N}_{1}$ and $\mathcal{N}_{2}$ share the general form $\mathcal{N}=\frac{1}{\hat{S}\hat{\sigma}\sqrt{2\pi}}\exp\left[-\frac{(\ln\hat{S}-\hat{\mu})^{2}}{2\hat{\sigma}^{2}}\right]$.
+The formation of skewed grain size distribution during the sintering process is demonstrated in Fig.
+6.
+The dopant concentration is $2\times 10^{26}$ $\text{m}^{-3}$, corresponding to a 1.1% site fraction, with segregation energies of oxygen vacancy and acceptor dopant set to -1.5 eV and -0.5 eV, respectively.
+At a sintering time of 1 h, the grain size distribution is nearly monomodal, as shown in Fig.
+6(a2), with only one peak observed in the fitted lognormal function.
+By 1.5 h, a transition from monomodal to skewed grain size distribution occurs, as illustrated in Fig.
+6(b2), where two distinct peaks emerge.
+At 2 h and 3 h, the grain size distribution is skewed and do not follow log-normal distributions.
+The peak positions of the two grain populations are shown in Fig.
+7(a).
+For small grains, the peak position increases from 2.55 at 1.5 h to 3.07 at 3 h, while for large grains, it rises from 6.21 to 10.64 over the same period.
+As sintering progresses, the size disparity between small and large grains becomes increasingly significant.
+In Fig.
+7(b), the time evolution of grain area is plotted for 10 small grains and 10 large grains.
+The unique grain numbers are labeled in Fig.
+6(d1), with black numbers indicating small grains and red numbers representing large grains.
+The results reveal that both small and large grains grow rapidly during the initial stage of sintering.
+However, as sintering progresses, the grain areas of small grains stabilize, while those of large grains continue to grow.
+Therefore, the growth velocity difference between two distinct grain populations leads to the formation of a skewed grain size distribution, which do not follow the normal grain growth.
+Moreover, the peak positions of the two grain populations in this skewed grain size distribution are influenced by various defect chemistry parameters, such as the segregation energies of dopant and oxygen vacancy, as well as the dopant concentration in the bulk phase.
+When these defect chemistry parameters enhance the solute drag effect, smaller grains become pinned earlier and remain stable during the sintering process, leading to a more noticeable difference between the two peak positions.
+In the following, we examine how the selection of defect chemistry parameters affects the grain growth processes.
+Figure 6: The microstructure evolution during the grain growth process is presented at (a1) 1 h, (b1) 1.5 h, (c1) 2 h, and (d1) 3 h, with defect chemistry parameters set to $\Delta g_{\text{V}_{\text{O}}^{..}}=-1.5$ eV, $\Delta g_{\text{dop}}=-0.5$ eV, and $c_{\text{dop},\text{b}}=2\times 10^{26}\text{m}^{-3}$.
 These parameters are consistent with those used in Fig.
-7(e).
-By 1h, a bimodal grain size distribution becomes apparent.
+8(f).
+The area fraction distribution is fitted by two lognormal distribution functions and shown in (a2)–(d2) to illustrate the skewed grain size distribution.
+The blue dashed lines represent the smaller grain population, while the red dashed lines correspond to the larger grain population.
+The brown solid lines indicate the combined distribution of both populations.
+Figure (a) also appears in Fig.
+8 (f) to compare with other simulations.
+Figure 7: (a) Illustration of the grain size peak positions for two grain populations: small grains and large grains.
+For sintering times shorter than 1 h, a monomodal grain size distribution is observed, with only a single peak present.
+As sintering progresses, a skewed grain size distribution emerges, characterized by two distinct peaks.
+The difference in peak positions between small and large grains becomes increasingly pronounced with longer sintering times.
+(b) Time evolution of grain areas for 20 specific grains.
+The unique grain numbers are labeled in Fig.
+6.
+Solid lines represent small grains, while dashed lines correspond to large grains.
+As sintering time increases, the grain areas of large grains grow significantly, while the grain areas of small grains remain pinned.
 5.3.2 Influence of segregation energies and bulk defect concentration
@@
 The microstructure of grains is presented in Fig.
-7(a)-(g) after 1 hour of sintering at 1623 K.
+8(a)-(g) after 1 hour of sintering at 1623 K.
 In Fig.
-7(a), (b) and (c), we increase $\Delta g_{\text{V}_{\text{O}}^{..}}$ from -1.5 eV to -0.5 eV when $\Delta g_{\text{dop}}$=-0.5 eV and $c_{\text{dop},\text{b}}=7\times 10^{25}$ $\text{m}^{-3}$.
-Bimodal grain size distribution is clearly observed in Fig.
-7(a2), (b2) and (c2).
-As $\Delta g_{\text{V}_{\text{O}}^{..}}$ becomes more negative, the peak in the range of $40<\frac{S}{<S>}<$ 4 induced by smaller grains becomes more pronounced.
-As predicted in Fig.
-5(a), solute drag effect is almost negligible for $\Delta g_{\text{V}_{\text{O}}^{..}}$=-0.5 eV.
-Nonetheless, the formation of solute cloud imprinted by vanished GBs is still detectable in Fig.
-7(c3) even though the solute drag effect is very weak.
+8(a), (b) and (c), we decrease the segregation energy of oxygen vacancy ($\lvert\Delta g_{\text{V}_{\text{O}}^{..}}\rvert$) from 1.5 eV to 0.5 eV when $\Delta g_{\text{dop}}$=-0.5 eV and $c_{\text{dop},\text{b}}=7\times 10^{25}$ $\text{m}^{-3}$.
+Skewed grain size distribution is clearly observed in Fig.
+8(a2).
+The reduction of oxygen vacancy segregation energy decreases not only the segregation concentration of oxygen vacancy but also that of the dopant in the GB core.
+Consequently, the solute drag effect is notably weakened.
+As shown in Fig.
+8(b2), when $\Delta g_{\text{V}_{\text{O}}^{..}}=-1.0\text{eV}$, two distinct grain populations are still observed.
+In contrast, Fig.
+8(c2) shows that the grain size distribution becomes nearly monomodal when $\Delta g_{\text{V}_{\text{O}}^{..}}=-0.5\text{eV}$, indicating that the solute drag effect is negligible in this case.
+This observation is consistent with the 1D phase-field simulation results presented in Fig.
+5.
+The concentration of dopant is presented in Fig.
+8(a3), (b3) and (c3), the formation of solute cloud imprinted by vanished GBs is clearly observed, especially in Fig.
+8(a3).
 Concentration of oxygen vacancy is detailed in Fig.
-7(a4), (b4) and (c4).
+8(a4), (b4) and (c4).
 Segregation is more significant for $\Delta g_{\text{V}_{\text{O}}^{..}}=-1.5$ eV.
@@
 The electrostatic potential distributions are depicted in Fig.
-7(a5), (b5) and (c5).
+8(a5), (b5) and (c5).
+The GB potential considerably decrease from (a5) to (c5).
 The negative potential bands can be also observed near the GB cores.
@@
 In Fig.
-7(a), (d) and (e), we keep $\Delta g_{\text{V}_{\text{O}}^{..}}$ and $c_{\text{dop},\text{b}}$ at -1.5 eV and $7\times 10^{25}$ $\text{m}^{-3}$ and test different $\Delta g_{\text{dop}}$ values.
-Whereas, in Fig.
-7(a), (f) and (g), $\Delta g_{\text{V}_{\text{O}}^{..}}$ and $\Delta g_{\text{dop}}$ are constant while $c_{\text{dop},\text{b}}$ increases from $2\times 10^{25}$ $\text{m}^{-3}$ to $2\times 10^{26}$ $\text{m}^{-3}$.
+8(a), (d), and (e), we fix $\Delta g_{\text{V}_{\text{O}}^{..}}$ = -1.5 eV and $c_{\text{dop},\text{b}}=7\times 10^{25}$ $\text{m}^{-3}$, and examine the influence of different $\Delta g_{\text{dop}}$ values.
+A larger $\lvert\Delta g_{\text{dop}}\rvert$ leads to more pronounced solute drag effects.
+In all cases, a skewed grain size distribution is observed, even when $\Delta g_{\text{dop}}$ = -0.2 eV, as shown in Fig.
+8(e).
+In Fig.
+8(d2), although the solute drag effect is more pronounced due to higher dopant segregation energy and the grain size distribution can be fitted by two lognormal distribution functions, the PDF($\hat{S}$) function still exhibits a monomodal shape.
+We expect that a skewed PDF($\hat{S}$) function that do not follow the normal grain growth will emerge if the sintering time is extended, as demonstrated in Fig.
+6.
+Additionally, the electrostatic potential distribution remains largely unchanged, as shown in Fig.
+8(a5), (d5), and (e5).
+For the low dopant segregation energy case ($\Delta g_{\text{dop}}=-0.2$ eV), the solute drag effect arises primarily from the electrochemical potential gradient terms (see Eq.
+65), which is contributed by the electrostatic potential part, rather than chemical part.
+In Fig.
+8(a), (f) and (g), $\Delta g_{\text{V}_{\text{O}}^{..}}$ and $\Delta g_{\text{dop}}$ are constant while $2\times 10^{25}$ $\text{m}^{-3}$, $7\times 10^{25}$ $\text{m}^{-3}$ and $2\times 10^{26}$ $\text{m}^{-3}$ are chosen for dopant concentration in bulk phase, corresponding to 0.11%, 0.42% and 1.1% site fractions.
 For all simulations, the solute drag effects from dopant are pronounced.
@@
 In Fig.
-7(e2) and (g2), the bimodal grain size distributions are presented.
-However, when $\Delta g_{\text{dop}}=-0.6$ eV in Fig.
-7(d2) and $c_{\text{dop},\text{b}}=2\times 10^{26}$ $\text{m}^{-3}$ in Fig.
-7(f2), the unimodal grain size distribution exhibits, due to the more pronounced probability of small grain.
-The simulation results in Fig.
-7(f) demonstrate the high grain number, where the large grains are surrounded by small grains.
-Figure 7: Phase-field simulation results of abnormal grain growth in Fe-doped STO after 1 h sintering under the influence of different defect chemistry parameters.
+8(e2) and (g2), the skewed grain size distributions are presented.
+However, due to the more pronounced probability of small grain, longer sintering time is required for $c_{\text{dop},\text{b}}(\infty)$ = $2\times 10^{26}$ $\text{m}^{-3}$ to exhibit the skewed grain size distribution that do not follow the normal grain size distribution (also see Fig.
+6).
+The influence of different defect chemistry parameters on grain growth process are elucidated in this part.
+First, the segregation oxygen vacancy plays a dominant role in solute drag induced skewed grain size distribution.
+Reduction of the segregation energy of oxygen vacancy is an effective way to eliminate solute drag effects and the skewed grain size distribution to enhance the sintering capability of oxide ceramics.
+In our case, when $\lvert\Delta g_{\text{V}_{\text{O}}^{..}}\rvert$< 0.5 eV, the skewed grain growth behavior can be nearly avoided.
+Second, the segregation energy of dopant impacts the SCL formation and GB concentration of dopant.
+Increase segregation energy of dopant strongly decrease the grain growth rate.
+However, very low $\lvert\Delta g_{\text{dop}}\rvert$< still leads to skewed grain size distribution.
+Third, dopant concentration dependent solute drag effect is also verified in the present work, even very small dopant concentration (0.11%) can lead to the skewed grain size distribution that do not follow log-normal distributions.
+Figure 8: Phase-field simulation results of different grain growth patterns in Fe-doped STO after 1 h sintering under the influence of different defect chemistry parameters.
 5.3.3 Grain boundary potential distribution
 In Fig.
-8, we present the statistics of the average grain boundary potential in the GB core, denoted as $\bar{\phi}_{\text{core}}$.
+9, we present the statistics of the average grain boundary potential in the GB core, denoted as $\bar{\phi}_{\text{core}}$.
 The value of $\bar{\phi}_{\text{core}}$ is calculated as an average of the electrostatic potential along the GB core indicated by $\eta_{i}=0.5$ over individal GB length.
 Fig.
-8(a) illustrates the distribution of $\bar{\phi}_{\text{core}}$ for $\Delta g_{\text{V}_{\text{O}}^{..}}=-0.5$ eV, $\Delta g_{\text{dop}}=-0.5$ eV, $c_{\text{dop}}(\infty)=7\times 10^{25}$ $\text{m}^{-3}$ (refer to Fig.
-7(c) for the associated phase-field simulation result).
+9(a) illustrates the distribution of $\bar{\phi}_{\text{core}}$ for $\Delta g_{\text{V}_{\text{O}}^{..}}=-0.5$ eV, $\Delta g_{\text{dop}}=-0.5$ eV, $c_{\text{dop}}(\infty)=7\times 10^{25}$ $\text{m}^{-3}$ (refer to Fig.
+8(c) for the associated phase-field simulation result).
 When the solute drag effect is negligible, $\bar{\phi}_{\text{core}}$ is smaller than 0.05 V and approximately follows the normal distribution.
 In Fig.
-8(b), the distribution of $\bar{\phi}_{\text{core}}$ is presented for $\Delta g_{\text{V}_{\text{O}}^{..}}=-1.5$ eV, $\Delta g_{\text{dop}}=-0.5$ eV, $c_{\text{dop}}(\infty)=7\times 10^{25}$ $\text{m}^{-3}$ (refer to Fig.
-7(a)).
+9(b), the distribution of $\bar{\phi}_{\text{core}}$ is presented for $\Delta g_{\text{V}_{\text{O}}^{..}}=-1.5$ eV, $\Delta g_{\text{dop}}=-0.5$ eV, $c_{\text{dop}}(\infty)=7\times 10^{25}$ $\text{m}^{-3}$ (refer to Fig.
+8(a)).
 Here, $\bar{\phi}_{\text{core}}$ significantly increases.
-In addition, the distribution of $\bar{\phi}_{\text{core}}$ changes from mono-modal to bimodal, with one peak at 0.31 V and another at 0.35 V.
+In addition, the distribution of $\bar{\phi}_{\text{core}}$ changes from mono-modal to skewed, with one peak at 0.31 V and another at 0.35 V.
 In Fig.
-8(c), the parameters $\Delta g_{\text{V}_{\text{O}}^{..}}=-1.5$ eV, $\Delta g_{\text{dop}}=-0.5$ eV, $c_{\text{dop}}(\infty)=2\times 10^{26}$ $\text{m}^{-3}$ are chosen (refer to Fig.
-7(f)).
+9(c), the parameters $\Delta g_{\text{V}_{\text{O}}^{..}}=-1.5$ eV, $\Delta g_{\text{dop}}=-0.5$ eV, $c_{\text{dop}}(\infty)=2\times 10^{26}$ $\text{m}^{-3}$ are chosen (refer to Fig.
+8(f)).
 In this case, the solute drag effect is the most pronounced in this study.
-Not only $\bar{\phi}_{\text{core}}$ increases but also the bimodal distribution of $\bar{\phi}_{\text{core}}$ becomes more distinct.
+Not only $\bar{\phi}_{\text{core}}$ increases but also the skewed distribution of $\bar{\phi}_{\text{core}}$ becomes more distinct.
 […]
 In addition, two distinct types of GBs can be observed Fig.
-9.
+10.
 The first type, characterized by a lower moving velocity, exhibits nearly symmetric concentration profiles across the GB core with a significantly higher segregation concentration of dopant.
@@
 Similarly, the oxygen vacancy distribution in this second GB core is also asymmetric, as illustrated in Fig.
-7(e).
+8(e).
 Fig.
-7(f) presents the electrostatic potential distributions across these two types of GB cores, where the second GB core displays a larger grain boundary potential and a more pronounced electrostatic potential difference.
+8(f) presents the electrostatic potential distributions across these two types of GB cores, where the second GB core displays a larger grain boundary potential and a more pronounced electrostatic potential difference.
+These two distinct types of GBs play different roles in the grain growth process.
+Fast moving GBs facilitate the formation of the large grain population during sintering, while low-moving GBs are responsible for pinning the small grain population.
 Figure 10: Two types of GBs are observed in the rectangular region highlighted in Fig.
-7(f).
+8(f).
 The distributions of dopant concentration, oxygen vacancy concentration, and electrostatic potential in this region are shown in (a), (b), and (c), respectively.
@@
 The relations between $v_{\text{c}}$ and $F_{\text{T}}$ as well as $\bar{\phi}_{\text{core}}$ and $F_{\text{T}}$ for each GB core are represented in Fig.
-10(b).
+11(b).
 Here, $F_{\text{T}}$ is the total driving force, numerically calculated via Eq.
-60.
+65.
 The solid lines indicate the phase-field simulation results in the bicrystal case (refer to the purple line in Fig.
@@
 In Fig.
-10(b), the relationships of $F_{\text{T}}$ vs.
+11(b), the relationships of $F_{\text{T}}$ vs.
 $v_{\text{c}}$ and $F_{\text{T}}$ vs.
@@
 The 14 different GBs, shown in Fig.
-9(a), are chosen from the rectangular region in Fig.
-7(f) in polycrystalline simulations.
+10(a), are chosen from the rectangular region in Fig.
+8(f) in polycrystalline simulations.
 The relations between $v_{\text{c}}$ and $F_{\text{T}}$ as well as $\bar{\phi}_{\text{core}}$ and $F_{\text{T}}$ are obtained from phase-field simulations in bicrystal and polycrystalline Fe-doped STO.
@@
 Other symbols represent 14 different GBs chosen in the polycrystalline case.
+In summary, we demonstrates the formation of a skewed grain size distribution in Fe-doped SrTiO3 during the sintering process, driven solely by the solute drag effect.
+Two types of GBs emerge during the growth process due to defect segregation in the GB core: fast-moving GBs and slow-moving GBs.
+Fast-moving GBs facilitate the growth of large grains, while slow-moving GBs act to pin small grains.
+The coexistence of these GBs ultimately leads to the skewed grain size distribution.
+Furthermore, variations in defect chemistry significantly influence the resulting grain size distribution.
+To improve the sintering process, reducing the segregation energy of oxygen vacancies proves to be an effective approach.
+Lowering the dopant segregation energy and bulk concentration also mitigate solute drag effects.
+However, even at low dopant segregation energy and bulk dopant concentration, the skewed grain size distribution can still occur.
 6 Conclusion
@@
   • 1.
-    The proposed phase-field model reproduces a series of benchmark results at equilibrium of Fe-doped STO bicrystal.
+    With the consideration of distinct segregation energies and available site densities of oxygen vacancy and acceptor dopant in bulk and GB core, the proposed phase-field model reproduces a series of benchmark results at equilibrium of Fe-doped STO bicrystal.
 Simulation results, including the symmetric SCLs, grain boundary potentials, concentrations of oxygen vacancies and acceptor dopants in the GB core and the bulk phase as well as the electrostatic distributions, show good agreement with the analytical predictions in the MS and GC models.
@@
     Simulations are carried out comprehensively for a GB core moving with constant velocities under quasi-equilibrium state.
-At low core velocity, the accumulation zone in the shrinking grain decreases, while it increases in the growing grain, leading to the asymmetric SCL formation.
+At low core velocity, the asymmetric SCL forms.
 With increasing GB core velocity, the dopant segregation diminishes.
@@
   • 3.
-    Abnormal grain growth is observed even without distinguishing any grain orientation or considering any anisotropic GB mobility in the phase-field simulations at 1623K.
-It indicates that the solute drag effect alone can lead to bimodal grain size distribution.
+    Skewed grain size distribution is observed even without distinguishing any grain orientation or considering any anisotropic GB mobility in the phase-field simulations at 1623K.
+It indicates that the solute drag effect alone can lead to skewed grain size distribution that do not follow the log-normal distribution in the normal grain growth.
 As the formation energy differences of oxygen vacancies and acceptor dopants ($\Delta g_{\text{V}_{\text{O}}^{..}}$ and $\Delta g_{\text{dop}}$) become more negative, the grain growth velocity decreases.
@@
 The second type of GB core shows asymmetric concentration profiles and a higher moving velocity, contributing predominantly to grain coarsening.
+The coexistence of these grain boundaries results in grain growth behavior that deviates from the typical normal grain growth pattern..
   • 5.
-    Abnormal grain growth resulting from solute drag effects leads to a large variation of grain boundary potentials.
+    To improve the sintering process, reducing the segregation energy of oxygen vacancies proves to be an effective approach.
+Lowering the dopant segregation energy and bulk concentration also mitigate solute drag effects.
+However, even at low dopant segregation energy and bulk dopant concentration, the skewed grain size distribution can still occur.
+  • 6.
+    The skewed grain size distribution resulting from solute drag effects leads to a large variation of grain boundary potentials.
 Typically, smaller grains exhibit higher grain boundary potentials at the later stages of grain growth and form more blocking GBs in term of ionic conductivity.
@@
 The proposed phase-field model offers naturally a series of promising extensions.
-We can e.g.
-incorporate additional species, such as neutral dopants, electrons, and electron holes, into the current phase-field model to quantitatively investigate the formation of the SCL.
-Furthermore, GB orientation plays a crucial role in SCL formation, and both the interface energy and the defect chemistry parameters $\Delta g_{\text{V}_{\text{O}}^{..}}$ and $\omega_{c}N_{\text{V}_{\text{O}}^{..},\text{c}}$ depend on GB orientation [49].
-These complexities can be included in the phase-field model in the near future.
+First, we can e.g.
+incorporate additional species, such as the electrons, holes and the dopants with different valance states, into the current phase-field model to quantitatively investigate the formation of the SCL.
+Second, GB orientation plays a crucial role in SCL formation and different grain growth patterns, and both the interface energy and the defect chemistry parameters $\Delta g_{\text{V}_{\text{O}}^{..}}$ and $\omega_{c}N_{\text{V}_{\text{O}}^{..},\text{c}}$ depend on GB orientation de2019effect .
+Third, the assumptions of dilute and non-interacting defects become invalid at high defect concentrations, where defect–defect interactions must be accounted for.
+Various approaches have been proposed to address this complexity.
+One approach involves using a regular solution model to formulate the interaction energy contribution lund2021thermodynamically , vikrant2020electrochemical , vikrant2020electrochemicallyActa .
+Another approach, proposed by Mebane and De Souza (MDS) mebane2015generalised, provides expressions for the electrochemical potentials of oxygen vacancy and acceptor dopant while incorporating defect–defect interactions.
+However, MDS model does not include an explicit free energy formulation, which is essential for direct integration into phase-field models.
 Moreover, we assume in this paper only the $\text{Fe}^{3+}$ and ignore $\text{Fe}^{4+}$, which is reasonable for high temperature scenarios.
@@
 The $\text{Fe}^{3+}$ ions in the GB core can not easily follow the movement of the GB core due to their low diffusivity.
-In contrast, the rapid movement of electron holes can adjust the local $\text{Fe}^{3+}$/$\text{Fe}^{4+}$ ratio and establish an equilibrium charge distribution.
+In contrast, the rapid movement of electron holes can adjust the local $\text{Fe}^{3+}$$\text{Fe}^{4+}$ ratio and establish an equilibrium charge distribution.
 The solute drag effect may differ when a significant amount of $\text{Fe}^{4+}$ is considered.